A Bibliography of Publications in IEEE Computer: 1972–1979

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: https://www.math.utah.edu/~beebe/

21 December 2023 Version 2.13

Title word cross-reference

10th [554]. 12 [1467]. 16-bit [1329]. 165 [759]. 16K [1264]. 1971 [96, 114, 133]. 1972 [119, 182, 208]. 1973 [341]. 1974 [743, 963]. 1975 [859, 905, 912, 1011]. 1976 [1224]. 1979 [1418, 1455, 1458]. 1st [1426].

20 [96]. 203 [7]. 20th [711]. 256-Bit [71]. 25th [1210]. 2nd [870, 903, 1211].

3000 [940]. 3000CX [816, 847, 887, 917].

48 [1374]. 480 [1286].

6502 [1060]. 68000 [1376].

7 [762]. 70 [31]. 70/FJCC [31]. ’71 [104].

’72 [146, 201, 209]. 73 [263]. ’74 [388, 548]. ’75 [814, 851, 878, 910]. ’76 [1107, 1114]. ’79 [1427].

8089 [1406]. 8O [1212]. 8th [94].

9 [1213].

able [853, 954]. abroad [788]. Access [26, 70, 71, 156, 1396]. Acclaimed [1193]. Accomplishments [972]. achieving [1240]. ACM [204, 209, 290, 1114]. Acquisition [1009, 1193]. Acquisition/Development [1193]. Activities [46, 145, 613]. additions [700]. address [1399]. addressable [1385]. Addressed [850]. Ads [126, 152, 336, 601, 632, 698, 721, 749, 767, 802, 900, 988, 1069, 1118, 1145]. Advance [299, 487, 684, 1002, 1030]. advanced [426, 1156, 1296, 1403]. Advances [24, 984, 1236]. Advantages [1105]. Advertisement [17, 18, 36, 37, 60, 132, 313, 315, 319, 324, 338, 451, 472, 494, 525, 543, 550, 553, 559, 567– 569, 575–577, 583, 586, 602– 604, 619, 621, 622, 628, 634–636, 644, 645, 651, 655, 667, 668, 670, 674, 679, 680, 682, 683, 687, 697, 713, 723, 725, 735, 738, 741, 770, 779, 781, 803, 819, 849, 865, 873, 876, 885, 895, 911, 935, 942, 944, 952, 956, 959, 996, 1047, 1051, 1074, 1079, 1094, 1100, 1111, 1125, 1134, 1151]. advertisement [828, 829, 835, 844, 863, 868, 881, 894, 898, 925, 927, 947, 953, 993, 1005, 1021, 1044, 1066, 1070, 1088, 1101, 1121, 1144, 1148, 1193, 1202, 1204, 1212, 1221, 1228]. Advertisements [389, 413, 436, 454, 560, 581, 584, 589, 605, 609, 610, 618, 656]. Advertiser [306, 335, 401, 431, 449, 666]. Advertisers [381, 510, 769, 801, 831, 1224]. Aerospace [49, 588]. Affordable [849, 885]. again [536]. against [394]. Age [849, 885]. Agony [975]. AHPL [782]. Aid [822]. Aided [55, 77, 105, 219, 554, 919, 1288, 1347, 1372]. Air [117]. Air-Pollution [117]. Alamos [591]. Algorithm [1189, 1378]. algorithms [758]. Almost [1202]. Alternative [426, 1284, 1290]. Ambulatory [818, 1012]. America [1139]. Analog [1171, 1172, 1451]. analog/hybrid [1172]. Analysis [271, 362, 415, 660, 1010, 1013, 1015, 1163, 1272, 1279, 1288, 1316, 1388, 1396, 1419, 1420]. Analytic [941, 1360]. and/or [975]. Angiocardiographic [1013]. Anniversary [1210]. Announcement [603]. Announcements [563, 597, 629, 662, 694, 719, 746, 797, 826, 858, 892, 921, 948, 991, 1019, 1042, 1064, 1086, 1116, 1142, 1218, 1432, 1433, 1464, 1465]. Announcing [568, 576, 866, 882, 1423]. Annual [94, 232, 554, 1017, 1039, 1109, 1138]. any [1211]. Anyone [112]. API [1267]. Application [368, 414, 620, 646, 893, 1077, 1078].

Applications

[112, 137, 138, 148, 175, 180, 199, 210, 250, 275, 304, 332, 353, 375, 378, 398, 428, 447, 469, 491, 508, 523, 540, 564, 586, 593, 595, 598, 630, 663, 685, 695, 701, 720, 736, 747, 763, 765, 798, 818, 827, 860, 922, 949, 992, 1011, 1020, 1043, 1065, 1082, 1087, 1117, 1143, 1155, 1219, 1255, 1263, 1278, 1320, 1327, 1381, 1401, 1402, 1434, 1466]. Applied [220, 740, 982, 1375]. Applying [969]. Approach [1168, 1258, 1332, 1365]. Approaches [974]. Architectural [362, 1387]. Architecture [43, 192, 443, 463, 465, 466, 489, 534, 538, 649, 652, 852, 1017, 1035, 1137, 1138, 1155, 1229, 1237, 1238, 1251, 1303, 1306, 1321, 1324, 1333, 1349, 1366, 1373, 1376, 1384, 1388, 1402]. Architectures [532, 533, 535, 1239, 1284, 1290, 1336]. Area [879]. Arithmetic [987, 1393]. array [1384]. Arrowhead [743, 963]. Art [196, 407, 1268]. Arterial [1009]. arts [1410]. artwork [502]. Aspects [142, 167, 168, 714]. Assembler [1080]. assemblers [1344, 1377]. Assemblies [918]. assembly [1296]. Assessment [967]. Assignment [1260]. Associates [898]. Association [1109]. Associative [1102, 1384]. Assurance [1415, 1422]. astronomers [1272]. Atlantic [64, 96]. Attacking [349]. Attend [929]. Author [272]. Authors [8, 27, 47, 73, 116, 141]. Automated [466, 501, 918, 943, 1013, 1231, 1420]. Automatic [1319, 1451]. Automating [556]. Automation [94, 143, 216–218, 232, 361, 394, 396, 500, 502, 503, 554, 913, 1155, 1295, 1296, 1408]. automobile [688]. availability [1292]. Available [160, 188, 226, 280, 307, 359, 384, 412, 1016].

B [344]. Babel [1256]. Back

[385, 405, 406, 432, 433, 452, 473, 495, 570, 606, 668, 671, 702, 726, 751, 771, 805]. Bangers [1056]. Barriers [1391]. Base [347, 1214, 1308, 1354, 1382, 1384, 1385, 1397, 1398]. Based [348, 822, 884, 886, 1181, 1258, 1279, 1377]. Bases [160, 188, 226, 280, 307, 359, 384, 412, 435, 478, 559, 583, 628, 687, 713, 738, 779, 832, 855, 915, 950, 1007, 1031, 1059, 1084, 1104, 1127, 1226, 1310, 1437]. BASIC [1080, 1278, 1334, 1351]. Basis [64]. Batch [267]. Battin [104]. Bayesian [1014]. be [966, 1204]. Beam [850, 1363]. Because [844]. Become [81]. becomes [233]. Behavior [1187–1190]. Behind [229]. Believes [163]. Better [163, 853, 876, 911, 935, 954, 959]. betting [739]. between [1034, 1172]. Bibliography [16, 35, 59, 80, 271, 654, 1110, 1178]. big [568, 576]. biological [1253]. Biomedical [595, 1165]. Bipolar [71, 1264]. Bit [71, 329, 652, 1329, 1333, 1344]. Bit-Oriented [329]. bit-slice [1344]. bit-sliced [652, 1333]. blocks [1385]. Blood [821]. Bloomington [706]. Blowing [121]. Board [53, 165, 323, 482, 688, 757, 1338, 1450]. Book [279, 770, 804, 833, 867, 899, 924, 995, 1091, 1149, 1227, 1439, 1471]. bookshelf [700, 834, 926, 994, 1046, 1093, 1122, 1150]. Brains [1228]. Branch [1403, 1460]. breaker [788]. bridge [1034]. Bridging [1133]. Bring [229]. Brings [201]. Britain [794]. Bubble [1105]. Budgets [257]. Building [854, 1202]. Builds [112]. Bus [1386, 1409]. Business [65, 221, 1181]. Bylaws [9, 123, 647, 648, 1004, 1225]. ByteShop [1202].

Cache [348]. Cache-based [348]. Calculator [568, 569, 576, 577]. Calculators [374, 644, 680]. Calendar [2, 21, 90, 120, 136, 158, 187, 213, 235, 259, 294, 320, 365, 391, 416, 438, 459, 479, 498, 514, 529, 551, 578, 614, 615, 641, 677, 692, 708, 732, 754, 776, 810, 839, 931, 960, 1028, 1067, 1090, 1112, 1146, 1222, 1436, 1469]. Calender [67]. California [835, 868, 1109]. Call [20, 41, 61, 82, 149, 164, 237, 260, 270, 293, 321, 343, 366, 387, 392, 417, 439, 460, 480, 499, 515, 530, 552, 579, 616, 642, 678, 689, 707, 709, 710, 733, 755, 777, 811, 812, 824, 840, 845, 864, 874, 906, 961, 1001, 1025, 1029, 1037, 1072, 1089, 1103, 1113, 1114, 1128, 1147, 1206, 1223, 1438, 1448, 1452, 1457, 1470]. Calls [263]. Can [84, 87, 154, 966]. Canada [789]. Capability [1395]. Capture [1079, 1100]. card [548, 669]. Cardiology [1006]. Cards [819, 865, 895]. Care [396, 818]. carrier [1299]. Case [362, 394, 1359]. cashless [484]. Catch [52, 873]. CDL [783]. CECC [1109]. CEDPA [1109]. cells [1264]. Cellular [1381]. Cellular-Logic [1381]. century [711]. Certification [1417]. CFA [1285]. Chairman [104, 879]. Challenge [143, 1079, 1100]. Challenges [1204]. Change [1, 109, 1134]. Changes [123, 292, 647]. changing [1376]. Channel [843, 880, 1215, 1283, 1429, 1461]. Chapter [145, 613, 879]. Chapters [186, 1460]. Characteristics [49, 1106]. characterization [243, 1168]. characters [1243]. Charge [1160, 1369]. Charge-coupled [1160, 1369]. CHDL [795, 1257]. Check [889]. Chinese [943]. chip [1230, 1299]. chips [1267]. Choose [129]. choosing [652]. Chord [1367]. choreography [1346]. Chronic [822]. Circuit [1402, 1404]. Circuits [63, 219, 715, 888, 1156, 1357, 1411]. City [64, 96]. Classification [1289]. Classified [126, 152, 336, 601, 632, 698, 721, 749, 767, 802, 900, 988, 1069, 1118, 1145]. clinical [393, 1326]. CMOS [1357]. Code [970, 981, 1252]. codes [1186]. coding [585]. Coherent [1370]. Colleges [221]. Combine [65]. Come [64]. Coming [1380]. Command [969]. commercialization [444]. Committee [879]. Committees [99]. common [1312]. Communication [74, 256, 519, 520, 626, 969, 1180, 1249, 1289, 1294]. Communications [4, 139, 189, 325, 327– 329, 517, 725, 929, 956, 1103, 1418, 1455]. comparing [1291]. comparison [1172]. Compcon [245, 263, 388, 487, 548, 655, 670]. Compcon’72 [170, 177, 202, 253, 285]. Compcon73 [308]. Compcon77 [1197]. Competition [486, 1056]. compleat [458]. Component [426, 1321]. composing [558]. composition [1348]. Computation [1173, 1345]. Computer [9, 20, 29, 32, 50, 55, 74, 77, 83, 88, 96, 101, 104, 105, 108, 110, 114, 128, 129, 131, 133, 135, 148, 153, 157, 162, 165, 178, 185, 189, 200, 206, 212, 214, 227, 229, 230, 239, 254, 256, 262, 268, 281, 283, 287, 292, 301, 309, 334, 350, 355, 368, 382, 402, 414, 451, 543, 568, 575, 576, 586, 644, 648, 669, 680, 763, 842, 849, 877, 885, 889, 905, 909, 932, 933, 996, 998, 1000, 1004, 1011, 1023, 1049, 1055, 1103, 1244]. Computer [6, 24, 43, 49, 92, 175, 190, 192, 194, 219, 220, 238, 242–244, 271, 325, 327, 329, 345– 348, 363, 377, 378, 380, 398, 400, 407, 423, 428, 430, 446, 447, 450, 467, 469, 490, 491, 517, 519, 554, 818, 820, 822, 837, 856, 870, 903, 914, 919, 987, 1012, 1017, 1040, 1061, 1074, 1105, 1108, 1132, 1133, 1135, 1137, 1138, 1148, 1165, 1167, 1176, 1178, 1179, 1192, 1196, 1202, 1220, 1225, 1255, 1259, 1288, 1289, 1300, 1304, 1346–1348, 1352, 1355, 1386, 1393, 1394, 1423, 1424, 1458– 1460]. computer [15, 19, 33, 38, 53, 58, 78, 106, 123, 124, 145, 150, 322, 370, 379, 390, 396, 397, 399, 421, 427, 429, 441, 442, 448, 461, 463, 470, 471, 481, 483, 492, 493, 518, 532, 547, 603, 711, 712, 736, 760, 780, 781, 787, 788, 813, 841, 853, 875, 908, 934, 954, 962, 1003, 1032, 1054, 1076, 1098, 1129, 1139, 1160, 1171, 1174, 1180, 1191, 1198, 1232, 1261, 1275, 1282, 1284–1286, 1291– 1293, 1305, 1306, 1311, 1316, 1321, 1334, 1335, 1338, 1340, 1341, 1349, 1359, 1360, 1388, 1407]. Computer [62, 85, 99, 111, 130, 155, 173, 179, 184, 203, 231, 233, 241, 255, 273, 288, 310, 360, 472, 494, 525, 546, 604, 636, 647, 836, 869, 902, 928, 955, 997, 1024, 1048, 1071, 1095, 1124, 1152, 1440, 1460, 1472]. Computer-Aided [55, 77, 105, 554, 919, 1288, 1347].

Computer-Based [822].

Computer-Communications [189].

Computer-Generated [1040]. Computer-Oriented [108, 287]. Computer-to-Computer [329]. computer/communication [1180]. Computerized [395, 817, 821, 1403]. Computers [10, 143, 163, 296, 357, 393, 462, 737, 815, 1006, 1131, 1172, 1210, 1250, 1257, 1287, 1326, 1330, 1383, 1427, 1453, 1456]. Computers* [1014]. Computing [42, 48, 119, 172, 182, 208, 290, 341, 358, 760, 859, 912, 1053, 1109, 1201, 1209, 1241, 1356, 1370, 1371, 1410, 1426]. computors [740]. Concept [1058]. Concepts [287, 964, 965, 1276, 1299, 1381]. Concerned [207, 1428]. Concerning [257]. Concerns [163]. Concluding [973]. Conclusions [986]. Concurrency [1183, 1371]. Concurrent [557, 1399]. Conference [20, 96, 97, 104, 114, 133, 227, 239, 256, 268, 603, 859, 870, 903, 929, 1053, 1109, 1211, 1426, 1437, 1458]. configuration [1368]. Congress [901]. Considerations [70, 71, 408, 535, 652, 1038, 1270, 1361, 1425]. Consortium [1109]. Constitution [9]. constrained [558]. content [1385]. content-addressable [1385]. Contents [3, 22, 40, 66, 89, 109, 113, 134, 159, 183, 205, 234, 258, 291, 317, 340, 367, 386, 411, 437, 455, 476, 496, 512, 527, 545, 573, 611, 639, 673, 705, 729, 753, 774, 808, 838, 871, 904, 930, 957, 999, 1026, 1050, 1073, 1097, 1126, 1194, 1412, 1443]. Contest [32, 619, 645]. Continuing [553, 634, 835, 868]. Continuous [1008]. Control [95, 329, 759, 969, 1185, 1280, 1281, 1286, 1294, 1343, 1375]. Controlled [558, 690, 1083]. controllers [688]. Cookbook [51, 100, 144, 191, 215, 251, 277, 468, 488, 537, 735]. Cooking [28]. Coronary [1009]. Corporation [588]. correcting [1186]. COS [328]. COSINE [376]. Cost [109, 686, 1060, 1074, 1161, 1388]. Councils [81]. coupled [1160, 1369]. Coupon [127].

Course [171, 1137]. Courses [553, 567, 602, 634, 635, 667, 697, 723, 803, 835, 868, 1120]. Cover [312, 339, 383, 385, 403– 406, 409, 410, 432–434, 452, 453, 473– 475, 495, 511, 526, 544, 570–572, 574, 606– 608, 637, 638, 668, 671, 672, 702–704, 726– 728, 751, 752, 771–773, 805–807, 1414, 1445]. Coverage [290]. covering [379, 399, 429, 448, 470, 492]. Credit [819, 865, 895]. Critique [1277, 1326, 1374]. Cross [1080, 1231]. Cross-Assembler [1080]. cross-system [1231]. CRT [691, 1082, 1362]. crunching [1393]. cryptanalysis [1262]. cure [29]. Curicula [1133]. Current [818, 1036, 1237, 1397]. Currently [1016]. curriculum [1232]. Cut [881, 927, 947]. cycle [1388]. cyclic [1158].

Data [4, 5, 7, 26, 160, 188, 226, 266, 280, 307, 327, 347, 359, 384, 412, 415, 435, 478, 519, 520, 559, 583, 621, 628, 660, 687, 713, 738, 779, 832, 855, 881, 915, 927, 947, 950, 956, 1007, 1031, 1059, 1084, 1104, 1107, 1109, 1127, 1214, 1226, 1243, 1262, 1273, 1286, 1290, 1301, 1303, 1308, 1310, 1318, 1342, 1354, 1382, 1384, 1385, 1396– 1398, 1418, 1437, 1455]. Data-base [1354]. Database [501, 1380]. database-driven [501]. databases [457]. DataCenters [816, 847, 887, 917]. Datapac [112]. Day [907]. Days [1047]. DDL [784]. December [1224]. Decision [820]. Dedicated [1008]. Deductive [888]. Definition [1136]. Definitions [964]. Delivery [311]. Departments [376, 1135]. Dept [28]. Derby [4]. Describe [345, 346]. Description [780, 781, 787– 794, 1011, 1256, 1260, 1327, 1375]. descriptions [1285, 1324]. Descriptive [1255]. Design [25, 46, 49, 55, 70, 71, 77, 84, 87, 94, 105, 216– 220, 229, 232, 287, 322, 361, 370, 377, 379, 390, 397, 399, 408, 415, 421, 427, 429, 441, 446, 448, 461, 467, 470, 481, 490, 492, 500, 502, 503, 534, 554, 556, 653, 660, 813, 841, 875, 908, 913, 916, 919, 934, 962, 980, 1003, 1032, 1051, 1054, 1076, 1077, 1079, 1098, 1100, 1129, 1136, 1155, 1170, 1179, 1198, 1204, 1230, 1239, 1247, 1257– 1259, 1270, 1310, 1340, 1359, 1361, 1378, 1425]. design-economic [408]. designer [380, 400, 430, 450, 471, 493]. Designing [1447]. designs [558]. detecting [1186]. Detection [1419]. Development [712, 850, 969, 983, 1193, 1236, 1331, 1332, 1365, 1377]. Developments [218, 500, 1169, 1170, 1191, 1308, 1349]. device [1369]. Devices [142, 167, 168, 1328, 1361, 1381, 1409]. Diagnosis [29, 44, 115, 1235]. Diagnostic [593]. Dial [156]. DIC [1083]. Did [674]. Different [758, 1161]. diffusion [1246]. Digital [44, 46, 115, 298, 396, 443, 501, 557, 582, 585, 587, 588, 591– 593, 595, 682, 883, 884, 886, 1140, 1171, 1172, 1259, 1270, 1271, 1274, 1294, 1358, 1403]. digital/analog [1171]. diode [1264]. Direct [1341, 1456]. Direct-Executing [1456]. direct-execution [1341]. Directed [740, 1363]. Directions [1015, 1349]. Directory [879, 943]. Disciplines [977]. discrimination [1288]. Disk [311]. Diskette [373]. Disks [846, 848, 852]. Display [1276, 1362, 1363, 1366]. displays [683]. Distance [690, 1187]. Distinguished [146]. Distributed [532–535, 983, 1280, 1281, 1301– 1303, 1308, 1343, 1426]. Distributed-function [532–534]. Distribution [1134]. Do [221, 780]. DoD [1312, 1411]. Does [876, 911, 935, 959]. don’t [1111]. Down [881, 927, 947]. Dramatic [68]. Driven [501, 884, 886]. drivers [1319]. Dues [257]. During [285]. Dynamic [1106, 1242, 1294, 1316, 1336].

earth [717]. Easier [876, 911, 935, 959]. Easy [357]. eating [1372]. economic [408]. Ecstacy [975]. editor [675, 676, 913, 1132]. editorial [1199]. Editors [1033]. EDP [207, 854]. Education [156, 553, 634, 835, 868, 1132, 1139, 1148, 1300, 1355]. Educational [1109, 1274]. Effect [194, 1321, 1340]. Effective [266]. Effects [24, 1450]. efficiency [1387]. efficient [1252]. effort [1312]. EG&G [592]. Eighth [1039]. Elected [323, 757]. Electrical [376, 1135, 1234]. Electrocardiograms [1010, 1012]. electrocardiographic [394]. Electron [850]. Electron-Beam-Addressed [850]. Electronic [690, 846, 848, 852, 918, 1158, 1265]. Electronics [1047]. Element [25]. elements [1394]. Eliminate [84, 87]. Else [112]. Emerging [823]. employment [1136]. emulation [1245, 1324]. Emulator [1035]. Encryption [1262]. Encyclopedia [1148]. End [257]. energy [1280]. Engine [979]. Engineer [1211]. Engineering [376, 553, 634, 714, 835, 861, 868, 916, 936, 938, 1133, 1135, 1211, 1232, 1300, 1307, 1322, 1353, 1355]. engineers [1234]. enhancement [1323]. enjoying [109]. enterprise [176]. entertainment [1311]. entry [1243]. Environment [519, 884, 886, 916, 1233]. Equipment [982]. Error [1186, 1419]. Evaluating [267]. Evaluation [43, 264, 265, 347, 1110, 1248, 1284, 1285, 1310]. Events [2, 21]. Ever [52, 240]. Everybody [877, 909, 933, 998, 1023, 1049, 1055]. Everything [284]. Evolution [217, 518, 1328, 1329, 1362, 1442]. Evolved [1379]. Evolving [445, 977]. exchange [1403]. Executing [1456]. execution [1320, 1341]. executive [458]. Exercise [1010]. Exhaustive [1262]. Experience [976, 1016, 1420]. Experienced [972]. Experiences [445]. Experimental [850, 1271, 1282, 1302, 1366]. Expert [350]. Exploiting [1371]. Exploring [1251]. Express [1134]. Expresses [163]. Extending [1306]. extension [1254]. Extraction [13]. eye [1282].

F8 [1277]. Faced [1204]. faces [462]. facilities [626, 760, 1309]. facility [1166, 1271, 1274, 1280]. Factory [939]. failure [1279]. Fall [655]. Family [1060, 1286, 1287, 1313, 1374, 1386]. Fast [93, 1456]. Fault [42, 43, 46, 48, 115, 119, 182, 208, 341, 912, 1189, 1235]. Fault-Tolerant [42, 46, 48, 182, 208, 341, 912]. Feature [13, 659, 742, 854, 943, 1154, 1157, 1350, 1425, 1456, 1459]. February [263]. Federal [791, 1134]. Fiction [717]. Field [172]. Fifth [891, 1437]. File [91, 1185]. Films [844]. Financial [1052]. Find [154]. firmware [1322, 1323]. First [64, 176, 1193, 1207]. Five [868]. FJCC [31]. Flank [349]. Flation [897]. flow [1342]. Flying [1407, 1424]. Forever [1134]. form [655, 668]. Formal [980, 1285, 1324, 1441]. Formerly [83, 110]. FORTRAN [505, 658, 966]. Fountains [84, 87]. Four [835, 1133]. Four-Year [1133]. Fourier [93]. Fourth [956, 1053, 1393]. France [790]. Francisco [229, 263]. Fred [140]. Freedom [143]. Freer [176]. Frequency [1189]. Front [312, 339, 383, 403, 404, 409, 410, 434, 453, 474, 475, 511, 526, 544, 571, 572, 607, 608, 637, 638, 640, 672, 703, 704, 727, 728, 730, 752, 772, 773, 806, 807]. FSD [971]. full [379, 399, 429, 448, 470, 492]. Fun [1202]. function [532–534]. Functions [265]. Future [5, 207, 425, 650, 916, 985, 1173, 1180, 1199].

Galindo [118]. gambling [736]. Garnet [844]. Gas [821, 1280]. gas-laser [1280]. GASP [1245]. GASP-PL [1245]. GASP-PL/I [1245]. Generalized [220]. Generated [1040]. Generation [1174, 1231, 1252, 1258]. Generations [195]. Geometric [1298]. Germany [791, 795]. Get [290, 674]. GGG [844]. Gift [284]. giveaway [568, 569, 576, 577]. Glossary [1339]. Go [52, 1330]. Goals [161, 938]. going [644, 680]. Good [844]. Got [93]. Governing [165, 323, 482, 757]. Governors [53]. Grade [83, 110]. Grand [224, 282]. Graphic [558, 1361]. Graphics [503, 742, 1159, 1175– 1178, 1259, 1305, 1311, 1363, 1407, 1424, 1473]. graphs [740]. Grass [273]. Great [4, 32, 794, 877, 909, 933, 998, 1023, 1049, 1055, 1220]. Group [1, 15, 33, 83, 110]. Grow [844]. Guest [913, 1132, 1199]. guide [224, 282, 372].

hand [1243, 1282]. hand-printed [1243]. hand/eye [1282]. handbook [550]. handicapped [1372]. Handle [64]. Happening [952]. Hard [129]. Hardware [244, 462, 536, 780, 781, 787– 794, 1011, 1034, 1082, 1137, 1205, 1208, 1255, 1256, 1260, 1323, 1327, 1363, 1392, 1446]. hardware/firmware/software [1323]. Hardware/Software [1082]. Hayne [488, 735]. Haynes [468]. Health [818]. heart [388]. Heaven [101]. held [795]. Hell [101]. Help [29, 221, 820, 1318]. Helps [221]. Here [996]. Heroes [166]. heterogeneous [533]. Hewlett [940]. Hewlett-Packard [940]. Hierarchically [1140]. hierarchies [1364, 1396]. Hierarchy [1161]. High [7, 1038, 1060, 1156, 1280, 1321, 1358, 1407, 1424, 1456]. high-energy [1280]. High-level [1456]. High-Speed [7, 1321, 1358]. Highlights [762, 963]. Highly [1193]. Hints [1318]. History [1209, 1345]. hobby [1241]. Holograms [1040]. home [686]. home/school [686]. homogeneous [533]. Honeymoon [1151]. Honeywell [1302]. host [1442]. host-to-host [1442]. Hotels [64]. HP [568, 569, 576, 577, 644, 680, 816, 847, 887, 917]. Hybrid [1171–1174].

I/O [464, 1406]. IBM [373, 621, 759, 971]. IBM-FSD [971]. IC [181, 563, 597, 629, 662, 694, 719, 746, 797, 826, 858, 892, 921, 948, 991, 1019, 1042, 1064, 1086, 1116, 1142, 1218, 1432, 1464]. identical [557].

IEEE [9, 20, 32, 74, 83, 88, 110, 114, 128, 131, 133, 135, 153, 157, 165, 178, 185, 200, 204, 206, 230, 241, 254, 281, 283, 290, 309, 334, 355, 368, 382, 402, 414, 568, 576, 644, 648, 680, 842, 877, 897, 905, 909, 932, 933, 998, 1000, 1004, 1023, 1049, 1055, 1196, 1210, 1225, 1423, 1460]. IEEE-Interest [290]. If [84, 87, 129, 154, 240, 853, 954, 1204]. IFIP [942]. II [595, 763, 967, 1424]. III [10, 659, 974]. Illness [822]. Image [582, 587, 588, 591– 593, 595, 742, 1013, 1178, 1270, 1272–1275]. images [1165]. Impact [68, 466, 823, 852, 977, 1057, 1058, 1392]. Implementation [970, 1337]. implemented [521]. Implementing [1038, 1309]. Implications [373, 484, 1265, 1427]. Important [241, 647]. Improve [1417]. Improved [978]. Improvement [971, 1057, 1190]. improving [1292]. Inc. [81, 898]. Index [272, 306, 335, 381, 401, 431, 449, 666, 769, 801, 831, 1224]. Indiana [706]. Individuality [143]. Industrial [138, 1266, 1375]. Industry [247, 486, 1311]. Information [162, 190, 212, 238, 262, 314, 322, 326, 337, 370, 390, 421, 441, 461, 481, 536, 813, 841, 875, 908, 934, 962, 1003, 1032, 1054, 1076, 1098, 1102, 1129, 1181, 1198, 1385]. initiative [1411]. Ink [84, 87]. Innovations [533]. Input [1185, 1238, 1361]. input/output [1185, 1238]. Inspection [918]. Instruction [346, 1388]. instruction-set [1388]. Instrumentation [242, 246, 465, 1279, 1317]. Integrated [219, 715, 1247, 1357, 1406, 1411]. Intel [1329, 1374, 1406]. intelligent [653, 1159]. intensive [396]. interaction [463]. Interactive [465, 503, 558, 742, 1162, 1178, 1259, 1271, 1272, 1305, 1407, 1424, 1473]. Intercon [897]. Interconnection [555, 1405]. Interest [290]. interface [1409]. interfacing [1234]. International [48, 97, 114, 119, 182, 208, 256, 341, 859, 891, 912, 929, 1011, 1211, 1426, 1437, 1458]. Interpretation [821]. Interprocess [626].

Interprocessor [1249]. Interviews [285]. Introducing [781–786, 1074]. Introduction [297, 649, 913, 987, 1132, 1169, 1250, 1255, 1327, 1378]. Invited [81]. IPAD [916]. ISP [346, 785]. Issue [1210]. issues [408, 1237]. Italy [792]. itself [853, 954]. IV [977].

Japan [227, 239, 268, 793, 870, 903, 1094, 1125, 1400, 1401]. JCCs [161, 257]. Join [185, 206, 1220]. Joint [96]. JOLT [1074]. Journal [63].

Keepsake [240]. Key [1191]. keyboard [1367]. Keynote [1399]. kill [29]. Kit [1212]. Know [547, 844, 914, 952]. Knows [284].

Labolink [1293]. Laboratory [298, 591, 593, 595, 742, 1140, 1233, 1253, 1293]. Lake [743, 963]. Landmarks [1107]. Language [978, 984, 1038, 1258, 1259, 1312, 1327, 1347, 1456]. Languages [780, 781, 787– 794, 1011, 1255, 1256, 1260, 1313, 1377]. Large [196, 219, 362, 443, 739, 1385, 1437]. Large-Scale [219]. laser [1280]. Last [64, 819, 865, 895]. Latin [1139]. law [483]. Lawrence [593, 595]. layered [1323]. Learning [1459]. Leave [204, 229]. left [388]. Legal [408, 484, 485]. Let [52]. Letters [11, 75, 102, 211, 236, 261, 295, 369, 422, 675, 1033]. Level [976, 1038, 1188, 1456]. Levels [1161]. Life [163, 1388]. Life-cycle [1388]. Limitations [1335, 1395]. Limits [154]. Line [763, 1010, 1057, 1243]. linked [1293]. List [318, 510]. literary [1347]. Literature [1217, 1431, 1463]. lithography [1156]. Livermore [593, 595]. LLL [1278]. LM2 [1099]. Local [285]. LOGAL [1257]. Logic [142, 167, 683, 711, 712, 714, 873, 884, 886, 888, 891, 1099, 1170, 1185, 1257, 1381, 1447]. Logical [168]. Logo [32]. Look [853, 954, 1199]. Lookahead [759]. Looking [651, 1005]. loom [558]. lost [819, 865, 895]. Low [686, 1060]. low-cost [686]. Lowest [1074]. LSI [169, 372, 504, 654, 1169, 1335, 1337, 1340, 1341, 1393].

M [1314]. Machine [625, 981, 1015, 1099, 1252, 1321, 1375, 1384]. Machine-Pattern [1015]. Machines [46, 1157, 1380, 1382, 1385, 1405]. Macro [1377]. Macro-assemblers [1377]. macro-based [1377]. Made [357]. magazine [233, 546]. Magnetic [1105]. Main [424]. Mainframe [1337, 1425]. Maintenance [91]. Major [163]. Make [163, 1136]. Making [820]. Man [284]. Management [466, 916, 1193, 1214, 1303, 1308, 1368, 1398]. Management/Engineering/Design [916]. Managing [822, 1331]. MarAd [919]. March [181]. marketplace [380, 400, 430, 450, 471, 493]. marriage [1281]. Mass [425, 744, 1309]. Mathematical [741, 1014]. matter [640, 730]. MAXC [1325]. May [96, 1204]. Mayer [224, 282]. McAuto [968]. MCS [1374]. MCS-48 [1374]. Mean [876, 911, 935, 959]. measurability [1167]. Measurement [242, 243, 1284]. Measurements [505, 682]. measures [1387]. Measuring [971]. Mechanical [1298]. Media [1242]. Medical [395, 820]. Medicine [393, 815, 1326]. meets [1268]. Member [81]. Members [53, 165, 323, 482, 568, 569, 576, 577, 644, 680, 757, 1460]. Membership [83, 110, 281, 368, 414, 620, 646, 877, 909, 933, 998, 1023, 1049, 1055]. Memories [50, 72, 1105, 1158, 1160, 1265, 1369]. Memory [23–25, 70, 71, 169, 269, 286, 424, 426, 466, 850, 1161, 1179, 1184, 1186, 1254, 1364, 1396, 1449]. Merge [204]. Message [39, 362, 676, 1444]. message-switching [362]. Messages [1413]. meter [690]. Method [1081, 1172]. Methods [26, 108, 976, 1240, 1291, 1292, 1345]. Metrics [1421]. Micro [762, 1213, 1279, 1281, 1458, 1467]. Micro-12 [1467]. MICRO-9 [1213]. micro-based [1279]. micro-mini [1281]. Microcomputer [538, 901, 1051, 1080, 1081, 1212, 1234, 1236, 1239, 1249, 1278, 1280, 1332, 1372, 1374, 1401]. Microcomputer-aided [1372]. Microcomputers [654, 1231, 1232]. Microdare [1456]. Microelectronics [1450]. Microfiche [223]. microprocessing [504]. Microprocessor [649, 653, 685, 686, 690, 1060, 1077, 1078, 1159, 1169, 1229, 1230, 1233, 1235, 1238, 1245, 1253, 1269, 1277, 1328, 1329, 1331, 1333, 1373, 1376, 1377, 1425]. Microprocessors [372, 650, 654, 688, 691, 1082, 1083, 1237, 1313, 1344, 1400]. Microprocessors-Controlled [1083]. Microprogram [1344]. microprogramable [652]. microprogrammable [1157, 1286]. Microprogramming [269, 361, 907, 1034, 1036, 1038, 1039, 1154]. Microsystems [1433, 1465]. Might [229]. migration [1323]. MIL [1287]. MIL-SPEC [1287]. military [1386, 1388]. Million [819, 865, 895]. mind [1230]. Mini [374, 816, 847, 887, 917, 1281, 1458]. Minicomputer [139, 373, 375, 742, 940, 1008, 1099, 1166, 1181]. Minicomputer-based [1181]. MiniComputers [72, 137, 138, 180, 210, 298, 371, 1254]. Minimal [1263]. minimize [1179]. Miscellanea [12, 30, 54, 76, 103, 122, 147]. MITS [937]. MOBIDIC [344]. MOBIDIC-B [344]. Model [759, 1133, 1246]. Modeling [1294, 1297, 1298, 1360]. Models [534, 941, 1187, 1408]. Modem [7]. Modern [837]. Modular [214, 301, 442, 465, 1314, 1338, 1340, 1341]. Module [850, 976]. Module-Level [976]. modules [443, 444, 1335, 1337]. Money [952]. Monitoring [117, 244, 817, 1008].

Monitors [1108]. monograph [1333]. monolithic [1358]. monthly [233]. MOS [109, 502, 1060]. Moses [140]. MOSFET [1156]. Motorola [1376]. MPU [1203]. MSFC [1273]. Much [221, 1202]. multi [714, 1249]. multi-microcomputer [1249]. multi-valued [714]. multiminiprocessor [521]. multiphasic [395]. Multiple [711, 712, 891, 1078]. Multiple-Valued [711, 712, 891]. multipliers [1358]. Multiprocessing [193]. multipurpose [668]. multistable [715]. MUMS [1229]. music [1348]. my [388].

Name [93, 1211]. Named [104]. Names [176]. nanoprogramming [1154]. National [172]. NBS [1262]. NCC [621, 814, 851, 878, 910, 1107, 1427]. Near [194]. nearly [557]. Need [166, 780, 978, 984, 1307]. Needs [207, 889, 985]. Netherlands [117]. Network [117, 362, 408, 518, 626, 1290, 1292, 1293]. networking [407]. Networks [6, 44, 92, 189, 415, 517, 547, 557, 586, 660, 716, 837, 914, 1269, 1289, 1294, 1304, 1338, 1342, 1404, 1405, 1447]. Never [52]. News [12, 30, 54, 76, 103, 122, 147, 1096, 1123, 1153]. Newsletter [55, 77, 105]. Next [195, 1047, 1131, 1174, 1200, 1212]. Next-generation [1174]. Ninth [232]. No [209]. No-Tea-Party [209]. Non [142, 167, 408, 1234]. non-electrical [1234]. Non-Logic [142, 167]. Non-technical [408]. Nonimpact [1062]. Notation [345, 346]. Notes [1096, 1123, 1153]. Notices [12, 30, 54, 76, 103, 122, 147]. NSF [1334]. Nuclear [1416].

O [464, 1406]. observations [443]. oceans [1350]. October [1114]. off [739, 1337]. off-the-shelf [1337]. off-track [739]. Offers [835, 868]. office [1366]. Officers [53, 165, 323, 482, 757]. ol [585]. on-board [688, 1338]. On-Line [763, 1010, 1243]. One [229, 907, 1070, 1202]. One-Day [907]. only [568, 569, 576, 577]. Open [843, 880, 1215, 1283, 1429, 1461]. Operating [171, 300, 463, 464, 623, 626, 1182, 1183, 1214, 1304, 1309, 1342]. Opportunity [143, 1136]. Optical [859, 1370]. optically [1293]. order [668]. Ordinary [505]. Organization [24, 26, 1158, 1321, 1341, 1352]. Oriented [98, 108, 287, 328, 329]. Other [408, 1211]. Our [154, 844, 1228]. Output [1040, 1185, 1238]. Outstanding [835, 868]. Overview [42, 1241, 1261, 1302, 1322, 1363, 1387, 1389].

package [1299]. Packaging [856]. Packard [940]. Page [758, 788, 1062, 1189]. Pages [1070]. Papers [20, 41, 61, 82, 149, 164, 237, 260, 263, 270, 293, 321, 343, 366, 379, 387, 392, 399, 417, 429, 439, 448, 460, 470, 480, 492, 499, 515, 530, 552, 579, 616, 642, 678, 689, 707, 709, 710, 733, 755, 777, 811, 812, 824, 840, 845, 864, 874, 906, 961, 1001, 1025, 1029, 1037, 1072, 1089, 1103, 1113, 1114, 1128, 1147, 1206, 1223, 1438, 1448, 1452, 1457, 1470]. Parallel [1102, 1275]. Parallelism [505]. Part [593, 595, 1424]. Parts [1298]. Partsbank [1134]. Party [209]. Pascal [1389, 1390]. Passing [344]. Past [425]. Patent [1410]. Path [1419]. Patient [817]. Patients [822]. Pattern [13, 91, 97, 108, 160, 188, 226, 280, 307, 359, 384, 412, 435, 457, 478, 559, 583, 594, 628, 687, 713, 738, 742, 779, 832, 855, 915, 950, 1007, 1015, 1031, 1059, 1084, 1104, 1127, 1162, 1166, 1226]. Paynes [537]. PBX [1403]. People [952]. Performance [242–244, 264– 266, 271, 347, 534, 758, 1060, 1110, 1161, 1323]. peripheral [1328]. peripherals [486]. Personal [1201, 1241, 1242, 1330, 1356]. perspective [1369]. Petition [1, 118]. philosophy [1239]. Physical [168]. Physiologic [1008]. Pick [1228]. Picture [585, 622, 674, 1163, 1166]. pipelining [521]. Pit [1133]. PL [1314]. PL/I [1245]. PL/M [1314]. Place [815]. Plan [929]. Plans [1422]. Pleasure [65]. Plot [1070]. PLZ [1313]. PMS [345, 786]. POD [1348]. Point [228, 252, 278, 354, 420, 653, 1353]. point-of-sale [653]. poised [1305]. policy [717]. Pollution [117]. Polyphonic [1348]. Portable [328, 1080]. Power [169]. Practical [686, 965]. practice [1170]. practicing [1318]. Pre [86]. Pre-register [86]. prefetching [1364]. Preliminary [1467]. Preparing [985]. Present [207, 425, 650, 1180]. presented [621]. Presents [897]. President [39, 289, 316, 364, 418, 440, 477, 497, 513, 528, 612, 613, 731, 775, 809, 872, 958, 1027, 1075, 1195]. Preview [814, 851, 878, 910]. Principles [171, 938, 1342]. printed [1243]. Printer [1056]. Printing [1057, 1058, 1061, 1062]. Printout [84, 87, 1070]. Privacy [300, 1428]. probes [658]. Problem [52, 229, 555, 1168, 1261]. Problems [974, 1009, 1336, 1352]. Procedure [44, 1417]. Procedures [329]. Process [10, 558, 938, 1281]. Processes [1298, 1343]. Processing [267, 327, 582, 588, 591– 593, 595, 742, 1012, 1102, 1107, 1109, 1165, 1166, 1178, 1270, 1273–1275, 1301, 1338, 1358]. Processor [1078, 1102, 1302, 1366, 1406]. Processors [1384, 1385, 1425]. Product [305, 333, 356, 380, 400, 430, 450, 471, 493, 509, 542, 566, 600, 633, 665, 699, 724, 750, 768, 800, 830, 862, 896, 923, 951, 989, 1022, 1045, 1068, 1092, 1119]. Products [14, 34, 57, 79, 107, 125, 151, 174, 198, 225, 249, 276, 303, 331, 352, 377, 397, 427, 446, 467, 490, 507, 522, 539, 562, 596, 627, 661, 693, 718, 745, 764, 796, 825, 857, 890, 920, 946, 990, 1018, 1041, 1063, 1085, 1115, 1141, 1216, 1430, 1462]. Profession [292]. Professional [162, 190, 212, 238, 262, 322, 370, 377, 390, 397, 421, 427, 441, 446, 461, 467, 481, 490, 813, 841, 875, 908, 934, 962, 1003, 1032, 1054, 1076, 1098, 1129, 1198]. Professionalism [161]. Profile [305, 333, 356, 380, 400, 430, 450, 471, 493, 509, 542, 566, 600, 633, 665, 699, 724, 750, 768, 800, 830, 862, 896, 923, 951, 989, 1022, 1045, 1068, 1092, 1119]. Prognosis [1255, 1327]. Program [31, 146, 298, 299, 456, 487, 590, 657, 684, 1002, 1016, 1030, 1187–1190, 1268, 1315, 1317, 1320, 1364, 1371, 1416, 1467]. programmable [1295]. Programmed [889]. Programmer [140, 1318]. Programming [743, 964, 965, 967– 969, 971, 972, 974, 975, 977, 978, 982, 1312– 1314, 1399]. Programming* [1081]. Programs [196, 505, 658, 1106, 1133, 1247, 1285]. Progress [590, 657, 1297, 1321]. Project [464, 1212]. projects [1331]. Proposed [486, 1081, 1421]. proprietary [485]. Protection [300, 485, 624, 1410, 1416]. protocol [1441, 1442]. Proudly [897]. Proven [1057]. Provisional [9]. Public [163, 1193, 1428]. publication [741]. Publications [128, 153, 178, 186, 200, 230, 241, 254, 281, 283, 309, 334, 355, 382, 402, 451, 472, 494, 525, 543, 604, 636, 942]. Pulmonary [1246]. Pulse [873]. purse [761].

Quality [1393, 1415, 1417, 1421, 1422]. Quantitative [967, 1172]. Quantitative/method [1172]. Questionnaire [56].

R [1398]. Race [877, 909, 933, 998, 1023, 1049, 1055]. racetrack [737]. radiography [593]. radix [715]. RAM [1263, 1267]. Ramdom [70]. Rams [69]. Random [71]. range [379, 399, 429, 448, 470, 492]. Reach [550]. Reactor [1416]. reader [669]. Readers [223]. readership [546]. ready [1204]. Real [244, 736, 945, 970, 1136, 1278, 1358]. Real-Time [244, 970, 1278, 1358]. Realtime [1079, 1100]. Reasons [64, 86]. Recently [1204]. Recognition [13, 97, 108, 160, 188, 226, 280, 307, 359, 384, 412, 435, 457, 478, 559, 583, 594, 628, 687, 713, 738, 742, 779, 832, 855, 915, 950, 1007, 1031, 1059, 1084, 1104, 1127, 1162–1164, 1166, 1226]. Reconfigurable [1229, 1338]. Recording [881, 927, 947]. records [395]. Recovery [941]. Reduction [169]. Redundancy [987]. references [575]. referendum [419]. register [86, 444, 1306]. register-transfer [1306]. Registration [655]. Related [712, 977]. Relational [1382, 1398]. Relays [1203]. Reliability [45, 1186, 1266, 1292, 1357, 1392]. reliable [1240]. Remarks [973]. Remote [267, 1279]. Repeat [1193]. Replacement [758, 1189]. Report [69, 548, 590, 613, 657, 744, 762, 823, 856, 1052, 1211, 1213]. Reporting [821]. Repository [15, 33, 58, 78, 106, 124, 150, 173, 197, 222, 248, 274, 302, 330, 351, 379, 399, 429, 448, 470, 492, 506, 524, 541, 565, 599, 631, 664, 696, 722, 748, 766, 799, 829, 863, 894, 925, 953, 993, 1021, 1044, 1066, 1088, 1121, 1144, 1221]. Representation [1009]. Representative [10]. Republic [791]. Requirements [83, 110, 817, 1177, 1179]. Research [625, 898, 1148, 1166, 1282, 1296, 1334, 1434, 1466]. Reservations [64]. Resource [172, 193]. Responding [292]. restoration [587]. retrieval [1383]. Review [115, 279, 407, 740, 965, 1091, 1164, 1289]. Reviews [290, 358, 833, 867, 899, 924, 995, 1149, 1227, 1439, 1471]. Revolution [1450]. Ribbon [1056]. Ribbons [84, 87]. Rickey [1212]. Right [129, 550, 1101]. Rising [52]. Robots [1408]. Rock [856]. Role [987, 1384]. Rollback [941]. Rolm [1403]. ROM [109, 1264]. ROMs [112]. Roots [273]. Roster [842, 905, 932, 1000, 1196]. Roundup [1220]. rules [619, 645].

S100 [1409]. Sale [184, 653]. Sampler [1404]. sampling [558]. San [229, 263]. satellite [1175]. Save [84, 87]. saving [109]. Scale [219, 739]. scene [1163]. Scheduling [1182, 1365]. Schemes [1289]. Scholarships [1460]. school [686]. Schottky [1264]. SCi [81]. Science [603, 711, 712, 760, 889, 898, 1133, 1135, 1212, 1300, 1334, 1355]. Scientific [591]. Scorched [717]. SDK [1212]. SDK-8O [1212]. searching [1263]. Second [1017, 1060]. secretary [458]. Security [519, 621, 1261]. See [154, 1111]. seismic [1288]. Selection [13, 243, 1318]. Selections [554]. Self [1453]. Self-Testing [1453]. Sells [112]. Semantic [1252]. Semiconductor [69, 70, 72, 286, 443, 1186]. sensors [1295]. Sequence [1258]. Sequential [91, 1364, 1396]. serial [1265]. Series [1193, 1286]. serve [853, 954]. service [669]. services [377, 397, 427, 446, 467, 490, 1291]. Session [964, 967, 974, 977]. Set [7, 1282, 1388]. set-up [1282]. Sets [346]. Seventies [50]. severely [1372]. Shaded [1311]. shape [1297]. Shared [193, 246]. Sharing [265]. shelf [770, 804, 1337]. shift [716]. Ship [919]. Shopping [1202]. Short [553, 567, 602, 634, 635, 667, 697, 723, 803, 835, 868, 1120]. Should [204, 966]. Shoulders [854]. signal [1358]. Silicon [68]. Silk [761]. SIMD [1405]. Simple [1282]. Simplified [538, 1081]. Simulating [888]. Simulation [81, 534, 557, 883, 884, 886, 1244, 1245, 1247, 1248, 1259, 1359]. Simulations [1459]. Sixth [1418, 1455]. Size [881, 927, 947, 1161]. sizes [758]. SJCC [86, 201]. slice [1344]. sliced [652, 1333]. Small [1181, 1456]. Snap [1051]. Social [408, 1427]. Society [1, 9, 20, 32, 53, 58, 74, 78, 83, 88, 99, 106, 114, 123, 124, 128, 131, 133, 135, 145, 150, 153, 157, 161, 165, 173, 178, 185, 200, 206, 230, 241, 254, 273, 281, 283, 309, 314, 323, 326, 334, 337, 355, 368, 382, 402, 414, 451, 472, 482, 484, 494, 525, 543, 568, 569, 576, 577, 604, 636, 644, 647, 648, 680, 842, 877, 905, 909, 932, 933, 998, 1000, 1004, 1023, 1049, 1055, 1196, 1225, 1423, 1460]. Soft [1366]. Software [45, 462, 466, 485, 535, 556, 658, 741, 861, 936, 938, 939, 969, 980, 983, 1034, 1079, 1082, 1100, 1136, 1176, 1193, 1204, 1211, 1231, 1240, 1248, 1307, 1316, 1317, 1319, 1323, 1331, 1332, 1339, 1353, 1365, 1368, 1377, 1379, 1415, 1417, 1421, 1422]. Solid [63]. Solid-State [63]. solutions [1336]. Some [24, 138, 161, 443, 965, 1425]. sorting [1263]. SOS [1200]. Southern [835, 868]. Soviet [46]. Space [184, 1350]. spacecraft [1338]. SPEC [1287]. Special [613, 659, 742, 854, 943, 1154, 1157, 1158, 1210, 1333, 1350, 1413, 1425, 1444, 1456, 1459]. Specification [976, 1409, 1441]. Speech [1164]. Speed [7, 1156, 1172, 1321, 1358]. SPG [266]. SPIE [651, 1005]. Split [856]. SPLM [297]. Spoken [996]. Spring [96, 487, 548, 1197]. Sr. [140]. Srtre [1012]. Stack [981, 1250–1252]. Staff [318]. Standard [91, 92, 1262, 1390, 1409, 1422]. standardization [1388]. Standards [95, 520, 1421]. Standing [854]. Staran [1275]. State [63, 407, 1375]. Statement [1260]. Static [1106, 1316, 1419]. Statistical [1291, 1365]. Statistics [1014]. Status [1, 69, 818, 1180, 1400]. statute [486]. Stay [828]. Step [1101]. stolen [819, 865, 895]. Stop [240]. Storage [423, 425, 744, 1263, 1309, 1362, 1395]. Strategies [362, 941]. Strengthens [145]. string [1187]. structural [1385]. Structured [743, 964– 970, 972, 974, 975, 977, 981, 982, 1140]. Structures [194, 345, 624, 1252, 1335]. Structuring [979]. Student [1460]. studies [1387]. Study [362, 621, 1137, 1246, 1347, 1359]. studying [1350]. Style [1451]. styles [1345]. Subjects [290]. Substrates [844]. Subsystem [25]. Success [854]. Sue [464]. Summary [46, 463, 504]. Summer [65, 553, 634]. Supercomputer [489]. supercomputers [561]. supervision [464]. support [1334]. surgical [396]. Survey [43, 376, 546, 585, 587, 625, 787, 788, 941, 1036, 1135, 1139, 1191, 1240, 1292, 1308, 1421]. switching [362, 1402, 1404]. Symbolic [220, 1188, 1320]. Symposium [13, 48, 119, 182, 189, 208, 286, 341, 891, 912, 956, 1011, 1017, 1138, 1393, 1418, 1455]. Synthesis [1040, 1257]. System [24, 242, 244, 265, 267, 328, 361, 379, 380, 399, 400, 429, 430, 445, 448, 450, 464, 465, 470, 471, 492, 493, 521, 558, 686, 759, 820, 821, 852, 916, 970, 1038, 1061, 1062, 1074, 1105, 1108, 1110, 1167, 1174, 1181, 1231, 1243, 1273, 1280, 1301, 1303, 1313, 1348, 1352, 1359, 1361, 1366, 1386, 1398, 1456]. System* [1083]. Systems [70, 98, 115, 171, 214, 246, 287, 300, 301, 311, 322, 329, 348, 370, 373, 377, 378, 390, 397, 398, 421, 426–428, 441, 442, 446, 447, 461, 463, 467, 469, 481, 490, 491, 623, 626, 714, 739, 813, 817, 841, 875, 883, 908, 934, 962, 983, 1003, 1032, 1053, 1054, 1076, 1078, 1098, 1129, 1140, 1160, 1162, 1171, 1175, 1176, 1182, 1183, 1198, 1211, 1214, 1235, 1236, 1249, 1259, 1270, 1304, 1308, 1309, 1323, 1325, 1328, 1342, 1354, 1360, 1371, 1388, 1397, 1416, 1426, 1450].

Table [3, 22, 40, 66, 89, 113, 134, 159, 183, 205, 234, 258, 291, 317, 340, 808, 838, 871, 884, 886, 904, 930, 957, 999, 1026, 1050, 1073, 1097, 1126, 1194, 1412, 1443]. Table-Driven [884, 886]. Tackling [1212]. takeoff [1305]. Tar [1133]. Tea [209]. teach [1306]. Teaching [296, 1140, 1234]. Technical [31, 99, 156, 379, 399, 408, 429, 448, 470, 492, 676, 684, 761]. Techniques [169, 220, 271, 888, 970, 974, 1162, 1178, 1182, 1254, 1263, 1339, 1375, 1416, 1441, 1442, 1449]. Technologies [426, 977]. Technology [24, 68, 194, 407, 423, 424, 443, 823, 945, 1016, 1056, 1057, 1180, 1191, 1205, 1276, 1306, 1340, 1351, 1391, 1392, 1395, 1425, 1446]. Telecommunications [518]. Teleconferencing [1473]. Teledisplay [95]. Telephone [943]. Teletraffic [189]. Term [194]. Terminal [98, 1082, 1159, 1177]. Terminal-Oriented [98]. Terminals [653, 691, 1362]. Test [44, 286, 1083, 1258, 1285, 1318, 1319, 1446, 1450]. Testability [1447]. testable [1170]. Testing [395, 501, 658, 982, 1016, 1266, 1268, 1315– 1317, 1320, 1416, 1417, 1419– 1421, 1447, 1449, 1451, 1453]. tests [1267]. Texas [1053]. Text [1383]. Texts [889]. Their [1011, 1255]. theory [1170, 1268, 1340]. There [86, 547, 914]. these [1204]. Things [651, 828]. Think [129, 1134]. Third [1138]. Thirty [1070]. Thoughts [161, 1201, 1209]. Three [195, 1387]. Time [244, 246, 265, 736, 884, 886, 970, 1136, 1278, 1358]. Time-Based [884, 886]. Time-Shared [246]. Time-Sharing [265]. timing [682]. Together [201]. Tolerant [42, 43, 46, 48, 119, 182, 208, 341, 912]. Tool [979, 1359]. Tools [1014, 1016, 1339]. Top [828]. TORC [659]. Total [1261]. Tough [1204]. Tower [1256]. TP [659]. track [739]. Tradeoffs [1082, 1161]. Traffic [5]. Transactions [1210]. transfer [444, 1306]. Transform [93]. Transition [979]. Translation [1038]. Trees [1009]. Trends [23, 25, 26, 49, 180, 210, 375, 586, 1180, 1335, 1362, 1397]. triggered [683]. Troubleshooters [181]. turbulence [518]. turning [1353]. Tutorial [489, 555, 907, 1078, 1105, 1108, 1158, 1164, 1175, 1244, 1263, 1316, 1368, 1423, 1454]. Two [241, 462].

U.S. [1135]. U.S.A. [787]. Ugly [127]. Un-Flation [897]. Undergraduate [171]. Unified [1332]. Uniform [716]. unique [1402]. University [706, 835, 868, 1233]. Update [162, 190, 212, 238, 247, 262, 322, 342, 370, 390, 421, 441, 461, 481, 516, 531, 549, 580, 617, 643, 681, 734, 756, 778, 813, 841, 875, 908, 934, 962, 1003, 1032, 1054, 1076, 1098, 1129, 1130, 1198, 1435, 1468]. USA [227, 239, 268, 870, 903]. USA-Japan [903]. Usage [266, 269]. Use [266, 658, 668, 688, 1248, 1266, 1295, 1335]. user [1230]. Using [240, 372, 1156, 1159, 1245, 1278, 1324, 1419]. Utility [265]. utilizing [1264].

Valued [711, 712, 714, 891]. Variable [715]. Variables [1008]. Vector [561]. verification [1247, 1441]. verify [1324]. Vertical [1323]. Very [1437]. via [981, 1285]. Vice [613]. Video [1013]. Video-Angiocardiographic [1013]. View [228, 252, 278, 354, 420]. Virtual [625, 1157, 1184]. vision [363, 1282]. Visitors [146]. Voices [1256]. Voideband [7]. Volume [5]. vs [462, 1154].

Waiting [311]. Washington [1130]. Waves [1136]. Way [651, 1005, 1134]. Weaving [761]. Week79 [1423, 1454]. Welcome [679]. were [819, 865, 895]. Wescon [679]. We’ve [881, 889, 927, 947]. while [109]. Whistle [121]. Whither [95]. Who [64, 284]. wife [229]. Will [29, 853, 954, 1134]. win [644, 680]. Wins [877, 909, 933, 998, 1023, 1049, 1055]. Word [1366]. Work [378, 398, 428, 447, 469, 491, 1036, 1330]. Workload [243, 1168]. Workshop [74, 94, 214, 232, 301, 463, 504, 554, 743, 744, 762, 763, 795, 823, 856, 963, 1039, 1213, 1214]. World [112, 945, 1074, 1148, 1376]. Worst [229]. Worth [1070]. Wrap [1427]. Wrap-up [1427]. Writing [196, 761].

Xerox [1061]. XPL [1106].

Year [819, 865, 895, 1133, 1205]. Years [1131, 1200, 1207]. you’re [350].

References

Status:1970:GPC

[1] Anonymous. On the Group’s petition for a change to Society status. Computer, 3(5):1, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:CEa

[2] Anonymous. Calendar of events 1970. Computer, 3(5):2–49, September/October 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:TCa

[3] Anonymous. Table of contents. Computer, 3(5):3, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:GDC

[4] Anonymous. The great data communications derby. Computer, 3(5): 5, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hough:1970:FDT

[5] R. W. Hough. Future data traffic volume. Computer, 3(5):6–12, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bell:1970:CN

[6] C. G. Bell, A. N. Habermann, J. Mc-Credie, R. Rutledge, and W. Wulf. Computer networks. Computer, 3 (5):13–23, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holzman:1970:DSN

[7] L. N. Holzman and W. J. Lawless. Data Set 203 a new high-speed Voideband modem. Computer, 3(5):24–30, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:AAa

[8] Anonymous. About the authors. Computer, 3(5):31–32, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:PIC

[9] Anonymous. Provisional IEEE Computer Society constitution and bylaws. Computer, 3(5):33–37, September/October 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larson:1970:CRP

[10] H. T. Larson. Computers in the representative process — III. Computer, 3(5):38–41, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:L

[11] Anonymous. Letters. Computer, 3(5): 42, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:NNMa

[12] Anonymous. News, notices and miscellanea. Computer, 3(5):43–46, September/October 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:SFE

[13] Anonymous. A symposium on feature extraction and selection in pattern recognition. Computer, 3(5):50– 52, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:NPa

[14] Anonymous. New products. Computer, 3(5):53–56, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:CGRa

[15] Anonymous. Computer Group repository. Computer, 3(5):57–61, September/October 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Ba

[16] Anonymous. Bibliography. Computer, 3(5):62–68, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Aa

[17] Anonymous. Advertisement. Computer, 3(5):68, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Ab

[18] Anonymous. Advertisement. Computer, 3(5):c2, September/October 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Ca

[19] Anonymous. Computer. Computer, 3 (5):c1, September/October 1970. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:CPI

[20] Anonymous. Call for papers 1971 IEEE Computer Society conference. Computer, 3(6):1, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:CEb

[21] Anonymous. Calendar of events. Computer, 3(6):2–34, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:TCb

[22] Anonymous. Table of contents. Computer, 3(6):4, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuller:1970:MT

[23] R. H. Fuller. Memory trends in the 70’s. Computer, 3(6):5–6, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gunderson:1970:SEA

[24] D. C. Gunderson. Some effects of advances in memory system technology on computer organization. Computer, 3(6):7–11, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Petschauer:1970:TME

[25] R. J. Petschauer. Trends in memory element and subsystem design in the 1970’s. Computer, 3(6):12–17, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patton:1970:TDO

[26] P. C. Patton. Trends in data organization and access methods. Computer, 3(6):18–24, November/December 1970.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:AAb

[27] Anonymous. About the authors. Computer, 3(6):25, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:WCD

[28] Anonymous. ”what’s cooking?” dept. Computer, 3(6):26, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:CDW

[29] Anonymous. Computer diagnosis — will it help cure you or kill you? Computer, 3(6):27–28, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:NNMb

[30] Anonymous. News, notices and miscellanea. Computer, 3(6):29, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:FTP

[31] Anonymous. 70/FJCC technical program. Computer, 3(6):30–31, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:GIC

[32] Anonymous. The Great IEEE Computer Society Logo Contest. Computer, 3(6):32, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:CGRb

[33] Anonymous. Computer Group repository. Computer, 3(6):35–37, November/December 1970. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:NPb

[34] Anonymous. New products. Computer, 3(6):38–41, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Bb

[35] Anonymous. Bibliography. Computer, 3(6):42–48, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Ac

[36] Anonymous. Advertisement. Computer, 3(6):48, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Ad

[37] Anonymous. Advertisement. Computer, 3(6):c2, November/December 1970. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1970:Cb

[38] Anonymous. Computer. Computer, 3 (6):c1, November/December 1970. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:PM

[39] Anonymous. President’s message. Computer, 4(1):1–2, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TCa

[40] Anonymous. Table of contents. Computer, 4(1):3, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CPa

[41] Anonymous. Call for papers. Computer, 4(1):4, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Avizienis:1971:FTC

[42] A. Avizienis. Fault-tolerant computing: An overview. Computer, 4(1):5–8, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carter:1971:SFT

[43] W. C. Carter and W. G. Bouricius. A survey of fault tolerant computer architecture and its evaluation. Computer, 4 (1):9–16, January/February 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McCluskey:1971:TDP

[44] E. J. McCluskey. Test and diagnosis procedure for digital networks. Computer, 4(1):17–20, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elspas:1971:SR

[45] B. Elspas, M. W. Green, and K. N. Levitt. Software reliability. Computer, 4(1):21–27, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Short:1971:SSA

[46] R. A. Short and J. Goldberg. A summary of Soviet activities in the design of fault-tolerant digital machines. Computer, 4(1):28–33, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:AAa

[47] Anonymous. About the authors. Computer, 4(1):34–36, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ISFa

[48] Anonymous. International symposium on fault-tolerant computing. Computer, 4(1):37–44, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baechler:1971:ACC

[49] D. O. Baechler. Aerospace computer characteristics and design trends. Computer, 4(1):45–57, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CMS

[50] Anonymous. Computer memories in the Seventies. Computer, 4(1):58–59, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ca

[51] Anonymous. The cookbook. Computer, 4(1):60, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CRP

[52] Anonymous. Catch a rising problem … and never ever let it go. Computer, 4(1):61–65, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSO

[53] Anonymous. Computer Society officers and members of the Board of Governors for 1971. Computer, 4(1):66–68, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NNMa

[54] Anonymous. News, notices and miscellanea. Computer, 4(1):69–73, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NCAa

[55] Anonymous. Newsletter for computeraided design. Computer, 4(1):74–77, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Q

[56] Anonymous. Questionnaire. Computer, 4(1):78, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NPa

[57] Anonymous. New products. Computer, 4(1):79–82, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSRa

[58] Anonymous. Computer Society repository. Computer, 4(1):83–87, January/February 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ba

[59] Anonymous. Bibliography. Computer, 4(1):88–96, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Aa

[60] Anonymous. Advertisement. Computer, 4(1):96, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CPb

[61] Anonymous. Call for papers. Computer, 4(1):96, January/February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cb

[62] Anonymous. Computer. Computer, 4 (1):c1, January/February 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:SSC

[63] Anonymous. Solid-state circuits journal. Computer, 4(1):c2, January/ February 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:LRA

[64] Anonymous. The last 64 reasons are the Atlantic City hotels, who handle reservations on a first come basis.

Computer, 4(2):1, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:SCB

[65] Anonymous. This summer? combine business with pleasure. Computer, 4 (2):2, March/April 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TCb

[66] Anonymous. Table of contents. Computer, 4(2):3, March/April 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cc

[67] Anonymous. Calender. Computer, 4 (2):4–38, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Allison:1971:DIS

[68] D. F. Allison. The dramatic impact of silicon technology. Computer, 4(2): 5, March/April 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Moore:1971:SRS

[69] G. E. Moore. Semiconductor rams — a status report. Computer, 4(2):6– 10, March/April 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kute:1971:DCS

[70] B. A. Kute. Design considerations for semiconductor ramdom access memory systems. Computer, 4(2):11– 17, March/April 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gray:1971:DCB

[71] J. Gray. Design considerations for a bipolar 256-bit random access memory. Computer, 4(2):18–22, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

House:1971:SMM

[72] D. L. House and R. A. Henzel. Semiconductor memories and minicomputers. Computer, 4(2):23–29, March/ April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:AAb

[73] Anonymous. About the authors. Computer, 4(2):30, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSa

[74] Anonymous. IEEE Computer Society workshop on computer communication. Computer, 4(2):31–36, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:La

[75] Anonymous. Letters. Computer, 4(2): 37, March/April 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NNMb

[76] Anonymous. News, notices and miscellanea. Computer, 4(2):38–41, March/ April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NCAb

[77] Anonymous. Newsletter for computeraided design. Computer, 4(2):42, March/April 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSRb

[78] Anonymous. Computer Society repository. Computer, 4(2):43–45, March/ April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NPb

[79] Anonymous. New products. Computer, 4(2):46–47, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Bb

[80] Anonymous. Bibliography. Computer, 4(2):48–54, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:YIS

[81] Anonymous. You are invited SCi to become a member of Simulation Councils, Inc. Computer, 4(2):55, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CPc

[82] Anonymous. Call for papers. Computer, 4(2):56, March/April 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSb

[83] Anonymous. IEEE Computer Society formerly IEEE Computer Group membership grade requirements. Computer, 4(2):56, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:WCYa

[84] Anonymous. What can you save if you eliminate ribbons and ink fountains from your printout design? Computer, 4(2):56, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cd

[85] Anonymous. Computer. Computer, 4(2):c1, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TRP

[86] Anonymous. There are 149 reasons to pre-register for the SJCC. Computer, 4(2):c2, March/April 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:WCYb

[87] Anonymous. What can you save if you eliminate ribbons and ink fountains from your printout design? Computer, 4(3):1, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSc

[88] Anonymous. IEEE Computer Society. Computer, 4(3):2, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TCc

[89] Anonymous. Table of contents. Computer, 4(3):3, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ce

[90] Anonymous. Calendar. Computer, 4 (3):4–44, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peterson:1971:SPS

[91] N. D. Peterson. A standard pattern for sequential file maintenance. Computer, 4(3):5–9, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bonn:1971:SCN

[92] T. H. Bonn. A standard for computer networks. Computer, 4(3):10– 14, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huang:1971:HFF

[93] T. S. Huang. How the Fast Fourier Transform got its name. Computer, 4 (3):15, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ADA

[94] Anonymous. 8th Annual Design Automation Workshop. Computer, 4(3): 16–24, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:WTC

[95] Anonymous. Whither teledisplay control standards? Computer, 4(3): 25, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:SJC

[96] Anonymous. 1971 Spring Joint Computer Conference May 18–20 in Atlantic City. Computer, 4(3):26–27, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICP

[97] Anonymous. International Conference on Pattern Recognition. Computer, 4 (3):28–29, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TOS

[98] Anonymous. Terminal-oriented systems. Computer, 4(3):30–31, May/ June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CST

[99] Anonymous. Computer Society technical committees. Computer, 4(3):32– 34, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cf

[100] Anonymous. The cookbook. Computer, 4(3):35, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CHH

[101] Anonymous. The computer: Hell or heaven. Computer, 4(3):36, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Lb

[102] Anonymous. Letters. Computer, 4 (3):37, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NNMc

[103] Anonymous. News, notices and miscellanea. Computer, 4(3):38, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:BNC

[104] Anonymous. Battin named Chairman of ’71 Computer Conference. Computer, 4(3):39, May/June 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NCAc

[105] Anonymous. Newsletter for computeraided design. Computer, 4(3):45– 46, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSRc

[106] Anonymous. Computer Society repository. Computer, 4(3):47–49, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NPc

[107] Anonymous. New products. Computer, 4(3):50–51, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:COM

[108] Anonymous. Computer-oriented methods in pattern recognition. Computer, 4(3):52, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CRC

[109] Anonymous. Change ROM contents … while enjoying MOS cost saving. Computer, 4(3):52, May/June 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICG

[110] Anonymous. IEEE Computer Group formerly IEEE Computer Group membership grade requirements. Computer, 4(3):52, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cg

[111] Anonymous. Computer. Computer, 4 (3):c1, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:DBS

[112] Anonymous. Datapac builds and sells more ROMs for more applications than anyone else in the world. Computer, 4 (3):c2, May/June 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TCd

[113] Anonymous. Table of contents. Computer, 4(4):1, July/August 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:IIC

[114] Anonymous. 1971 IEEE International Computer Society Conference. Computer, 4(4):2–11, July/August 1971.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bennetts:1971:FDD

[115] R. G. Bennetts and D. W. Lewin. Fault diagnosis of digital systems — a review. Computer, 4(4):12–20, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:AAc

[116] Anonymous. About the authors. Computer, 4(4):21, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilting:1971:APM

[117] J. J. Wilting and H. van den Berge. Air-pollution monitoring network in The Netherlands. Computer, 4(4):22– 27, July/August 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:GP

[118] Anonymous. The Galindo Petition. Computer, 4(4):28–29, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ISFb

[119] Anonymous. 1972 International Symposium on Fault Tolerant Computing. Computer, 4(4):30–32, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ch

[120] Anonymous. Calendar. Computer, 4 (4):33, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Larson:1971:WB

[121] H. T. Larson. On whistle blowing. Computer, 4(4):34, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NNMd

[122] Anonymous. News, notices and miscellanea. Computer, 4(4):35–37, July/ August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CCS

[123] Anonymous. Changes to Computer Society bylaws. Computer, 4(4):38– 39, July/August 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSRd

[124] Anonymous. Computer Society repository. Computer, 4(4):40–43, July/ August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NPd

[125] Anonymous. New products. Computer, 4(4):44–46, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CAa

[126] Anonymous. Classified ads. Computer, 4(4):47, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CU

[127] Anonymous. This coupon is ugly. Computer, 4(4):48, July/August 1971.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSd

[128] Anonymous. IEEE Computer Society publications. Computer, 4(4): 48, July/August 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:IYT

[129] Anonymous. If you think it’s hard to choose the right computer. Computer, 4(4):48, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ci

[130] Anonymous. Computer. Computer, 4(4):c1, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSe

[131] Anonymous. IEEE Computer Society. Computer, 4(4):c2, July/August 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ab

[132] Anonymous. Advertisement. Computer, 4(5):1, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSf

[133] Anonymous. 1971 IEEE Computer Society Conference. Computer, 4(5): 2, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:TCe

[134] Anonymous. Table of contents. Computer, 4(5):3, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSg

[135] Anonymous. IEEE Computer Society. Computer, 4(5):4, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cj

[136] Anonymous. Calendar. Computer, 4 (5):5, September/October 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

House:1971:AM

[137] D. L. House. The applications of minicomputers. Computer, 4(5):6, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Henzel:1971:SIA

[138] R. A. Henzel. Some industrial applications of minicomputers. Computer, 4(5):7–12, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ball:1971:CM

[139] C. J. Ball. Communications and the minicomputer. Computer, 4(5):13–21, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:FMS

[140] Anonymous. Fred Moses Sr. programmer. Computer, 4(5):22–26, September/October 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:AAd

[141] Anonymous. About the authors. Computer, 4(5):27, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NLA

[142] Anonymous. Non-logic aspects of logic devices. Computer, 4(5):28–29, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CFI

[143] Anonymous. Computers and freedom, individuality and automation: Challenge and opportunity. Computer, 4(5):30–31, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Ck

[144] Anonymous. The cookbook. Computer, 4(5):32–33, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSS

[145] Anonymous. Computer Society strengthens chapter activities. Computer, 4(5):34–36, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:DVP

[146] Anonymous. Distinguished Visitors Program 1971–72. Computer, 4(5):36– 38, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NNMe

[147] Anonymous. News, notices and miscellanea. Computer, 4(5):39–40, September/October 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CAb

[148] Anonymous. Computer applications. Computer, 4(5):41, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CPd

[149] Anonymous. Call for papers. Computer, 4(5):42, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CSRe

[150] Anonymous. Computer Society repository. Computer, 4(5):43–45, September/October 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:NPe

[151] Anonymous. New products. Computer, 4(5):46–47, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:CAc

[152] Anonymous. Classified ads. Computer, 4(5):48, September/October 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:ICSh

[153] Anonymous. IEEE Computer Society publications. Computer, 4(5):48, September/October 1971. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:SIY

[154] Anonymous. See if you can find our limits. Computer, 4(5):48, September/October 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:Cl

[155] Anonymous. Computer. Computer, 4 (5):c1, September/October 1971. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1971:DAT

[156] Anonymous. Dial access technical education. Computer, 4(5):c2, September/October 1971. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSa

[157] Anonymous. IEEE Computer Society. Computer, 5(1):1, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ca

[158] Anonymous. Calendar. Computer, 5 (1):2, January/February 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:TCa

[159] Anonymous. Table of contents. Computer, 5(1):3, January/February 1972.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PRDa

[160] Anonymous. Pattern recognition data bases available. Computer, 5(1):4–5, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:STJ

[161] Anonymous. Some thoughts on the JCCs, professionalism and Society goals. Computer, 5(1):6, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UICa

[162] Anonymous. Update information for the computer professional. Computer, 5 (1):7–14, January/February 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PBC

[163] Anonymous. Public believes computers make life better but expresses major concerns. Computer, 5(1):15–16, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CPa

[164] Anonymous. Call for papers. Computer, 5(1):17, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:OMI

[165] Anonymous. Officers, members of IEEE Computer Society Governing Board. Computer, 5(1):18–21, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Unger:1972:NH

[166] S. H. Unger. The need for heroes. Computer, 5(1):22–23, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NLA

[167] Anonymous. Non-logic aspects of logic devices. Computer, 5(1):24, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lundstrom:1972:PAL

[168] S. F. Lundstrom and B. E. Cowart. Physical aspects of logical devices. Computer, 5(1):25–30, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Greene:1972:PRT

[169] F. S. Greene. Power reduction techniques for LSI memory. Computer, 5(1):31–39, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cb

[170] Anonymous. Compcon72. Computer, 5(1):40, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UCO

[171] Anonymous. An undergraduate course on operating systems principles. Computer, 5(1):40–59, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CFN

[172] Anonymous. The computing field — a national resource. Computer, 5(1):60– 62, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CSR

[173] Anonymous. Computer Society repository. Computer, 5(1):63–70, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NPa

[174] Anonymous. New products. Computer, 5(1):71–74, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frost:1972:CA

[175] C. Frost. Computer applications. Computer, 5(1):75–78, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:FNF

[176] Anonymous. First names in freer enterprise. Computer, 5(1):79, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cc

[177] Anonymous. Compcon72. Computer, 5(1):80, January/February 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSb

[178] Anonymous. IEEE Computer Society publications. Computer, 5(1):80, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cd

[179] Anonymous. Computer. Computer, 5 (1):c1, January/February 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:MTAa

[180] Anonymous. Minicomputers — trends and applications. Computer, 5(1):c2, January/February 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ITM

[181] Anonymous. The IC troubleshooters march on. Computer, 5(2):1, March/ April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ISFa

[182] Anonymous. The 1972 International Symposium on Fault-Tolerant Computing. Computer, 5(2):2, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:TCb

[183] Anonymous. Table of contents. Computer, 5(2):3, March/April 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:SSC

[184] Anonymous. This space for sale in Computer. Computer, 5(2):4, March/ April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:JICa

[185] Anonymous. To join the IEEE Computer Society. Computer, 5(2):5, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PC

[186] Anonymous. On publications, chapters. Computer, 5(2):6, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ce

[187] Anonymous. Calendar. Computer, 5 (2):7, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PRDb

[188] Anonymous. Pattern recognition data bases available. Computer, 5(2):8, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:SCC

[189] Anonymous. Symposium on computercommunications networks and teletraffic. Computer, 5(2):9, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UICb

[190] Anonymous. Update information for the computer professional. Computer, 5 (2):10–15, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cf

[191] Anonymous. Cookbook. Computer, 5 (2):16–17, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foster:1972:CA

[192] C. C. Foster. Computer architecture. Computer, 5(2):18–19, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flynn:1972:SRM

[193] M. J. Flynn and A. Podvin. Shared resource multiprocessing. Computer, 5 (2):20–28, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bell:1972:ETN

[194] C. G. Bell, R. Chen, and S. Rege. Effect of technology on near term computer structures. Computer, 5(2):29– 38, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foster:1972:NTG

[195] C. C. Foster. The next three generations. Computer, 5(2):39–42, March/ April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simmons:1972:AWL

[196] D. B. Simmons. The art of writing large programs. Computer, 5(2):43– 49, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ra

[197] Anonymous. Repository. Computer, 5 (2):50–55, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NPb

[198] Anonymous. New products. Computer, 5(2):56–61, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NAa

[199] Anonymous. New applications. Computer, 5(2):62–63, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSc

[200] Anonymous. IEEE Computer Society publications. Computer, 5(2): 64, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:SBI

[201] Anonymous. ’72 SJCC brings it all together …. Computer, 5(2): 64, March/April 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cg

[202] Anonymous. Compcon72. Computer, 5(2):c2, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ch

[203] Anonymous. Computer. Computer, 5(2):c1, March/April 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:SWM

[204] Anonymous. Should we merge with the ACM and leave the IEEE? Computer, 5(3):1, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:TCc

[205] Anonymous. Table of contents. Computer, 5(3):2, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:JICb

[206] Anonymous. To join the IEEE Computer Society. Computer, 5(3):3, May/ June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amdahl:1972:CAP

[207] L. Amdahl. Concerned about present or future EDP needs. Computer, 5 (3):4–5, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ISFb

[208] Anonymous. The 1972 International Symposium on Fault-Tolerant Computing. Computer, 5(3):6–7, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NTP

[209] Anonymous. For a no-tea-party ACM 72. Computer, 5(3):8, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:MTAb

[210] Anonymous. Minicomputers: Trends and applications. Computer, 5(3):9– 10, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:La

[211] Anonymous. Letters. Computer, 5 (3):11, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UICc

[212] Anonymous. Update information for the computer professional. Computer, 5(3):11–14, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ci

[213] Anonymous. Calendar. Computer, 5 (3):14, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:WMC

[214] Anonymous. Workshop on modular computer systems. Computer, 5(3): 15–17, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cj

[215] Anonymous. Cookbook. Computer, 5 (3):17, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Russo:1972:DA

[216] R. L. Russo. Design automation. Computer, 5(3):18–20, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Case:1972:EDA

[217] P. W. Case. Evolution of design automation. Computer, 5(3):21–22, May/ June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Breuer:1972:RDD

[218] M. A. Breuer. Recent developments in design automation. Computer, 5(3): 23–35, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lynn:1972:CAD

[219] D. K. Lynn. Computer aided design for large-scale integrated circuits. Computer, 5(3):36–45, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reed:1972:SDT

[220] I. S. Reed. Symbolic design techniques applied to a generalized computer. Computer, 5(3):46–52, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:DCH

[221] Anonymous. Do colleges help business as much as business helps colleges? Computer, 5(3):53, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Rb

[222] Anonymous. Repository. Computer, 5(3):54–57, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:MR

[223] Anonymous. Microfiche readers. Computer, 5(3):58–60, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:MGGa

[224] Anonymous. “Mayer’s Grand Guide”. Computer, 5(3):61, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NPc

[225] Anonymous. New products. Computer, 5(3):62–63, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PRDc

[226] Anonymous. Pattern recognition data bases available. Computer, 5(3): 64, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UJCa

[227] Anonymous. The USA–Japan Computer Conference. Computer, 5(3): 65, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PVa

[228] Anonymous. Point of view. Computer, 5(3):66–67, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:BYW

[229] Anonymous. Bring your worst computer design problem to San Francisco. bring your wife. you might leave one behind. Computer, 5(3):68, May/June 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSd

[230] Anonymous. IEEE Computer Society publications. Computer, 5(3): 68, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ck

[231] Anonymous. Computer. Computer, 5 (3):c1, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NAD

[232] Anonymous. Ninth Annual Design Automation Workshop. Computer, 5(3): c2, May/June 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CMB

[233] Anonymous. Computer Magazine becomes a monthly. Computer, 5(4): 1, July/August 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:TCd

[234] Anonymous. Table of contents. Computer, 5(4):2, July/August 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cl

[235] Anonymous. Calendar. Computer, 5 (4):3, July/August 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Lb

[236] Anonymous. Letters. Computer, 5(4): 4–5, July/August 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CPb

[237] Anonymous. Call for papers. Computer, 5(4):6, July/August 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UICd

[238] Anonymous. Update information for the computer professional. Computer, 5(4):7–15, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UJCb

[239] Anonymous. The USA–Japan Computer Conference. Computer, 5(4): 9, July/August 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:KIY

[240] Anonymous. Keepsake (if you ever stop using it). Computer, 5(4):11, July/August 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:TIN

[241] Anonymous. Two important new publications from the IEEE Computer Society. Computer, 5(4):12, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shemer:1972:CSI

[242] J. E. Shemer. Computer system instrumentation and performance measurement. Computer, 5(4):16–17, July/ August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferrari:1972:WCS

[243] D. Ferrari. Workload characterization and selection in computer performance measurement. Computer, 5(4): 18–24, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Arndt:1972:HMR

[244] F. R. Arndt and C. G. M. Oliver. Hardware monitoring of real-time computer system performance. Computer, 5(4): 25–29, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cm

[245] Anonymous. Compcon. Computer, 5 (4):30–38, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shemer:1972:ITS

[246] J. E. Shemer and J. B. Robertson. Instrumentation of time-shared systems. Computer, 5(4):39–48, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mrazek:1972:IU

[247] D. Mrazek. Industry update. Computer, 5(4):49–53, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Rc

[248] Anonymous. Repository. Computer, 5 (4):54–57, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NPd

[249] Anonymous. New products. Computer, 5(4):58–60, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NAb

[250] Anonymous. New applications. Computer, 5(4):61–64, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cn

[251] Anonymous. Cookbook. Computer, 5 (4):65–66, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PVb

[252] Anonymous. Point of view. Computer, 5(4):67, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Co

[253] Anonymous. Compcon72. Computer, 5(4):68, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSe

[254] Anonymous. IEEE Computer Society publications. Computer, 5(4): 68, July/August 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cq

[255] Anonymous. Computer. Computer, 5(4):c1, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICC

[256] Anonymous. International Conference on Computer Communication. Computer, 5(4):c2, July/August 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CDB

[257] Anonymous. Concerning dues, budgets and the end of the JCCs. Computer, 5 (5):1, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:TCe

[258] Anonymous. Table of contents. Computer, 5(5):2, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cr

[259] Anonymous. Calendar. Computer, 5 (5):3, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CPc

[260] Anonymous. Call for papers. Computer, 5(5):4, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Lc

[261] Anonymous. Letters. Computer, 5(5): 5, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UICe

[262] Anonymous. Update information for the computer professional. Computer, 5(5):5–11, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CPd

[263] Anonymous. Compcon 73 calls for papers; San Francisco in February. Computer, 5(5):7, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shirley:1972:PE

[264] R. R. Shirley. Performance evaluation. Computer, 5(5):12–15, September/October 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grochow:1972:UFT

[265] J. M. Grochow. Utility functions for time-sharing system performance evaluation. Computer, 5(5):16–19, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cooperman:1972:SEU

[266] J. A. Cooperman, H. W. Lynch, and W. H. Tetzlaff. SPG: An effective use of performance and usage data. Computer, 5(5):20–23, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Amiot:1972:ERB

[267] L. Amiot, N. K. Natarajan, and R. A. Aschenbrenner. Evaluating a remote batch processing system. Computer, 5(5):24–29, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:UJCc

[268] Anonymous. The USA–Japan Computer Conference. Computer, 5(5): 30, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Luderer:1972:MUU

[269] G. W. R. Luderer. On memory usage under microprogramming. Computer, 5(5):31–34, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CPe

[270] Anonymous. Call for papers. Computer, 5(5):35–38, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1972:BTC

[271] E. F. Miller. Bibliography on techniques of computer performance analysis. Computer, 5(5):39–40, September/October 1972. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:IA

[272] Anonymous. Index author. Computer, 5(5):41–47, September/October 1972.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CSG

[273] Anonymous. The Computer Society’s grass roots. Computer, 5(5):48–51, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Rd

[274] Anonymous. Repository. Computer, 5(5):52–56, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NAc

[275] Anonymous. New applications. Computer, 5(5):57–58, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:NPe

[276] Anonymous. New products. Computer, 5(5):58–62, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Cs

[277] Anonymous. Cookbook. Computer, 5 (5):63, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PVc

[278] Anonymous. Point of view. Computer, 5(5):64–65, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1972:BR

[279] H. S. Stone. Book review. Computer, 5 (5):66, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:PRDd

[280] Anonymous. Pattern recognition data bases available. Computer, 5(5):67, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSf

[281] Anonymous. IEEE Computer Society membership & publications. Computer, 5(5):68, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:MGGb

[282] Anonymous. “Mayer’s Grand Guide”. Computer, 5(5):69, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:ICSg

[283] Anonymous. IEEE Computer Society publications. Computer, 5(5):70– 71, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:GMW

[284] Anonymous. The gift for the man who knows everything. Computer, 5(5): 71, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:LID

[285] Anonymous. Local interviews during Compcon’72. Computer, 5(5):71, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:SMT

[286] Anonymous. Semiconductor memory test symposium. Computer, 5(5):71, September/October 1972. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:CDC

[287] Anonymous. Concepts in design of computer-oriented systems. Computer, 5(5):c2, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1972:Ct

[288] Anonymous. Computer. Computer, 5 (5):c1, September/October 1972. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Pa

[289] Anonymous. From the President. Computer, 6(1):1–2, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:IIS

[290] Anonymous. IEEE-interest subjects to get more coverage in ACM Computing Reviews. Computer, 6(1):3, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:TCa

[291] Anonymous. Table of contents. Computer, 6(1):4, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RCC

[292] Anonymous. Responding to the changes in the computer profession. Computer, 6(1):5, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPa

[293] Anonymous. Call for papers. Computer, 6(1):6, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ca

[294] Anonymous. Calendar. Computer, 6 (1):7, January/February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:La

[295] Anonymous. Letters. Computer, 6(1): 8–15, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CT

[296] Anonymous. Computers for teaching. Computer, 6(1):16–22, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hokom:1973:IS

[297] R. Hokom, A. Grebert, F. Gerbstadt, and R. Bock. Introduction to SPLM. Computer, 6(1):23–27, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Taylor:1973:MDL

[298] L. Booth Taylor, C. G. Bell, C. H. Coker, R. M. Glorioso, E. J. Mc-Cluskey, F. J. Mowle, and D. M. Robinson. Minicomputers in the Digital Laboratory Program. Computer, 6(1):28– 42, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AP

[299] Anonymous. Advance program. Computer, 6(1):30, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Branstan:1973:PPO

[300] D. K. Branstan. Privacy and protection in operating systems. Computer, 6(1):43–46, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:WMC

[301] Anonymous. Workshop on Modular Computer Systems. Computer, 6(1):47, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ra

[302] Anonymous. Repository. Computer, 6(1):48–51, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NPa

[303] M. G. Bienhoff. New products. Computer, 6(1):52–54, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NAa

[304] M. G. Bienhoff. New applications. Computer, 6(1):55, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPa

[305] Anonymous. Product profile. Computer, 6(1):56–57, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AIa

[306] Anonymous. Advertisers’ index. Computer, 6(1):58, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDa

[307] Anonymous. Pattern recognition data bases available. Computer, 6(1):59, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cb

[308] Anonymous. Compcon73. Computer, 6(1):60, January/February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSa

[309] Anonymous. IEEE Computer Society publications. Computer, 6(1):60, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cc

[310] Anonymous. Computer. Computer, 6 (1):c1, January/February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:WDD

[311] Anonymous. Waiting for delivery of disk systems? Computer, 6(1):c2, January/February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCa

[312] Anonymous. [front cover]. Computer, 6(2):c1, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Aa

[313] Anonymous. [advertisement]. Computer, 6(2):nil1, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:IASa

[314] Anonymous. Information about the Society. Computer, 6(2):nil2–nil3, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ab

[315] Anonymous. [advertisement]. Computer, 6(2):1, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoagland:1973:P

[316] Albert S. Hoagland. From the President. Computer, 6(2):2–3, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:TCb

[317] Anonymous. Table of contents. Computer, 6(2):4, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:LS

[318] Anonymous. List of staff. Computer, 6(2):4, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ac

[319] Anonymous. [advertisement]. Computer, 6(2):5, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cd

[320] Anonymous. Calendar. Computer, 6 (2):6, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPb

[321] Anonymous. Call for papers. Computer, 6(2):6–48, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICa

[322] Anonymous. Update: information for the computer systems design professional. Computer, 6(2):7–8, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NSOa

[323] Anonymous. New Society officers, Governing Board members elected. Computer, 6(2):9–10, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ad

[324] Anonymous. [advertisement]. Computer, 6(2):11–12, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hartung:1973:CC

[325] Albert F. Hartung. Computer communications. Computer, 6(2):13, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:IASb

[326] Anonymous. Information about the society. Computer, 6(2):14, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matteson:1973:CPD

[327] Ronald G. Matteson. Computer processing for data communications. Computer, 6(2):15–19, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Willner:1973:CPC

[328] Erich R. Willner and Joseph L. Lindinger. COS: A portable communications oriented system. Computer, 6 (2):20–26, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tymann:1973:BOC

[329] Brian Tymann. Bit-oriented control procedures for computer-to-computer communications systems. Computer, 6 (2):27–31, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Rb

[330] Anonymous. Repository. Computer, 6 (2):32–34, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NP

[331] Anonymous. New products. Computer, 6(2):35–39, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NA

[332] Anonymous. New applications. Computer, 6(2):39–40, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPb

[333] Anonymous. Product profile. Computer, 6(2):41–45, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSb

[334] Anonymous. IEEE Computer Society publications. Computer, 6(2):46– 47, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AIb

[335] Anonymous. Advertiser’s index. Computer, 6(2):48, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CAa

[336] Anonymous. Classified ads. Computer, 6(2):48, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:IASc

[337] Anonymous. Information about the Society. Computer, 6(2):nil4–nil5, February 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ae

[338] Anonymous. [advertisement]. Computer, 6(2):nil6, February 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCb

[339] Anonymous. [front cover]. Computer, 6(2):c2, February 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:TCc

[340] Anonymous. Table of contents. Computer, 6(3):5, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ISF

[341] Anonymous. 1973 International Symposium on Fault-Tolerant Computing. Computer, 6(3):6, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:U

[342] Anonymous. Update. Computer, 6 (3):7–9, May/June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPc

[343] Anonymous. Call for papers. Computer, 6(3):9, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abrams:1973:PMB

[344] M. D. Abrams and R. Rosenthal. On the passing of MOBIDIC-B. Computer, 6(3):10–18, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barbacci:1973:PND

[345] M. Barbacci, C. G. Bell, and D. Siewiorek. PMS: a notation to describe computer structures. Computer, 6(3):19–21, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barbacci:1973:IND

[346] M. Barbacci, C. G. Bell, and D. Siewiorek. ISP: a notation to describe a computer’s instruction sets. Computer, 6(3):22–24, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Winder:1973:DBC

[347] R. O. Winder. A data base for computer performance evaluation. Computer, 6(3):25–29, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kaplan:1973:CBC

[348] K. R. Kaplan and R. O. Winder. Cache-based computer systems. Computer, 6(3):30–36, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AF

[349] Anonymous. Attacking on the flank. Computer, 6(3):37–39, May/June 1973.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:YCE

[350] Anonymous. So you’re a computer expert! Computer, 6(3):40, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Rc

[351] Anonymous. Repository. Computer, 6(3):41–42, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NPb

[352] M. G. Bienhoff. New products. Computer, 6(3):43–46, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NAb

[353] M. G. Bienhoff. New applications. Computer, 6(3):47–48, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PVa

[354] Anonymous. Point of view. Computer, 6(3):49, May/June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSc

[355] Anonymous. IEEE Computer Society publications. Computer, 6(3):50– 51, May/June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPc

[356] Anonymous. Product profile. Computer, 6(3):52–56, May/June 1973.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CME

[357] Anonymous. Computers made easy. Computer, 6(3):56, May/June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CR

[358] Anonymous. Computing reviews. Computer, 6(3):56, May/June 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDb

[359] Anonymous. Pattern recognition data bases available. Computer, 6(3): 56, May/June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ce

[360] Anonymous. Computer. Computer, 6 (3):c1, May/June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:DAS

[361] Anonymous. Design automation system for microprogramming. Computer, 6(3):c2, May/June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hopewell:1973:AAS

[362] L. Hopewell, W. S. Chou, and H. Frank. Analysis of architectural strategies for a large message-switching network: a case study. Computer, 6 (4):26–34, April 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Binford:1973:CV

[363] T. O. Binford and J. M. Tenenbaum. Computer vision. Computer, 6(5):19–24, May 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Pb

[364] Anonymous. From the President. Computer, 6(6):1, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cf

[365] Anonymous. Calendar. Computer, 6(6):2–3, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPd

[366] Anonymous. Call for papers. Computer, 6(6):3, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cg

[367] Anonymous. Contents. Computer, 6(6):4, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSd

[368] Anonymous. IEEE Computer Society membership application. Computer, 6(6):5–6, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Lb

[369] Anonymous. Letters. Computer, 6 (6):7–8, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICb

[370] Anonymous. Update information for the computer systems design professional. Computer, 6(6):8–10, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:M

[371] Anonymous. Minicomputers. Computer, 6(6):11–12, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schultz:1973:GUL

[372] Gaymond W. Schultz, Raymond M. Holt, and Harold L. McFarland. A guide to using LSI microprocessors. Computer, 6(6):13–20, June 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zschau:1973:IDI

[373] Edwin V. W. Zschau. The IBM Diskette and its implications for minicomputer systems. Computer, 6(6): 21–26, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CM

[374] Anonymous. Calculators to mini’s. Computer, 6(6):27, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blanc:1973:MTA

[375] Robert Blanc. Minicomputer trends and applications 1973. Computer, 6 (6):28–29, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sloan:1973:CSE

[376] M. E. Sloan, C. L. Coates, and E. J. McCluskey. COSINE survey of electrical engineering departments. Computer, 6(6):30–39, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NPPa

[377] Anonymous. New products products and services for the computer systems design professional. Computer, 6(6):40–45, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NACa

[378] Anonymous. New applications computer systems at work. Computer, 6 (6):46–47, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RTPa

[379] Anonymous. Repository: Technical papers covering the full range of computer system design. Computer, 6 (6):48–51, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPCa

[380] Anonymous. Product profile: The computer system designer’s marketplace. Computer, 6(6):52–55, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AIc

[381] Anonymous. Advertisers’s index. Computer, 6(6):55–59, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSe

[382] Anonymous. IEEE Computer Society publications. Computer, 6(6): 60–61, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCc

[383] Anonymous. [front cover]. Computer, 6(6):c1, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDc

[384] Anonymous. Pattern recognition data bases available. Computer, 6(6):c2, June 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BCa

[385] Anonymous. [back cover]. Computer, 6(6):c4, June 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ch

[386] Anonymous. Contents. Computer, 6(7):1, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPe

[387] Anonymous. Call for papers. Computer, 6(7):1, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:LMH

[388] Anonymous. ‘I left my heart … ” [COMPCON 74]. Computer, 6(7):1, July 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ag

[389] Anonymous. [advertisements]. Computer, 6(7):1–2, July 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICc

[390] Anonymous. Update information for the computer systems design professional. Computer, 6(7):6–9, July 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ci

[391] Anonymous. Calendar. Computer, 6(7):10, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPf

[392] Anonymous. Call for papers. Computer, 6(7):11–12, July 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kempner:1973:CCM

[393] Kenneth M. Kempner. Computers in clinical medicine. Computer, 6 (7):13–14, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Caceres:1973:CAE

[394] C. A. Caceres. The case against electrocardiographic automation. Computer, 6(7):15–22, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Collen:1973:CMR

[395] Morris F. Collen and Lou S. Davis. Computerized medical records in multiphasic testing. Computer, 6(7):23–28, July 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sheppard:1973:DCS

[396] Louis C. Sheppard, Nicholas T. Kouchoukos, and John W. Kirklin. The digital computer in surgical intensive care automation. Computer, 6 (7):29–34, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NPPa

[397] Milton G. Bienhoff. New products: Products and services for the computer systems design professional. Computer, 6(7):35–38, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NACa

[398] Milton G. Bienhoff. New applications: Computer systems at work. Computer, 6(7):39–41, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RTPb

[399] Anonymous. Repository: Technical papers covering the full range of computer system design. Computer, 6 (7):42–43, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPCb

[400] Anonymous. Product profile: The computer system designer’s marketplace. Computer, 6(7):44–47, July 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AId

[401] Anonymous. Advertiser’s index. Computer, 6(7):47, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSf

[402] Anonymous. IEEE Computer Society publications. Computer, 6(7): 48–52, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCd

[403] Anonymous. [front cover]. Computer, 6(7):c1, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FICa

[404] Anonymous. [front inside cover]. Computer, 6(7):c2, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BICa

[405] Anonymous. [back inside cover]. Computer, 6(7):c3, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BCb

[406] Anonymous. [back cover]. Computer, 6(7):c4, July 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pyke:1973:CNT

[407] T. N. Pyke, Jr. and R. P. Blanc. Computer networking technology — a state of the art review. Computer, 6(8): 12–19, August 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Enslow:1973:NTI

[408] P. H. Enslow, Jr. Non-technical issues in network design-economic, legal, social, and other considerations. Computer, 6(8):20–30, August 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCe

[409] Anonymous. [front cover]. Computer, 6(9):c1, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FICb

[410] Anonymous. [front inside cover]. Computer, 6(9):c2, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cj

[411] Anonymous. Contents. Computer, 6 (9):1, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDd

[412] Anonymous. Pattern recognition data bases available. Computer, 6(9):1, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ah

[413] Anonymous. [advertisements]. Computer, 6(9):2, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:ICSg

[414] Anonymous. IEEE Computer Society membership application. Computer, 6(9):3–4, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:DNA

[415] Anonymous. Data networks: Analysis and design. Computer, 6(9):5, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ck

[416] Anonymous. Calendar. Computer, 6 (9):6–7, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPg

[417] Anonymous. Call for papers. Computer, 6(9):7, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Pc

[418] Anonymous. From the President. Computer, 6(9):8, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Rd

[419] Anonymous. The referendum. Computer, 6(9):9–10, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PVb

[420] Anonymous. Point of view. Computer, 6(9):11, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICd

[421] Anonymous. Update information for the computer systems design professional. Computer, 6(9):12–16, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Lc

[422] Anonymous. Letters. Computer, 6(9): 17, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laliotis:1973:CST

[423] Theodore A. Laliotis. Computer storage technology. Computer, 6(9):18–19, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laliotis:1973:MMT

[424] Theodore A. Laliotis. Main memory technology. Computer, 6(9):20–27, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoagland:1973:MSP

[425] A. S. Hoagland. Mass storage: Past, present and future. Computer, 6(9): 28–33, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hodges:1973:ACT

[426] David A. Hodges. Alternative component technologies for advanced memory systems. Computer, 6(9):34–37, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NPPb

[427] Milton G. Bienhoff. New products: Products and services for the computer systems design professional. Computer, 6(9):38–44, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bienhoff:1973:NACb

[428] Milton G. Bienhoff. New applications: Computer systems at work. Computer, 6(9):45–46, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RTPc

[429] Anonymous. Repository: Technical papers covering the full range of computer system design. Computer, 6(9):47–49, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPCc

[430] Anonymous. Product profile: The computer system designer’s marketplace. Computer, 6(9):50–53, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AIe

[431] Anonymous. Advertiser’s index. Computer, 6(9):53–58, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BICb

[432] Anonymous. [back inside cover]. Computer, 6(9):c3, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BCc

[433] Anonymous. [back cover]. Computer, 6(9):c4, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCf

[434] Anonymous. [front cover]. Computer, 6(10):c1, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDe

[435] Anonymous. Pattern recognition data bases. Computer, 6(10):c2, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Aj

[436] Anonymous. [advertisements]. Computer, 6(10):1–2, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cl

[437] Anonymous. Contents. Computer, 6 (10):3, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cn

[438] Anonymous. Calendar. Computer, 6 (10):4, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPh

[439] Anonymous. Call for papers. Computer, 6(10):5, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Pd

[440] Anonymous. From the President. Computer, 6(10):6, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICe

[441] Anonymous. Update information for the computer systems design professional. Computer, 6(10):7–12, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ellis:1973:MCS

[442] Robert A. Ellis. Modular computer systems. Computer, 6(10):13–14, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuller:1973:SOS

[443] Samuel H. Fuller and Daniel P. Siewiorek. Some observations on semiconductor technology and the architecture of large digital modules. Computer, 6(10):15–22, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grason:1973:CRT

[444] John Grason, C. Gordon Bell, and John Eggert. The commercialization of register transfer modules. Computer, 6(10):23–28, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ellis:1973:EES

[445] Robert A. Ellis. Experiences with an evolving system. Computer, 6(10):29– 34, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NPPb

[446] Anonymous. New products: Products and services for the computer systems design professional. Computer, 6(10):35–39, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NACb

[447] Anonymous. New applications computer systems at work. Computer, 6(10):40–41, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RTPd

[448] Anonymous. Repository: Technical papers covering the full range of computer system design. Computer, 6(10):42–52, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:AIf

[449] Anonymous. Advertiser’s index. Computer, 6(10):52, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPCd

[450] Anonymous. Product profile: The computer system designer’s marketplace. Computer, 6(10):53–55, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CSPa

[451] Anonymous. Computer society publications [advertisement]. Computer, 6(10):56–57, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BCd

[452] Anonymous. [back cover]. Computer, 6(10):c4, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCg

[453] Anonymous. [front cover]. Computer, 6(11):c1, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Ak

[454] Anonymous. [advertisements]. Computer, 6(11):c2, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Co

[455] Anonymous. Contents. Computer, 6 (11):1, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Pe

[456] Anonymous. Program. Computer, 6 (11):2–3, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDf

[457] Anonymous. Pattern recognition databases. Computer, 6(11):4–6, 1973.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CES

[458] Anonymous. The compleat executive secretary. Computer, 6(11):7, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cq

[459] Anonymous. Calendar. Computer, 6(11):7–8, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPi

[460] Anonymous. Call for papers. Computer, 6(11):9, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICf

[461] Anonymous. Update information for the computer systems design professional. Computer, 6(11):10–14, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jensen:1973:HVS

[462] E. Douglas Jensen. Hardware vs software: The two faces of computers. Computer, 6(11):15, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Browne:1973:IOS

[463] J. C. Browne and J. H. Howard. The interaction of operating systems and computer architecture: a workshop summary. Computer, 6(11):16–18, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atwood:1973:SPS

[464] J. William Atwood. I/O supervision in the Project Sue operating system. Computer, 6(11):19–24, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferrari:1973:AIM

[465] Domenico Ferrari. Architecture and instrumentation in a modular interactive system. Computer, 6(11):25–30, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gordon:1973:IAM

[466] Robert L. Gordon. The impact of automated memory management on software architecture. Computer, 6(11):31– 36, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NPPc

[467] Anonymous. New products: Products and services for the computer systems design professional. Computer, 6(11):37–41, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:HCa

[468] Anonymous. Haynes’ cookbook. Computer, 6(11):42–43, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NACc

[469] Anonymous. New applications computer systems at work. Computer, 6(11):44–46, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RTPe

[470] Anonymous. Repository: Technical papers covering the full range of computer system design. Computer, 6(11):47–51, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPCe

[471] Anonymous. Product profile: The computer system designer’s marketplace. Computer, 6(11):52–58, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CSPb

[472] Anonymous. Computer Society publications [advertisement]. Computer, 6(11):59–61, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BCe

[473] Anonymous. [back cover]. Computer, 6(11):c4, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FCh

[474] Anonymous. [front cover]. Computer, 6(12):c1, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:FICc

[475] Anonymous. [front inside cover]. Computer, 6(12):c2, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cs

[476] Anonymous. Contents. Computer, 6 (12):1, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Pf

[477] Anonymous. From the President. Computer, 6(12):2–6, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PRDg

[478] Anonymous. Pattern recognition data bases. Computer, 6(12):7, 1973. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:Cu

[479] Anonymous. Calendar. Computer, 6 (12):8, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CPj

[480] Anonymous. Call for papers. Computer, 6(12):9, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:UICg

[481] Anonymous. Update information for the computer systems design professional. Computer, 6(12):10–13, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NSOb

[482] Anonymous. New Society officers, governing board members. Computer, 6(12):14–17, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buxbaum:1973:CL

[483] Richard M. Buxbaum and Howard A. Latin. The computer and the law. Computer, 6(12):18–20, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fischer:1973:LIC

[484] L. Richard Fischer. Legal implications of a cashless society. Computer, 6(12):21–26, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fromholz:1973:LPP

[485] Haley J. Fromholz. Legal protection of proprietary software. Computer, 6(12):27–32, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fromholz:1973:CPI

[486] Haley J. Fromholz, William T. Barker, Michael S. Drucker, and Martin Mattes. Competition in the peripherals industry: a proposed statute. Computer, 6(12):33–34, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CSA

[487] Anonymous. COMPCON 74 Spring advance program. Computer, 6(12):35– 44, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:HCb

[488] Anonymous. Haynes’ cookbook. Computer, 6(12):45–47, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Higbie:1973:TSA

[489] L. C. Higbie. Tutorial: Supercomputer architecture. Computer, 6(12):48–58, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NPPd

[490] Anonymous. New products: Products and services for the computer systems design professional. Computer, 6(12):59–63, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:NACd

[491] Anonymous. New applications computer systems at work. Computer, 6(12):64–66, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:RTPf

[492] Anonymous. Repository: Technical papers covering the full range of computer system design. Computer, 6(12):67–68, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:PPCf

[493] Anonymous. Product profile: The computer system designer’s marketplace. Computer, 6(12):69–74, 1973. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:CSPc

[494] Anonymous. Computer Society publications [advertisement]. Computer, 6(12):75–77, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1973:BCf

[495] Anonymous. [back cover]. Computer, 6(12):c4, 1973. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ca

[496] Anonymous. Contents. Computer, 7 (1):1, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Pa

[497] Anonymous. From the President. Computer, 7(1):2–3, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cb

[498] Anonymous. Calendar. Computer, 7 (1):4, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPa

[499] Anonymous. Call for papers. Computer, 7(1):5–8, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Collmeyer:1974:DDA

[500] Arthur J. Collmeyer. Developments in design automation. Computer, 7 (1):9–12, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Toth:1974:ADD

[501] Attila Toth and Chip Holt. Automated database-driven digital testing. Computer, 7(1):13–20, January 1974.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mattison:1974:DAM

[502] Roland L. Mattison. Design automation of MOS artwork. Computer, 7 (1):21–28, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wieczner:1974:IGD

[503] Simon Wieczner. Interactive graphics in design automation. Computer, 7(1): 29–33, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Waaben:1974:LMW

[504] S. Waaben and G. C. Feth. LSI microprocessing workshop summary. Computer, 7(1):34–36, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kuck:1974:MPO

[505] David J Kuck, Paul P. Budnik, Shyh-Ching Chen, Duncan H. Lawrie, Ross A. Towle, Richard E. Strebendt, Edward W. Davis, Joseph Han, Paul W. Kraska, and Yoichi Muraoka. Measurements of parallelism in ordinary FORTRAN programs. Computer, 7(1):37–46, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ra

[506] Anonymous. Repository. Computer, 7 (1):47–48, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPa

[507] Anonymous. New products. Computer, 7(1):49–53, January 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAa

[508] Anonymous. New applications. Computer, 7(1):54–56, January 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPa

[509] Anonymous. Product profile. Computer, 7(1):57–61, January 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:LA

[510] Anonymous. List of advertisers. Computer, 7(1):62–66, January 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCa

[511] Anonymous. [front cover]. Computer, 7(1):c1, January 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cc

[512] Anonymous. Contents. Computer, 7 (2):1, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Pb

[513] Anonymous. From the President. Computer, 7(2):2–3, February 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cd

[514] Anonymous. Calendar. Computer, 7 (2):4, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPb

[515] Anonymous. Call for papers. Computer, 7(2):5, February 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ua

[516] Anonymous. Update. Computer, 7 (2):6–10, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hartung:1974:CNC

[517] A. F. Hartung. Computer networks and communications. Computer, 7(2): 11–12, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Doll:1974:TTC

[518] Dixon R. Doll. Telecommunications turbulence and the computer network evolution. Computer, 7(2):13– 22, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Winkler:1974:DSC

[519] Stanley Winkler and Lee Danner. Data security in the computer communication environment. Computer, 7(2): 23–31, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schutz:1974:DCS

[520] Gerald C. Schutz and George E. Clark. Data communication standards. Computer, 7(2):32–41, February 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shar:1974:MSI

[521] Leonard E. Shar and Edward S. Davidson. A multiminiprocessor system implemented through pipelining. Computer, 7(2):42–51, February 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPb

[522] Anonymous. New products. Computer, 7(2):52–56, February 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAb

[523] Anonymous. New applications. Computer, 7(2):57–59, February 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rb

[524] Anonymous. Repository. Computer, 7 (2):60–69, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CSPa

[525] Anonymous. Computer Society publications [advertisement]. Computer, 7 (2):70–72, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCb

[526] Anonymous. [front cover]. Computer, 7(2):c1, February 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ce

[527] Anonymous. Contents. Computer, 7(3):1, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Pc

[528] Anonymous. From the President. Computer, 7(3):2–3, March 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cf

[529] Anonymous. Calendar. Computer, 7(3):4, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPc

[530] Anonymous. Call for papers. Computer, 7(3):5, March 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ub

[531] Anonymous. Update. Computer, 7 (3):6–14, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nilsen:1974:DFC

[532] Ragnar N. Nilsen. Distributed-function computer architectures. Computer, 7 (3):15–16, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Joseph:1974:IHH

[533] Earl C Joseph. Innovations in heterogeneous and homogeneous distributedfunction architectures. Computer, 7 (3):17–24, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Baer:1974:MDS

[534] Jean-Loup Baer. Models for the design, simulation, and performance of distributed-function architecture. Computer, 7(3):25–30, March 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Farber:1974:SCD

[535] David J. Farber. Software considerations in distributed architectures. Computer, 7(3):31–35, March 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Poppelbaum:1974:IHA

[536] W. J. Poppelbaum. Information hardware again. Computer, 7(3):36–37, March 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PCa

[537] Anonymous. Paynes’ cookbook. Computer, 7(3):38–40, March 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weisbecker:1974:SMA

[538] Joe Weisbecker. A simplified microcomputer architecture. Computer, 7 (3):41–47, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPc

[539] Anonymous. New products. Computer, 7(3):48–52, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAc

[540] Anonymous. New applications. Computer, 7(3):53–54, March 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rc

[541] Anonymous. Repository. Computer, 7 (3):55–56, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPb

[542] Anonymous. Product profile. Computer, 7(3):57–65, March 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CSPb

[543] Anonymous. Computer society publications [advertisement]. Computer, 7 (3):66–68, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCc

[544] Anonymous. [front cover]. Computer, 7(3):c1, March 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cg

[545] Anonymous. Contents. Computer, 7(4):1, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CMR

[546] Anonymous. Computer magazine’s 1974 readership survey. Computer, 7 (4):1–2, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:WTK

[547] Anonymous. What is there to know about computer networks? Computer, 7(4):2–3, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RCC

[548] Anonymous. Report card on Compcon ’74 spring. Computer, 7(4):4–5, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Uc

[549] Anonymous. Update. Computer, 7 (4):6–10, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RRH

[550] Anonymous. Reach for the right handbook! [advertisement]. Computer, 7(4):11, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ch

[551] Anonymous. Calendar. Computer, 7(4):12, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPd

[552] Anonymous. Call for papers. Computer, 7(4):13, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CEEa

[553] Anonymous. Continuing engineering education: Summer short courses [advertisement]. Computer, 7(4):14–16, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Galey:1974:CAD

[554] J. Michael Galey. Computer-aided design: Selections from the 10th annual design automation workshop. Computer, 7(4):17, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hightower:1974:IPT

[555] David W. Hightower. The interconnection problem: a tutorial. Computer, 7(4):18–32, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Freeman:1974:ASD

[556] Peter Freeman. Automating software design. Computer, 7(4):33–38, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ulrich:1974:CSN

[557] E. G. Ulrich and T. Baker. Concurrent simulation of nearly identical digital networks. Computer, 7(4):39–44, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dooner:1974:IGP

[558] Nitta P. Dooner, Janice R. Lourie, and Pat Velderman. An interactive graphic and process controlled system for composing and sampling loom constrained designs. Computer, 7(4):45–49, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDa

[559] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (4):50, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Aa

[560] Anonymous. [advertisements]. Computer, 7(4):51, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Theis:1974:VS

[561] D. J. Theis. Vector supercomputers. Computer, 7(4):52–61, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPd

[562] Anonymous. New products. Computer, 7(4):62–65, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAa

[563] Anonymous. Recent IC announcements. Computer, 7(4):66, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAd

[564] Anonymous. New applications. Computer, 7(4):67–69, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rd

[565] Anonymous. Repository. Computer, 7(4):70–71, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPc

[566] Anonymous. Product profile. Computer, 7(4):72–75, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAa

[567] Anonymous. Short courses [advertisement]. Computer, 7(4):76–77, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ABHa

[568] Anonymous. Announcing the big HP calculator giveaway! For IEEE Computer Society members only! [advertisement]. Computer, 7(4):78, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:HCGa

[569] Anonymous. HP calculator giveaway! Society members only! [advertisement]. Computer, 7(4):79–81, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCa

[570] Anonymous. [back cover]. Computer, 7(4):c4, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCd

[571] Anonymous. [front cover]. Computer, 7(4):c1, April 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCa

[572] Anonymous. [inside front cover]. Computer, 7(4):c2, April 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ci

[573] Anonymous. Contents. Computer, 7(5):1, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:AC

[574] Anonymous. About the cover. Computer, 7(5):2, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CRAa

[575] Anonymous. Computer references [advertisement]. Computer, 7(5):3–5, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ABHb

[576] Anonymous. Announcing the big HP calculator giveaway! for IEEE Computer Society members only! [advertisement]. Computer, 7(5):6, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:HCGb

[577] Anonymous. HP calculator giveaway! Society members only! [advertisement]. Computer, 7(5):7, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cj

[578] Anonymous. Calendar. Computer, 7(5):8, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPe

[579] Anonymous. Call for papers. Computer, 7(5):9, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ford:1974:U

[580] Gerald R. Ford. Update. Computer, 7(5):10–13, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ab

[581] Anonymous. [advertisements]. Computer, 7(5):14–16, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andrews:1974:DIP

[582] Harry C. Andrews. Digital image processing. Computer, 7(5):17–19, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDb

[583] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (5):20, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ad

[584] Anonymous. [advertisements]. Computer, 7(5):21, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Habibi:1974:SOD

[585] Ali Habibi and Guner S. Robinson. A survey ol digital picture coding. Computer, 7(5):22–34, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CNT

[586] Anonymous. Computer networks: Trends and applications [advertisement]. Computer, 7(5):35, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andrews:1974:DIR

[587] Harry C. Andrews. Digital image restoration: a survey. Computer, 7 (5):36–45, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Patterson:1974:DIP

[588] C. L. Patterson and Guenther Buechler. Digital image processing at the Aerospace Corporation. Computer, 7 (5):46–52, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ae

[589] Anonymous. [advertisements]. Computer, 7(5):53, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPRa

[590] Anonymous. Program progress report. Computer, 7(5):54–56, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hunt:1974:DIP

[591] B. R. Hunt and D. H. Janney. Digital image processing at Los Alamos Scientific Laboratory. Computer, 7 (5):57–62, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wecksung:1974:DIP

[592] George W. Wecksung and Katherine Campbell. Digital image processing at EG&G. Computer, 7(5):63–71, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ekstrom:1974:DIP

[593] M. P. Ekstrom. Digital image processing at Lawrence Livermore Laboratory. Part I: Diagnostic radiography applications. Computer, 7(5):72–78, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PR

[594] Anonymous. Pattern recognition. Computer, 7(5):79–80, May 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mayall:1974:DIP

[595] Brian H. Mayall. Digital image processing at Lawrence Livermore Laboratory: Part II: Biomedical applications. Computer, 7(5):81–87, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPe

[596] Anonymous. New products. Computer, 7(5):88–91, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAb

[597] Anonymous. Recent IC announcements. Computer, 7(5):92, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAe

[598] Anonymous. New applications. Computer, 7(5):93–95, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Re

[599] Anonymous. Repository. Computer, 7(5):96–97, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPd

[600] Anonymous. Product profile. Computer, 7(5):98–102, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAa

[601] Anonymous. Classified ads. Computer, 7(5):103, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAb

[602] Anonymous. Short courses [advertisement]. Computer, 7(5):104, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ACS

[603] Anonymous. Announcement computer science conference [advertisement]. Computer, 7(5):105, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CSPc

[604] Anonymous. Computer Society publications [advertisement]. Computer, 7 (5):106–108, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Af

[605] Anonymous. [advertisements]. Computer, 7(5):109, May 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCb

[606] Anonymous. [back cover]. Computer, 7(5):c4, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCe

[607] Anonymous. [front cover]. Computer, 7(5):c1, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCb

[608] Anonymous. [inside front cover]. Computer, 7(5):c2, May 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ah

[609] Anonymous. [advertisements]. Computer, 7(6):1, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ai

[610] Anonymous. [advertisements]. Computer, 7(6):1, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ck

[611] Anonymous. Contents. Computer, 7(6):1, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Pd

[612] Anonymous. From the President. Computer, 7(6):2, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simmons:1974:SRV

[613] Dick Simmons. A special report from the Vice President for Chapter Activities. Computer, 7(6):3, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cl

[614] Anonymous. Calendar. Computer, 7(6):4, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cm

[615] Anonymous. Calendar. Computer, 7(6):4–5, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPf

[616] Anonymous. Call for papers. Computer, 7(6):5, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ud

[617] Anonymous. Update. Computer, 7 (6):6–11, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ag

[618] Anonymous. [advertisements]. Computer, 7(6):12, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CRAb

[619] Anonymous. Contest rules [advertisement]. Computer, 7(6):13, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:MAa

[620] Anonymous. Membership application. Computer, 7(6):14, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IDS

[621] Anonymous. IBM data security study presented at NCC [advertisement]. Computer, 7(6):15–17, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:YPY

[622] Anonymous. Are you in the picture yet? [advertisement]. Computer, 7 (6):18–20, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Muntz:1974:OS

[623] Richard R. Muntz. Operating systems. Computer, 7(6):21, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Popek:1974:PS

[624] Gerald J. Popek. Protection structures. Computer, 7(6):22–33, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Goldberg:1974:SVM

[625] Robert P. Goldberg. Survey of virtual machine research. Computer, 7 (6):34–45, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Akkoyunlu:1974:ICF

[626] Eralp Akkoyunlu, Arthur Bernstein, and Richard Schantz. Interprocess communication facilities for network operating systems. Computer, 7(6): 46–55, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPf

[627] Anonymous. New products. Computer, 7(6):56–59, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDc

[628] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (6):60, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAc

[629] Anonymous. Recent IC announcements. Computer, 7(6):61, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAf

[630] Anonymous. New applications. Computer, 7(6):62–63, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rf

[631] Anonymous. Repository. Computer, 7(6):64–67, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAb

[632] Anonymous. Classified ads. Computer, 7(6):67, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPe

[633] Anonymous. Product profile. Computer, 7(6):68–71, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CEEb

[634] Anonymous. Continuing engineering education: Summer short courses [advertisement]. Computer, 7(6):72, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAc

[635] Anonymous. Short courses [advertisement]. Computer, 7(6):73–74, June 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CSPd

[636] Anonymous. Computer Society publications [advertisement]. Computer, 7 (6):75–76, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCf

[637] Anonymous. [front cover]. Computer, 7(6):c1, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCc

[638] Anonymous. [inside front cover]. Computer, 7(6):c2, June 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cn

[639] Anonymous. Contents. Computer, 7(7):1, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FMa

[640] Anonymous. [front matter]. Computer, 7(7):1, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Co

[641] Anonymous. Calendar. Computer, 7(7):2, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPg

[642] Anonymous. Call for papers. Computer, 7(7):3, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ue

[643] Anonymous. Update. Computer, 7 (7):4–9, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ICSa

[644] Anonymous. IEEE Computer Society members are going to win HP calculators [advertisement]. Computer, 7(7):10, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CRAc

[645] Anonymous. Contest rules [advertisement]. Computer, 7(7):11, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:MAb

[646] Anonymous. Membership application. Computer, 7(7):12, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ICSb

[647] Anonymous. Important Computer Society bylaws changes. Computer, 7 (7):13, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ICSc

[648] Anonymous. IEEE Computer Society bylaws. Computer, 7(7):14–18, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nelson:1974:MAI

[649] Donald R. Nelson. Microprocessor architecture introduction. Computer, 7(7):19, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laliotis:1974:MPF

[650] Theodore A. Laliotis. Microprocessors present and future. Computer, 7 (7):20–24, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SWL

[651] Anonymous. SPIE has a way of looking at things [advertisement]. Computer, 7(7):25, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reyling:1974:CCM

[652] George Reyling. Considerations in choosing a microprogramable bit-sliced architecture. Computer, 7(7):26–29, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Herr:1974:MDI

[653] Joseph R. Herr. Microprocessor design for intelligent point-of-sale terminals. Computer, 7(7):30–34, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ward:1974:LMM

[654] Ann R. Ward. LSI microprocessors and microcomputers: a bibliography. Computer, 7(7):35–36, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CFR

[655] Anonymous. [Compcon 74 Fall registration form — advertisement]. Computer, 7(7):36–39, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Aj

[656] Anonymous. [advertisements]. Computer, 7(7):37, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPRb

[657] Anonymous. Program progress report. Computer, 7(7):38–39, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paige:1974:USP

[658] M. R. Paige and J. P. Benson. The use of software probes in testing FOR-TRAN programs. Computer, 7(7): 40–47, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sanderford:1974:SFT

[659] H. Britton Sanderford. Special feature: The TORC TP III. Computer, 7(7):48–51, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:DNA

[660] Anonymous. Data networks: Analysis and design. Computer, 7(7):52–54, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPg

[661] Anonymous. New products. Computer, 7(7):55–58, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAd

[662] Anonymous. Recent IC announcements. Computer, 7(7):59, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAg

[663] Anonymous. New applications. Computer, 7(7):60–61, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rg

[664] Anonymous. Repository. Computer, 7(7):62–64, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPf

[665] Anonymous. Product profile. Computer, 7(7):65–69, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:AIa

[666] Anonymous. Advertiser’s index. Computer, 7(7):70, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAd

[667] Anonymous. Short courses [advertisement]. Computer, 7(7):71, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:OUM

[668] Anonymous. To order: Use the multipurpose order form inside back cover [advertisement]. Computer, 7(7):72, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CRS

[669] Anonymous. Computer reader service card. Computer, 7(7):73–74, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAc

[670] Anonymous. Compcon [advertisement]. Computer, 7(7):75, July 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCc

[671] Anonymous. [back cover]. Computer, 7(7):c4, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCg

[672] Anonymous. [front cover]. Computer, 7(7):c1, July 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cq

[673] Anonymous. Contents. Computer, 7(8):1, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:DYG

[674] Anonymous. Did you get in the picture? [advertisement]. Computer, 7 (8):1–2, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:LE

[675] Anonymous. Letters to the editor. Computer, 7(8):2, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:MTE

[676] Anonymous. A message from the technical editor. Computer, 7(8):2, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cr

[677] Anonymous. Calendar. Computer, 7(8):3, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPh

[678] Anonymous. Call for papers. Computer, 7(8):4, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:WWA

[679] Anonymous. Welcome to Wescon! [advertisement]. Computer, 7(8):5, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ICSd

[680] Anonymous. IEEE Computer Society members are going to win HP calculators [advertisement]. Computer, 7(8):6, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Uf

[681] Anonymous. Update. Computer, 7 (8):7–11, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:DTM

[682] Anonymous. Digital timing measurements [advertisement]. Computer, 7 (8):12, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:LTD

[683] Anonymous. Logic triggered displays [advertisement]. Computer, 7(8):

13–14, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoeffler:1974:ATP

[684] James D. Schoeffler. Advance technical program. Computer, 7(8):15–18, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nelson:1974:MA

[685] Donald R. Nelson. Microprocessor applications. Computer, 7(8):19, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Weisbecker:1974:PLC

[686] Joe Weisbecker. A practical, lowcost, home/school microprocessor system. Computer, 7(8):20–31, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDd

[687] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (8):32, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Temple:1974:UMA

[688] R. H. Temple and S. S. Devlin. The use of microprocessors as automobile on-board controllers. Computer, 7(8): 33–36, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPi

[689] Anonymous. Call for papers. Computer, 7(8):37–40, August 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clark:1974:MCE

[690] Richard J. Clark. A microprocessor controlled electronic distance meter. Computer, 7(8):41–47, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cropper:1974:MCT

[691] L. C. Cropper and J. W. Whiting. Microprocessors in CRT terminals. Computer, 7(8):48–53, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cs

[692] Anonymous. Calendar. Computer, 7 (8):54, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPh

[693] Anonymous. New products. Computer, 7(8):55–59, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAe

[694] Anonymous. Recent IC announcements. Computer, 7(8):60, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAh

[695] Anonymous. New applications. Computer, 7(8):61–63, August 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rh

[696] Anonymous. Repository. Computer, 7 (8):66–68, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAe

[697] Anonymous. Short courses [advertisement]. Computer, 7(8):68–69, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAd

[698] Anonymous. Classified ads. Computer, 7(8):69–70, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPg

[699] Anonymous. Product profile. Computer, 7(8):71–74, August 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RAB

[700] Anonymous. Recent additions to the bookshelf. Computer, 7(8):75–77, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ak

[701] Anonymous. Applications. Computer, 7(8):77, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCd

[702] Anonymous. [back cover]. Computer, 7(8):c4, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCh

[703] Anonymous. [front cover]. Computer, 7(8):c1, August 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCd

[704] Anonymous. [inside front cover]. Computer, 7(8):c2, August 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ct

[705] Anonymous. Contents. Computer, 7 (9):1, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IUB

[706] Anonymous. Indiana University, Bloomington. Computer, 7(9):1, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPj

[707] Anonymous. Call for papers. Computer, 7(9):2–3, September 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cu

[708] Anonymous. Calendar. Computer, 7 (9):4–5, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPk

[709] Anonymous. Call for papers. Computer, 7(9):6–15, September 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPl

[710] Anonymous. Call for papers. Computer, 7(9):17, September 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rine:1974:MVL

[711] David C. Rine. Multiple-valued logic and computer science and the 20th century. Computer, 7(9):18–19, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Epstein:1974:DMV

[712] George Epstein, Gideon Frieder, and David C. Rine. The development of multiple-valued logic as related to computer science. Computer, 7(9):20– 32, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDe

[713] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (9):33, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vranesic:1974:EAM

[714] Z. G. Vranesic and K. C. Smith. Engineering aspects of multi-valued logic systems. Computer, 7(9):34– 41, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Abraham:1974:VRM

[715] George Abraham. Variable radix multistable integrated circuits. Computer, 7(9):42–59, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davis:1974:USN

[716] Robert L. Davis. Uniform shift networks. Computer, 7(9):60–71, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Race:1974:FSE

[717] John Race. Fiction: Scorched earth policy. Computer, 7(9):72–73, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPi

[718] Anonymous. New products. Computer, 7(9):74–77, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAf

[719] Anonymous. Recent IC announcements. Computer, 7(9):78, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAi

[720] Anonymous. New applications. Computer, 7(9):79–82, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAe

[721] Anonymous. Classified ads. Computer, 7(9):82, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ri

[722] Anonymous. Repository. Computer, 7 (9):83–85, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAf

[723] Anonymous. Short courses [advertisement]. Computer, 7(9):86–87, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPh

[724] Anonymous. Product profile. Computer, 7(9):88–90, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAf

[725] Anonymous. Communications [advertisement]. Computer, 7(9):91–93, September 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCe

[726] Anonymous. [back cover]. Computer, 7 (9):c4, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCi

[727] Anonymous. [front cover]. Computer, 7(9):c1, September 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCe

[728] Anonymous. [inside front cover]. Computer, 7(9):c2, September 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cv

[729] Anonymous. Contents. Computer, 7 (10):1, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FMb

[730] Anonymous. [front matter]. Computer, 7(10):1–3, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Pe

[731] Anonymous. From the President. Computer, 7(10):2–3, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cw

[732] Anonymous. Calendar. Computer, 7 (10):4–5, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPm

[733] Anonymous. Call for papers. Computer, 7(10):5, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ug

[734] Anonymous. Update. Computer, 7 (10):6–15, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:HCA

[735] Anonymous. Haynes’ cookbook [advertisement]. Computer, 7(10):16– 17, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liebowitz:1974:RTC

[736] Burt H. Liebowitz. Real time computer applications in gambling. Computer, 7(10):18–19, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levine:1974:CR

[737] David A. Levine. Computers at the racetrack. Computer, 7(10):20–28, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDf

[738] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (10):29, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dunik:1974:LST

[739] Stephen L. Dunik. Large scale off-track betting systems. Computer, 7(10): 30–38, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stigall:1974:RDG

[740] Paul D. Stigall and Ömür Tasar. A review of directed graphs as applied to computors. Computer, 7(10):39– 47, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPM

[741] Anonymous. New publication on mathematical software [advertisement].

Computer, 7(10):48, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jarvis:1974:SFI

[742] R. A. Jarvis. Special feature: an interactive minicomputer laboratory for graphics, image processing, and pattern recognition. Computer, 7(10): 49–60, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1974:LAW

[743] R. R. Brown. 1974 Lake Arrowhead Workshop on structured programming. Computer, 7(10):61–63, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bacon:1974:MSW

[744] G. C. Bacon. Mass storage workshop report. Computer, 7(10):64– 65, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPj

[745] Anonymous. New products. Computer, 7(10):66–69, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAg

[746] Anonymous. Recent IC announcements. Computer, 7(10):70, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAj

[747] Anonymous. New applications. Computer, 7(10):71–72, October 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rj

[748] Anonymous. Repository. Computer, 7(10):73–76, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAg

[749] Anonymous. Classified ads. Computer, 7(10):76–78, October 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPi

[750] Anonymous. Product profile. Computer, 7(10):79–81, October 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCf

[751] Anonymous. [back cover]. Computer, 7(10):c4, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCj

[752] Anonymous. [front cover]. Computer, 7(10):c1, October 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cx

[753] Anonymous. Contents. Computer, 7 (11):1, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cy

[754] Anonymous. Calendar. Computer, 7 (11):2–3, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPn

[755] Anonymous. Call for papers. Computer, 7(11):3, November 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Uh

[756] Anonymous. Update. Computer, 7(11): 4–10, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NOG

[757] Anonymous. New officers and governing board members elected. Computer, 7(11):11–13, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1974:PRA

[758] W. W. Chu and H. Opderbeck. Performance of replacement algorithms with different page sizes. Computer, 7(11): 14–23, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Flores:1974:LCI

[759] Ivan Flores. Lookahead control in the IBM System 370 model 165. Computer, 7(11):24–38, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Desautels:1974:CFC

[760] E. J. Desautels. On computing facilities for computer science. Computer, 7(11): 39–49, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bethke:1974:TWW

[761] Frederick J. Bethke. Technical writing: Weaving the silk purse. Computer, 7 (11):50–51, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1974:WRM

[762] Louise Jones, Richard Kleir, Alan Salisbury, John Tartar, and Larry Wear. Workshop report: Micro 7 highlights. Computer, 7(11):52–54, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAW

[763] Anonymous. Computer applications workshop II: On-line computer applications. Computer, 7(11):55–57, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPk

[764] Anonymous. New products. Computer, 7(11):58–62, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAk

[765] Anonymous. New applications. Computer, 7(11):63–65, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rk

[766] Anonymous. Repository. Computer, 7 (11):66–70, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAh

[767] Anonymous. Classified ads. Computer, 7(11):70, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPj

[768] Anonymous. Product profile. Computer, 7(11):71–73, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:AIb

[769] Anonymous. Advertisers index. Computer, 7(11):73–76, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BSA

[770] Anonymous. The book shelf [advertisement]. Computer, 7(11):77, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCg

[771] Anonymous. [back cover]. Computer, 7 (11):c4, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCk

[772] Anonymous. [front cover]. Computer, 7 (11):c1, November 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCf

[773] Anonymous. [inside front cover]. Computer, 7(11):c2, November 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cz

[774] Anonymous. Contents. Computer, 7 (12):1, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Pf

[775] Anonymous. From the President. Computer, 7(12):2–3, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Cba

[776] Anonymous. Calendar. Computer, 7 (12):4, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CPo

[777] Anonymous. Call for papers. Computer, 7(12):5, December 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Ui

[778] Anonymous. Update. Computer, 7(12): 6–16, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PRDg

[779] Anonymous. Pattern recognition data bases [advertisement]. Computer, 7 (12):17, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:WDW

[780] Anonymous. Why do we need computer hardware description languages? Computer, 7(12):18–26, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:ICH

[781] Anonymous. Introducing computer hardware description languages [advertisement]. Computer, 7(12):27, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:1974:IA

[782] Frederick J. Hill. Introducing AHPL. Computer, 7(12):28–30, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1974:IC

[783] Yaohan Chu. Introducing CDL. Computer, 7(12):31–33, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dietmeyer:1974:ID

[784] D. L. Dietmeyer. Introducing DDL. Computer, 7(12):34–38, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siewiorek:1974:II

[785] Dan Siewiorek. Introducing ISP. Computer, 7(12):39–41, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siewiorek:1974:IP

[786] Dan Siewiorek. Introducing PMS. Computer, 7(12):42–44, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Su:1974:SCH

[787] Stephen Y. H. Su. A survey of computer hardware description languages in the U.S.A. Computer, 7(12):45– 51, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCH

[788] Anonymous. A survey of computer hardware description languages abroad [breaker page]. Computer, 7(12): 52, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Vaucher:1974:HDL

[789] Jean Vaucher. Hardware description languages in Canada. Computer, 7(12): 53–54, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mermet:1974:HDL

[790] Jean Mermet. Hardware description languages in France. Computer, 7(12): 55–56, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Piloty:1974:HDL

[791] Robert Piloty. Hardware description languages in the Federal Republic of Germany. Computer, 7(12):57– 59, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Marcoz:1974:HDL

[792] F. Marcoz and O. Tedone. Hardware description languages in Italy. Computer, 7(12):60–61, December 1974.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Watanabe:1974:HDL

[793] Hitoshi Watanabe and Kiichi Fujino. Hardware description languages in Japan. Computer, 7(12):62–63, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dervisoglu:1974:HDL

[794] Bulent Dervisoglu. Hardware description languages in Great Britain. Computer, 7(12):64–66, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Magnuson:1974:CWH

[795] Waldo G. Magnuson. CHDL workshop held in Germany. Computer, 7 (12):67, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NPl

[796] Anonymous. New products. Computer, 7(12):68–72, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:RIAh

[797] Anonymous. Recent IC announcements. Computer, 7(12):72, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:NAl

[798] Anonymous. New applications. Computer, 7(12):73–75, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:Rl

[799] Anonymous. Repository. Computer, 7 (12):76–80, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:PPk

[800] Anonymous. Product profile. Computer, 7(12):81–82, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:AIc

[801] Anonymous. Advertisers index. Computer, 7(12):82, December 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:CAi

[802] Anonymous. Classified ads. Computer, 7(12):82, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:SCAg

[803] Anonymous. Short courses [advertisement]. Computer, 7(12):83–84, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BS

[804] Anonymous. The book shelf. Computer, 7(12):85, December 1974. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:BCh

[805] Anonymous. [back cover]. Computer, 7 (12):c4, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:FCl

[806] Anonymous. [front cover]. Computer, 7 (12):c1, December 1974. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1974:IFCg

[807] Anonymous. [inside front cover]. Computer, 7(12):c2, December 1974. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCa

[808] Anonymous. Table of contents. Computer, 8(1):1, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Pa

[809] Anonymous. From the President. Computer, 8(1):2–3, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ca

[810] Anonymous. Calendar. Computer, 8 (1):4–70, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPa

[811] Anonymous. Call for papers. Computer, 8(1):5, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPb

[812] Anonymous. Call for papers. Computer, 8(1):6–7, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICa

[813] Anonymous. Update: information for the computer systems design professional. Computer, 8(1):8–17, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NPa

[814] Anonymous. ’75 NCC preview. Computer, 8(1):13, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gardner:1975:PCM

[815] R. M. Gardner. The place of computers in medicine. Computer, 8(1): 18–19, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:HMDa

[816] Anonymous. HP 3000CX mini datacenters. Computer, 8(1):20–21, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Drazen:1975:RCP

[817] E. Drazen, A. Wechsler, K. Wiig, and A. D. Little. Requirements for computerized patient monitoring systems. Computer, 8(1):22–27, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Giebink:1975:CSA

[818] G. A. Giebink, L. L. Hurst, C. L. Foreman, and D. Elser. Current status of ambulatory health care computer applications. Computer, 8(1): 28–32, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:LYMa

[819] Anonymous. Last year more than 20 million credit cards were lost or stolen! [advertisement]. Computer, 8 (1):33, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pryor:1975:HCS

[820] T. A. Pryor, J. D. Morgan, S. J. Clark, W. A. Miller, and H. R. Warner. Help a computer system for medical decision making. Computer, 8(1): 34–38, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gardner:1975:CBG

[821] R. M. Gardner, G. H. Cannon, A. H. Morris, K. R. Olsen, and W. G. Price. Computerized blood gas interpretation and reporting system. Computer, 8(1): 39–45, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Starmer:1975:CBA

[822] C. F. Starmer and R. A. Rosati. Computer-based aid to managing patients with chronic illness. Computer, 8 (1):46–50, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Feth:1975:WRI

[823] G. C. Feth. Workshop report: Impact of emerging technology. Computer, 8 (1):51–54, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPc

[824] Anonymous. Call for papers. Computer, 8(1):55, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPa

[825] D. A. Michalopoulos. New products. Computer, 8(1):56–60, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAa

[826] Anonymous. Recent IC announcements. Computer, 8(1):61, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAa

[827] D. A. Michalopoulos. New applications. Computer, 8(1):62–64, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:STT

[828] Anonymous. Stay on top of things [advertisement]. Computer, 8(1):65, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAa

[829] Anonymous. Repository [advertisement]. Computer, 8(1):66–68, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPa

[830] Anonymous. Product profile. Computer, 8(1):71–72, January 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:AI

[831] Anonymous. Advertisers index 1974. Computer, 8(1):73, January 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDa

[832] Anonymous. Pattern recognition data bases. Computer, 8(1):74, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRa

[833] Anonymous. Book reviews. Computer, 8(1):75–76, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ba

[834] Anonymous. The bookshelf. Computer, 8(1):77, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:USCa

[835] Anonymous. University of Southern California offers four outstanding continuing engineering education short courses [advertisement]. Computer, 8 (1):77, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cb

[836] Anonymous. Computer. Computer, 8(1):c1, January 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:MCN

[837] Anonymous. Modern computer networks. Computer, 8(1):c2, January 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCb

[838] Anonymous. Table of contents. Computer, 8(2):1, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cc

[839] Anonymous. Calendar. Computer, 8 (2):2–79, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPd

[840] Anonymous. Call for papers. Computer, 8(2):4–80, February 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICb

[841] Anonymous. Update: information for the computer systems design professional. Computer, 8(2):6–11, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ICSa

[842] Anonymous. 1975 IEEE Computer Society roster. Computer, 8(2):12– 13, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:OCa

[843] Anonymous. The open channel. Computer, 8(2):14–16, February 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WKO

[844] Anonymous. We know our GGG substrates grow good garnet films because we grow good garnet films on our GGG substrates [advertisement]. Computer, 8(2):17, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPe

[845] Anonymous. Call for papers. Computer, 8(2):18–19, February 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wensley:1975:ED

[846] J. H. Wensley. Electronic disks. Computer, 8(2):20–21, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:HMDb

[847] Anonymous. HP 3000CX Mini Data-Centers. Computer, 8(2):22–23, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Martin:1975:ED

[848] R. R. Martin and H. D. Frankel. Electronic disks in the 1980’s. Computer, 8 (2):24–30, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:AACa

[849] Anonymous. The age of the affordable computer [advertisement]. Computer, 8(2):31, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kelly:1975:DEE

[850] J. Kelly. The development of an experimental electron-beam-addressed memory module. Computer, 8(2):32– 42, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NPb

[851] Anonymous. ’75 NCC preview. Computer, 8(2):43, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wensley:1975:IED

[852] J. H. Wensley. The impact of electronic disks on system architecture. Computer, 8(2):44–48, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CWSa

[853] Anonymous. The computer will serve you better if it is able to look out for itself! Computer, 8(2):49, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harper:1975:SFB

[854] W. L. Harper. Special feature: Building EDP success by standing on shoulders. Computer, 8(2):50–56, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDb

[855] Anonymous. Pattern recognition data bases. Computer, 8(2):57, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Matcovitch:1975:WRS

[856] T. J. Matcovitch. Workshop report: Split Rock Workshop on Computer Packaging. Computer, 8(2):58– 59, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPb

[857] D. A. Michalopoulos. New products. Computer, 8(2):60–66, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAb

[858] Anonymous. Recent IC announcements. Computer, 8(2):67, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:IOC

[859] Anonymous. International Optical Computing Conference 1975. Computer, 8(2):68–69, February 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAb

[860] D. A. Michalopoulos. New applications. Computer, 8(2):70–71, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:SE

[861] Anonymous. Software engineering. Computer, 8(2):72, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPb

[862] Anonymous. Product profile. Computer, 8(2):73–74, February 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAb

[863] Anonymous. Repository [advertisement]. Computer, 8(2):75–77, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPf

[864] Anonymous. Call for papers. Computer, 8(2):78, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:LYMb

[865] Anonymous. Last year more than 20 million credit cards were lost or stolen! [advertisement]. Computer, 8 (2):81, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Aa

[866] Anonymous. Announcing. Computer, 8(2):82, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRb

[867] Anonymous. Book reviews. Computer, 8(2):83–84, February 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:USCb

[868] Anonymous. University of Southern California offers five outstanding continuing engineering education short courses [advertisement]. Computer, 8 (2):84, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cd

[869] Anonymous. Computer. Computer, 8 (2):c1, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UJCa

[870] Anonymous. The 2nd USA–Japan Computer Conference. Computer, 8 (2):c2, February 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCc

[871] Anonymous. Table of contents. Computer, 8(3):1, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Pb

[872] Anonymous. From the President. Computer, 8(3):2–3, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CLP

[873] Anonymous. Catch a logic pulse [advertisement]. Computer, 8(3):4–78, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPg

[874] Anonymous. Call for papers. Computer, 8(3):5–79, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICc

[875] Anonymous. Update: information for the computer systems design professional. Computer, 8(3):6–16, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:DEMa

[876] Anonymous. Does easier mean better? [advertisement]. Computer, 8 (3):9, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGa

[877] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8(3):10, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NPc

[878] Anonymous. ’75 NCC preview. Computer, 8(3):14, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ACCa

[879] Anonymous. Area Committee Chairman, Chapter Chairman Directory. Computer, 8(3):17–79, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:OCb

[880] Anonymous. The open channel. Computer, 8(3):19–20, March 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WCDa

[881] Anonymous. We’ve cut data recording down to size [advertisement]. Computer, 8(3):21, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ab

[882] Anonymous. Announcing. Computer, 8(3):22, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Szygenda:1975:DSS

[883] S. A. Szygenda. Digital systems simulation. Computer, 8(3):23, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Szygenda:1975:DLS

[884] S. A. Szygenda and E. W. Thompson. Digital logic simulation in a time-based, table-driven environment. Computer, 8(3):24–36, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:AACb

[885] Anonymous. The age of the affordable computer [advertisement]. Computer, 8(3):37, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thompson:1975:DLS

[886] E. W. Thompson and S. A. Szygenda. Digital logic simulation in a time-based, table-driven environment. Computer, 8(3):38–49, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:HMDc

[887] Anonymous. HP 3000CX Mini Data-Centers. Computer, 8(3):50–51, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chang:1975:DTS

[888] H. Y. Chang and S. G. Chappell. Deductive techniques for simulating logic circuits. Computer, 8(3):52–59, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CTW

[889] Anonymous. Check the texts we’ve programmed for your computer science needs. Computer, 8(3):60, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPc

[890] D. A. Michalopoulos. New products. Computer, 8(3):61–64, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:FIS

[891] Anonymous. Fifth International Symposium on Multiple-Valued Logic. Computer, 8(3):65, March 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAc

[892] Anonymous. Recent IC announcements. Computer, 8(3):66, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAc

[893] D. A. Michalopoulos. New application. Computer, 8(3):67–69, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAc

[894] Anonymous. Repository [advertisement]. Computer, 8(3):70–72, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:LYMc

[895] Anonymous. Last year more than 20 million credit cards were lost or stolen! [advertisement]. Computer, 8(3):73, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPc

[896] Anonymous. Product profile. Computer, 8(3):74, March 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:IIP

[897] Anonymous. IEEE Intercon proudly presents un-flation. Computer, 8 (3):76, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:SRAa

[898] Anonymous. Science Research Associates, Inc. [advertisement]. Computer, 8(3):80, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRc

[899] Anonymous. Book reviews. Computer, 8(3):81–83, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CAa

[900] Anonymous. Classified ads. Computer, 8(3):84, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:MC

[901] Anonymous. Microcomputer congress. Computer, 8(3):84, March 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ce

[902] Anonymous. Computer. Computer, 8(3):c1, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UJCb

[903] Anonymous. The 2nd USA-Japan Computer Conference. Computer, 8 (3):c2, March 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCd

[904] Anonymous. Table of contents. Computer, 8(4):1, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ICSb

[905] Anonymous. 1975 IEEE Computer Society roster. Computer, 8(4):2–65, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPh

[906] Anonymous. Call for papers. Computer, 8(4):4–5, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ODT

[907] Anonymous. A one-day tutorial on microprogramming. Computer, 8(4):6, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICd

[908] Anonymous. Update: information for the computer systems design professional. Computer, 8(4):7–14, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGb

[909] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8(4):9, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NPd

[910] Anonymous. ’75 NCC preview. Computer, 8(4):13, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:DEMb

[911] Anonymous. Does easier mean better? [advertisement]. Computer, 8 (4):15, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ISF

[912] Anonymous. 1975 International Symposium on Fault-Tolerant Computing. Computer, 8(4):16–17, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wall:1975:DAG

[913] H. M. Wall. Design automation [guest editor’s introduction]. Computer, 8 (4):18–19, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WTK

[914] Anonymous. What is there to know about computer networks? Computer, 8(4):20–21, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDc

[915] Anonymous. Pattern recognition data bases. Computer, 8(4):22, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garrocq:1975:ISF

[916] C. A. Garrocq and M. J. Hurley. The IPAD system: a future management/ engineering/design environment. Computer, 8(4):23–33, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:HMDd

[917] Anonymous. HP 3000CX Mini Data-Centers. Computer, 8(4):34–35, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Harlow:1975:AIE

[918] C. A. Harlow, S. E. Henderson, D. A. Rayfield, R. J. Johnston, and S. J. Dwyer. Automated inspection of electronic assemblies. Computer, 8(4): 36–45, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woodyard:1975:CAS

[919] A. H. Woodyard. Computer-aided ship design at MarAd. Computer, 8 (4):46–51, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPd

[920] D. A. Michalopoulos. New products. Computer, 8(4):52–57, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAd

[921] Anonymous. Recent IC announcements. Computer, 8(4):58, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAd

[922] D. A. Michalopoulos. New applications. Computer, 8(4):59–61, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPd

[923] Anonymous. Product profile. Computer, 8(4):62, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRd

[924] Anonymous. Book reviews. Computer, 8(4):66, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAd

[925] Anonymous. Repository [advertisement]. Computer, 8(4):67–68, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Bb

[926] Anonymous. The bookshelf. Computer, 8(4):68, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WCDb

[927] Anonymous. We’ve cut data recording down to size [advertisement]. Computer, 8(4):68, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cf

[928] Anonymous. Computer. Computer, 8(4):c1, April 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PNA

[929] Anonymous. Plan now to attend the International Conference on Communications. Computer, 8(4):c2, April 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCe

[930] Anonymous. Table of contents. Computer, 8(5):1, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cg

[931] Anonymous. Calendar. Computer, 8(5):2–76, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ICSc

[932] Anonymous. 1975 IEEE Computer Society roster. Computer, 8(5):4–5, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGc

[933] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8(5):6, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICe

[934] Anonymous. Update: information for the computer systems design professional. Computer, 8(5):7–14, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:DEMc

[935] Anonymous. Does easier mean better? [advertisement]. Computer, 8 (5):11, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yeh:1975:SE

[936] R. T. Yeh. Software engineering. Computer, 8(5):15, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:M

[937] Anonymous. MITS. Computer, 8 (5):16, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ross:1975:SEP

[938] D. T. Ross, J. B. Goodenough, and C. A. Irvine. Software engineering: Process, principles, and goals. Computer, 8(5):17–27, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bratman:1975:SF

[939] H. Bratman and T. Court. The software factory. Computer, 8(5):28–37, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:HPM

[940] Anonymous. The Hewlett-Packard 3000 is a minicomputer? Computer, 8(5):38–39, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chandy:1975:SAM

[941] K. M. Chandy. A survey of analytic models of rollback and recovery strategies. Computer, 8(5):40–47, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NIP

[942] Anonymous. New IFIP publications [advertisement]. Computer, 8(5):48, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chin:1975:SFA

[943] Y. H. Chin, J. W. Jou, W. H. Peng, and C. C. Yang. Special feature: an automated Chinese telephone directory. Computer, 8(5):49–54, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ac

[944] Anonymous. [advertisement]. Computer, 8(5):55, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eklund:1975:TRW

[945] M. H. Eklund. Technology in the real world. Computer, 8(5):56–57, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPe

[946] D. A. Michalopoulos. New products. Computer, 8(5):58–64, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WCDc

[947] Anonymous. We’ve cut data recording down to size [advertisement]. Computer, 8(5):65, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAe

[948] Anonymous. Recent IC announcements. Computer, 8(5):66, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAe

[949] D. A. Michalopoulos. New applications. Computer, 8(5):67–68, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDd

[950] Anonymous. Pattern recognition data bases. Computer, 8(5):69, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPe

[951] Anonymous. Product profile. Computer, 8(5):70, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:MPK

[952] Anonymous. The money people have to know what’s happening [advertisement]. Computer, 8(5):71, May 1975.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAe

[953] Anonymous. Repository [advertisement]. Computer, 8(5):72–75, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CWSb

[954] Anonymous. The computer will serve you better if it is able to look out for itself! Computer, 8(5):76, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ch

[955] Anonymous. Computer. Computer, 8(5):c1, May 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:FDC

[956] Anonymous. Fourth Data Communications Symposium [advertisement]. Computer, 8(5):c2, May 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCf

[957] Anonymous. Table of contents. Computer, 8(6):1, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Pc

[958] Anonymous. From the President. Computer, 8(6):2–3, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:DEMd

[959] Anonymous. Does easier mean better? [advertisement]. Computer, 8(6):4, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ci

[960] Anonymous. Calendar. Computer, 8 (6):5–79, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPi

[961] Anonymous. Call for papers. Computer, 8(6):8–79, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICf

[962] Anonymous. Update: information for the computer systems design professional. Computer, 8(6):9–19, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ross:1975:HLA

[963] W. F. Ross. Highlights of the 1974 Lake Arrowhead Workshop. Computer, 8(6):20–22, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Naughton:1975:SSP

[964] J. Naughton. Session I structured programming: Concepts and definitions. Computer, 8(6):23–24, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McGowan:1975:SPR

[965] C. McGowan. Structured programming: a review of some practical concepts. Computer, 8(6):25–30, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Horowitz:1975:FCI

[966] E. Horowitz. Fortran can it be structured — should it be? Computer, 8 (6):30–37, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1975:SIS

[967] B. Boehm. Session II structured programming: a quantitative assessment. Computer, 8(6):38–40, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holmes:1975:SPM

[968] C. E. Holmes. Structured programming at McAuto. Computer, 8(6): 41–43, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Katkus:1975:ASP

[969] G. R. Katkus. Applying structured programming to command, control, and communication software development. Computer, 8(6):43–47, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Romanos:1975:ISC

[970] J. P. Romanos. An implementation of structured code techniques on a real-time system. Computer, 8(6): 48–49, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McHenry:1975:MPI

[971] R. McHenry. Measuring programming improvement at IBM-FSD. Computer, 8(6):49, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gordon:1975:EAS

[972] E. K. Gordon. Experienced and accomplishments with structured programming. Computer, 8(6):50–53, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boehm:1975:CR

[973] B. W. Boehm. Concluding remarks. Computer, 8(6):53–54, June 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1975:SIS

[974] R. R. Brown. Session III structured programming: Problems, approaches, and techniques. Computer, 8(6):55, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Town:1975:SPA

[975] J. R. Town. Structured programming: Agony and/or ecstacy. Computer, 8 (6):56–57, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Braken:1975:EML

[976] J. W. Braken. Experience with modulelevel specification methods. Computer, 8(6):58–61, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

deBalbine:1975:SII

[977] G. de Balbine. Session IV impact of structured programming on evolving technologies and related programming disciplines. Computer, 8(6):62, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zahn:1975:NIP

[978] C. T. Zahn. The need for improved programming language. Computer, 8 (6):63–64, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

deBalbine:1975:SET

[979] G. de Balbine. The structuring engine: a transition tool. Computer, 8 (6):64–65, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Robinson:1975:FDS

[980] L. Robinson. A formal design for software. Computer, 8(6):66, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bigelow:1975:SCS

[981] R. Bigelow. Structured code via stack machine. Computer, 8(6):67, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Towaley:1975:SPA

[982] J. Towaley. Structured programming applied to equipment testing. Computer, 8(6):68, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Faber:1975:SDD

[983] D. J. Faber. Software development for distributed systems. Computer, 8 (6):68–69, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Liddle:1975:NLA

[984] A. S. Liddle. The need for language advances. Computer, 8(6):69, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Prywes:1975:PFN

[985] N. Prywes. Preparing for future needs. Computer, 8(6):70–72, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

deBalbine:1975:C

[986] G. de Balbine. Conclusions. Computer, 8(6):72–73, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkins:1975:IRR

[987] D. E. Atkins. Introduction to the role of redundancy in computer arithmetic. Computer, 8(6):74–77, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CAb

[988] Anonymous. Classified ads. Computer, 8(6):78, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPf

[989] Anonymous. Product profile. Computer, 8(6):78, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPf

[990] D. A. Michalopoulos. New products. Computer, 8(6):80–84, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAf

[991] Anonymous. Recent IC announcements. Computer, 8(6):85, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAf

[992] D. A. Michalopoulos. New applications. Computer, 8(6):86–87, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAf

[993] Anonymous. Repository [advertisement]. Computer, 8(6):88–91, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Bc

[994] Anonymous. The bookshelf. Computer, 8(6):92, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRe

[995] Anonymous. Book reviews. Computer, 8(6):92, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CSH

[996] Anonymous. Computer spoken here [advertisement]. Computer, 8(6):92, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cj

[997] Anonymous. Computer. Computer, 8(6):c1, June 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGd

[998] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8(6):c2, June 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCg

[999] Anonymous. Table of contents. Computer, 8(7):1, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ICSd

[1000] Anonymous. 1975 IEEE Computer Society roster. Computer, 8(7):2–92, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPj

[1001] Anonymous. Call for papers. Computer, 8(7):4–73, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:APa

[1002] Anonymous. Advance program. Computer, 8(7):5–11, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICg

[1003] Anonymous. Update: information for the computer systems design professional. Computer, 8(7):12–19, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ICSe

[1004] Anonymous. IEEE Computer Society bylaws. Computer, 8(7):20–24, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:SWL

[1005] Anonymous. SPIE has a way of looking at … [advertisement]. Computer, 8(7):25, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kempner:1975:CC

[1006] K. M. Kempner. Computers in cardiology. Computer, 8(7):26–28, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDe

[1007] Anonymous. Pattern recognition data bases. Computer, 8(7):29, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Thomas:1975:CMPa

[1008] Lewis J. Thomas, Jr., G. James Blaine, V. W. Gerth, Jr., and Ronald W. Hagen. Continuous monitoring of physiologic variables with a dedicated minicomputer. Computer, 8(7):30–35, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Starmer:1975:PAR

[1009] C. Frank Starmer and William M. Smith. Problems in acquisition and representation of coronary arterial trees. Computer, 8(7):36–41, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Simoons:1975:LAEa

[1010] Maarten L. Simoons, Ed Smallenburg, Cees Zeelenberg, and Willem A. H. Engelse. On-line analysis of exercise electrocardiograms. Computer, 8 (7):42–45, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ISC

[1011] Anonymous. 1975 International Symposium on Computer Hardware Description Languages and their Applications. Computer, 8(7):46–47, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fitzgerald:1975:SCP

[1012] John W. Fitzgerald, Robert J. Clappier, and Donald C. Harrison. Srtre computer processing of ambulatory electrocardiograms. Computer, 8(7): 48–54, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Heintzen:1975:AVAa

[1013] Paul H. Heintzen, Ruediger Brennecke, Joachim H. Buersch, Peter Lange, Vitus Malerczyk, Klaus Moldenhauer, and Dietrich Onnasch. Automated video-angiocardiographic image analysis. Computer, 8(7):55–64, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raeside:1975:BSM

[1014] David E. Raeside. Bayesian statistics: Mathematical tools for computers*. Computer, 8(7):65–73, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Parrish:1975:NDM

[1015] E. A. Parrish, G. Nagy, Y. T. Chien, and A. Klinger. New directions in machine-pattern analysis. Computer, 8(7):74–76, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stucki:1975:CAP

[1016] L. G. Stucki. Currently available program testing tools: Technology and experience. Computer, 8(7):77–78, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:SAS

[1017] Anonymous. Second annual symposium on computer architecture. Computer, 8(7):79–80, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPg

[1018] D. A. Michalopoulos. New products. Computer, 8(7):81–84, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAg

[1019] Anonymous. Recent IC announcements. Computer, 8(7):85, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAg

[1020] D. A. Michalopoulos. New applications. Computer, 8(7):86–88, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAg

[1021] Anonymous. Repository [advertisement]. Computer, 8(7):89–91, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPg

[1022] Anonymous. Product profile. Computer, 8(7):91, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGe

[1023] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8(7):92, July 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ck

[1024] Anonymous. Computer. Computer, 8(7):c1, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPk

[1025] Anonymous. Call for papers. Computer, 8(7):c2, July 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCh

[1026] Anonymous. Table of contents. Computer, 8(8):1, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Pd

[1027] Anonymous. From the President. Computer, 8(8):2–3, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cl

[1028] Anonymous. Calendar. Computer, 8 (8):4–74, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPl

[1029] Anonymous. Call for papers. Computer, 8(8):6–75, August 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:APb

[1030] Anonymous. Advance program. Computer, 8(8):7–13, August 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDf

[1031] Anonymous. Pattern recognition data bases. Computer, 8(8):14, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICh

[1032] Anonymous. Update: information for the computer systems design professional. Computer, 8(8):15–21, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:LE

[1033] Anonymous. Letters to the editors. Computer, 8(8):22, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Galey:1975:MBB

[1034] J. M. Galey. Microprogramming: The bridge between hardware and software. Computer, 8(8):23, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mallach:1975:EA

[1035] E. G. Mallach. Emulator architecture. Computer, 8(8):24–32, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1975:SCW

[1036] L. H. Jones. A survey of current work in microprogramming. Computer, 8 (8):33–38, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPm

[1037] Anonymous. Call for papers. Computer, 8(8):39, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mallett:1975:CIH

[1038] P. W. Mallett and T. G. Lewis. Considerations for implementing a high level microprogramming language translation system. Computer, 8(8):40–52, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EAW

[1039] Anonymous. Eighth Annual Workshop on Microprogramming. Computer, 8 (8):53, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Newell:1975:SHC

[1040] J. C. Newell, D. O. Dickman, and M. A. Winkler. Synthesis of holograms from computer-generated output. Computer, 8(8):54–56, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPh

[1041] D. A. Michalopoulos. New products.

Computer, 8(8):57–63, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAh

[1042] Anonymous. Recent IC announcements. Computer, 8(8):64, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAh

[1043] D. A. Michalopoulos. New applications. Computer, 8(8):65–66, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAh

[1044] Anonymous. Repository [advertisement]. Computer, 8(8):67–72, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPh

[1045] Anonymous. Product profile. Computer, 8(8):73, August 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Bd

[1046] Anonymous. The bookshelf. Computer, 8(8):76, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:END

[1047] Anonymous. Electronics in the next 1000 days [advertisement]. Computer, 8(8):76, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cm

[1048] Anonymous. Computer. Computer, 8(8):c1, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGf

[1049] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8 (8):c2, August 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCi

[1050] Anonymous. Table of contents. Computer, 8(9):1, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:MDS

[1051] Anonymous. Microcomputer design is a snap [advertisement]. Computer, 8 (9):2–3, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:FR

[1052] Anonymous. Financial report for 1974. Computer, 8(9):4–6, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:FTC

[1053] Anonymous. Fourth Texas Conference on Computing Systems. Computer, 8 (9):7, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICi

[1054] Anonymous. Update: information for the computer systems design professional. Computer, 8(9):8–13, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:EWGg

[1055] Anonymous. Everybody wins … in the Great IEEE Computer Society membership race. Computer, 8(9): 12, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Busch:1975:PTC

[1056] D. F. Busch. Printer technology: Competition for the Ribbon Bangers. Computer, 8(9):14–15, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lennemann:1975:ILP

[1057] E. Lennemann and W. Sakmann. Impact line printing: Improvement of a proven technology. Computer, 8(9): 16–27, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tschinkel:1975:NCI

[1058] J. Tschinkel. A new concept in impact printing. Computer, 8(9):28– 36, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDg

[1059] Anonymous. Pattern recognition data bases. Computer, 8(9):37, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:MSL

[1060] Anonymous. MOS 6502 the second of a low cost high performance microprocessor family. Computer, 8(9): 38–39, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mitchell:1975:XCP

[1061] J. N. Mitchell. The Xerox 1200 Computer Printing System. Computer, 8 (9):40–48, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Borelli:1975:NPP

[1062] R. F. Borelli, R. B. Bayless, and E. R. Truax. A nonimpact page printing system. Computer, 8(9):49–57, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPi

[1063] D. A. Michalopoulos. New products. Computer, 8(9):58–63, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAi

[1064] Anonymous. Recent IC announcements. Computer, 8(9):64, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAi

[1065] D. A. Michalopoulos. New applications. Computer, 8(9):65–66, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAi

[1066] Anonymous. Repository [advertisement]. Computer, 8(9):67–70, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cn

[1067] Anonymous. Calendar. Computer, 8 (9):73–75, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPi

[1068] Anonymous. Product profile. Computer, 8(9):75, September 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CAc

[1069] Anonymous. Classified ads. Computer, 8(9):76–78, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:OPW

[1070] Anonymous. One plot is worth thirty pages of printout [advertisement]. Computer, 8(9):78, September 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Co

[1071] Anonymous. Computer. Computer, 8 (9):c1, September 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPn

[1072] Anonymous. Call for papers. Computer, 8(9):c2, September 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCj

[1073] Anonymous. Table of contents. Computer, 8(10):1, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:IWL

[1074] Anonymous. Introducing … the world’s lowest cost computer system JOLT [advertisement]. Computer, 8 (10):2–3, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Pe

[1075] Anonymous. From the President. Computer, 8(10):4–5, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICj

[1076] Anonymous. Update: information for the computer systems design professional. Computer, 8(10):6–19, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Collins:1975:MDA

[1077] D. C. Collins. Microprocessor design and application. Computer, 8(10): 20–21, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TMA

[1078] Anonymous. Tutorial: Microprocessor application in multiple processor systems. Computer, 8(10):22–30, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CRSa

[1079] Anonymous. Capture! realtime software design challenge [advertisement]. Computer, 8(10):31, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Conley:1975:PMC

[1080] S. W. Conley. Portable microcomputer cross-assembler in basic. Computer, 8(10):32–42, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Korn:1975:PMS

[1081] G. A. Korn. A proposed method for simplified microcomputer programming*. Computer, 8(10):43–52, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gray:1975:MCT

[1082] M. T. Gray. Microprocessors in CRT terminal applications: Hardware/ software tradeoffs. Computer, 8(10): 53–59, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Laliotis:1975:MCD

[1083] T. A. Laliotis and T. D. Brumett.

A microprocessors-controlled DIC test system*. Computer, 8(10):60–67, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDh

[1084] Anonymous. Pattern recognition data bases. Computer, 8(10):68, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPj

[1085] D. A. Michalopoulos. New products. Computer, 8(10):69–74, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAj

[1086] Anonymous. Recent IC announcements. Computer, 8(10):75, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAj

[1087] D. A. Michalopoulos. New applications. Computer, 8(10):76–78, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAj

[1088] Anonymous. Repository [advertisement]. Computer, 8(10):81–88, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPo

[1089] Anonymous. Call for papers. Computer, 8(10):87, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cq

[1090] Anonymous. Calendar. Computer, 8 (10):89–92, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRf

[1091] Anonymous. Book review. Computer, 8(10):93, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPj

[1092] Anonymous. Product profile. Computer, 8(10):93, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Be

[1093] Anonymous. The bookshelf. Computer, 8(10):94, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:JAa

[1094] Anonymous. Japan? [advertisement].

Computer, 8(10):94, October 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cr

[1095] Anonymous. Computer. Computer, 8 (10):c1, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NNa

[1096] Anonymous. News & notes. Computer, 8(10):c2, October 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCk

[1097] Anonymous. Table of contents. Computer, 8(11):1, November 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICk

[1098] Anonymous. Update: information for the computer systems design professional. Computer, 8(11):2–11, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kehl:1975:LLM

[1099] T. H. Kehl, C. Moss, and L. Dunkel. LM2: a logic machine minicomputer. Computer, 8(11):12–22, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CRSb

[1100] Anonymous. Capture! Realtime software design challenge [advertisement]. Computer, 8(11):23, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:SRAb

[1101] Anonymous. Step right up [advertisement]. Computer, 8(11):24, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lea:1975:IPA

[1102] R. M. Lea. Information processing with an associative parallel processor. Computer, 8(11):25–32, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPCa

[1103] Anonymous. Call for papers on computer communications. Computer, 8 (11):33, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDi

[1104] Anonymous. Pattern recognition data bases. Computer, 8(11):34, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhandarkar:1975:TCS

[1105] D. P. Bhandarkar and J. E. Juliussen.

Tutorial: Computer system advantages of magnetic bubble memories. Computer, 8(11):35–40, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alexander:1975:SDC

[1106] W. G. Alexander and D. B. Wortman.

Static and dynamic characteristics of XPL programs. Computer, 8(11):41– 46, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NLD

[1107] Anonymous. ’76 NCC … landmarks in data processing. Computer, 8(11): 47–50, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCS

[1108] Anonymous. Tutorial: Computer system monitors. Computer, 8(11):51– 61, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:ACCb

[1109] Anonymous. Annual Conference of California Educational Data Processing Association (CEDPA) and California Educational Computing Consortium (CECC). Computer, 8(11): 62, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agajanian:1975:BSP

[1110] A. H. Agajanian. A bibliography on system performance evaluation. Computer, 8(11):63–74, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NYS

[1111] Anonymous. Now you see it, now you don’t [advertisement]. Computer, 8 (11):75, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cs

[1112] Anonymous. Calendar. Computer, 8 (11):76–79, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPp

[1113] Anonymous. Call for papers. Computer, 8(11):80–82, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPAa

[1114] Anonymous. Call for papers — ACM ’76 October 20–22, 1976…. Computer, 8(11):81, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPk

[1115] D. A. Michalopoulos. New products. Computer, 8(11):85–89, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAk

[1116] Anonymous. Recent IC announcements. Computer, 8(11):90, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAk

[1117] D. A. Michalopoulos. New applications. Computer, 8(11):91–92, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CAd

[1118] Anonymous. Classified ads. Computer, 8(11):93, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PPk

[1119] Anonymous. Product profile. Computer, 8(11):93, November 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:SC

[1120] Anonymous. Short courses. Computer, 8(11):94, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAk

[1121] Anonymous. Repository [advertisement]. Computer, 8(11):95–97, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Bf

[1122] Anonymous. The bookshelf. Computer, 8(11):98, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NNb

[1123] Anonymous. News & notes. Computer, 8(11):98, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Ct

[1124] Anonymous. Computer. Computer, 8 (11):c1, November 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:JAb

[1125] Anonymous. Japan? [advertisement]. Computer, 8(11):c2, November 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TCl

[1126] Anonymous. Table of contents. Computer, 8(12):1, December 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:PRDj

[1127] Anonymous. Pattern recognition data bases. Computer, 8(12):2, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPq

[1128] Anonymous. Call for papers. Computer, 8(12):3–6, December 1975. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:UICl

[1129] Anonymous. Update: information for the computer systems design professional. Computer, 8(12):7–12, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WU

[1130] Anonymous. Washington update.

Computer, 8(12):13–25, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CNY

[1131] Anonymous. Computers the next 5 years. Computer, 8(12):15–22, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Irwin:1975:CEG

[1132] D. Irwin and C. V. Ramamoorthy.

Computer education [guest editor’s introduction]. Computer, 8(12):26– 27, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mulder:1975:MCF

[1133] M. C. Mulder. Model curicula for four-year computer science and engineering programs: Bridging the tar pit. Computer, 8(12):28–33, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:FEP

[1134] Anonymous. The Federal Express Partsbank. It will change the way you think about distribution forever. [advertisement]. Computer, 8(12):

34, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sloan:1975:SEE

[1135] M. E. Sloan. Survey of electrical engineering and computer science departments in the U.S. Computer, 8(12): 35–42, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:MWR

[1136] Anonymous. Make waves! in real time software definition and design [employment opportunity]. Computer, 8(12): 43, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rossmann:1975:CSC

[1137] G. E. Rossmann, M. J. Flynn, S. H. Fuller, C. G. Bell, F. P. Brooks, and H. Hellerman. A course of study in computer hardware architecture. Computer, 8(12):44–63, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:TAS

[1138] Anonymous. Third Annual Symposium on Computer Architecture. Computer, 8(12):64–65, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barquin:1975:SCE

[1139] R. C. Barquin. A survey of computer education in Latin America. Computer, 8(12):66–72, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Grason:1975:THS

[1140] J. Grason and D. P. Siewiorek. Teaching with a hierarchically structured digital systems laboratory. Computer, 8 (12):73–81, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NPl

[1141] D. A. Michalopoulos. New products. Computer, 8(12):82–86, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RIAl

[1142] Anonymous. Recent IC announcements. Computer, 8(12):87, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1975:NAl

[1143] D. A. Michalopoulos. New applications. Computer, 8(12):88–90, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:RAl

[1144] Anonymous. Repository [advertisement]. Computer, 8(12):91–97, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CAe

[1145] Anonymous. Classified ads. Computer, 8(12):98, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cu

[1146] Anonymous. Calendar. Computer, 8 (12):99–102, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:CPr

[1147] Anonymous. Call for papers. Computer, 8(12):105–107, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:WEC

[1148] Anonymous. World Encyclopedia of Computer Education and Research [advertisement]. Computer, 8(12):108, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:BRg

[1149] Anonymous. Book reviews. Computer, 8(12):109–110, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Bg

[1150] Anonymous. The bookshelf. Computer, 8(12):110, December 1975. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:HA

[1151] Anonymous. The honeymoon is over [advertisement]. Computer, 8(12): 110, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:Cv

[1152] Anonymous. Computer. Computer, 8 (12):c1, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1975:NNc

[1153] Anonymous. News & notes. Computer, 8(12):c2, December 1975. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Casaglia:1976:SFN

[1154] G. Franco Casaglia. Special feature:

nanoprogramming vs. microprogramming. Computer, 9(1):54–58, January 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosenbluth:1976:DAA

[1155] William Rosenbluth. Design automation architecture and applications. Computer, 9(2):12–17, February 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Critchlow:1976:HSM

[1156] D. L. Critchlow. High speed MOS-FET circuits using advanced lithography. Computer, 9(2):31–37, February 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bagley:1976:SFM

[1157] John D. Bagley. Special feature:

microprogrammable virtual machines. Computer, 9(2):38–42, February 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1976:STO

[1158] Harold S. Stone. Special tutorial: the organization of electronic cyclic memories. Computer, 9(3):45–50, March 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Raymond:1976:UMI

[1159] Jacques Raymond and Dilip K. Banerji. Using a microprocessor in an intelligent graphics terminal. Computer, 9(4):18–25, April 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Panigrahi:1976:CMC

[1160] G. Panigrahi. Charge-coupled memories for computer systems. Computer, 9(4):33–42, April 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rege:1976:CPS

[1161] S. L. Rege. Cost, performance and size tradeoffs for different levels in a memory hierarchy. Computer, 9 (4):43–51, April 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chien:1976:IPR

[1162] Y. T. Chien. Interactive pattern recognition: techniques and systems. Computer, 9(5):11–25, May 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosenfeld:1976:PRS

[1163] Azriel Rosenfeld and Joan S. Weszka. Picture recognition and scene analysis. Computer, 9(5):28–38, May 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

White:1976:SRT

[1164] George M. White. Speech recognition: a tutorial review. Computer, 9 (5):40–53, May 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Preston:1976:CPB

[1165] Kendall Preston, Jr. Computer processing of biomedical images. Computer, 9(5):54–68, May 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Persoon:1976:MFP

[1166] Eric Persoon and K. S. Fu. Minicomputer facility for picture processing and pattern recognition research. Computer, 9(5):70–77, May 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Svobodova:1976:CSM

[1167] Liba Svobodova. Computer system measurability. Computer, 9(6):9–17, June 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Agrawala:1976:AWC

[1168] A. K. Agrawala, J. M. Mohr, and R. M.

Bryant. Approach to the workload characterization problem. Computer, 9(6):18–32, June 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williman:1976:ILM

[1169] A. O. Williman and H. J. Jelinek. Introduction to LSI microprocessor developments. Computer, 9(6):34–46, June 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bennetts:1976:RDT

[1170] R. G. Bennetts and R. V. Scott. Recent developments in the theory and practice of testable logic design. Computer, 9(6):47–63, June 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Landauer:1976:HDC

[1171] J. Paul Landauer. Hybrid digital/ analog computer systems. Computer, 9(7):15–24, July 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Howe:1976:QSC

[1172] R. M. Howe and J. Paul Landauer. Quantitative/method of speed comparison between analog/hybrid and digital computers. Computer, 9(7):31–36, July 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rubin:1976:HCF

[1173] Arthur I. Rubin and J. Bruce Mawson. Hybrid computation 1976 and its future. Computer, 9(7):37–46, July 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Graber:1976:NHC

[1174] G. F. Graber and E. J. Fadden. Next-generation hybrid computer system. Computer, 9(7):55–62, July 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foley:1976:TSG

[1175] James D. Foley. Tutorial on satellite graphics systems. Computer, 9 (8):14–21, August 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lucido:1976:SSC

[1176] Anthony P. Lucido. Software systems for computer graphics. Computer, 9 (8):23–32, August 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carlson:1976:GTR

[1177] Eric D. Carlson. Graphics terminal requirements for the 1970’s. Computer, 9 (8):37–45, August 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pooch:1976:CGI

[1178] Udo W. Pooch. Computer graphics, interactive techniques, and image processing 1970–1975: a bibliography. Computer, 9(8):46–64, August 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hehner:1976:CDM

[1179] Eric C. R. Hehner. Computer design to minimize memory requirements. Computer, 9(8):65–70, August 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fraser:1976:PSF

[1180] A. G. Fraser. The present status and future trends in computer/ communication technology. Computer, 9(9):10–19, September 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fulton:1976:MIS

[1181] David L. Fulton and Richard T.

Thomas. Minicomputer-based information system for a small business. Computer, 9(9):24–28, September 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bunt:1976:STO

[1182] Rick B. Bunt. Scheduling techniques for operating systems. Computer, 9 (10):10–18, October 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atwood:1976:COS

[1183] J. William Atwood. Concurrency in operating systems. Computer, 9(10): 18–26, October 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Doran:1976:VM

[1184] Robert Doran. Virtual memory. Computer, 9(10):27–37, October 1976. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dependahl:1976:FIC

[1185] Robert H. Dependahl, Jr. and Leon Presser. File input/output control logic. Computer, 9(10):38–42, October 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Levine:1976:SMR

[1186] Len Levine and Ware Myers. Semiconductor memory reliability with error detecting and correcting codes. Computer, 9(10):43–50, October 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spirn:1976:DSM

[1187] Jeffrey R. Spirn. Distance string models for program behavior. Computer, 9 (11):14–20, November 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Batson:1976:PBA

[1188] Alan P. Batson. Program behavior at the symbolic level. Computer, 9(11): 21–26, November 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1976:PBP

[1189] Wesley W. Chu and Holger Opderbeck. Program behavior and the page fault frequency replacement algorithm. Computer, 9(11):29–38, November 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ferrari:1976:IPB

[1190] Domenico Ferrari. Improvement of program behavior. Computer, 9(11):39– 47, November 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1976:KDC

[1191] Ware Myers. Key developments in computer technology: a survey. Computer, 9(11):48–75, November 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:Ca

[1192] Anonymous. Computer. Computer, 9 (12):c1, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:RHA

[1193] Anonymous. Repeat of highly acclaimed first public series software acquisition/development management [advertisement]. Computer, 9(12): c2, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:TC

[1194] Anonymous. Table of contents. Computer, 9(12):1, December 1976. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:P

[1195] Anonymous. From the President. Computer, 9(12):2, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:ICSa

[1196] Anonymous. 1976 IEEE Computer Society roster. Computer, 9(12):3– 4, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:SC

[1197] Anonymous. Spring Compcon77. Computer, 9(12):5–6, December 1976. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:UIC

[1198] Anonymous. Update: information for the computer systems design professional. Computer, 9(12):7–101, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hobbs:1976:LFG

[1199] L. C. Hobbs. A look at the future [guest editorial]. Computer, 9(12):9– 10, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tomash:1976:SNY

[1200] E. Tomash. SOS: The next 11111 years. Computer, 9(12):11–21, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Osborne:1976:PTP

[1201] T. E. Osborne. Personal thoughts on personal computing. Computer, 9(12): 22–24, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:SCB

[1202] Anonymous. Shopping for a computer at the ByteShop is almost as much fun as building one [advertisement]. Computer, 9(12):25, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Noyce:1976:RM

[1203] R. N. Noyce. From relays to mpu’s. Computer, 9(12):26–29, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:IYF

[1204] Anonymous. If you have faced tough software design challenges recently … you may be ready for these! [advertisement]. Computer, 9(12):30, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bremer:1976:HTY

[1205] J. W. Bremer. Hardware technology in the year 2001. Computer, 9(12): 31–36, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:CPa

[1206] Anonymous. Call for papers. Computer, 9(12):37–40, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:FY

[1207] Anonymous. The first 25 years. Computer, 9(12):41–53, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:HTN

[1208] Anonymous. Hardware then and now. Computer, 9(12):54–57, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:THC

[1209] Anonymous. Thoughts on the history of computing. Computer, 9(12): 58–65, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:SAI

[1210] Anonymous. Special 25th anniversary issue of the IEEE Transactions on Computers. Computer, 9(12): 66, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:CRI

[1211] Anonymous. Conference report 2nd International Conference on Software Engineering or a systems engineer by any other name …. Computer, 9(12): 67–71, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:RTS

[1212] Anonymous. Rickey’s tackling the SDK-8O microcomputer kit for his next science project [advertisement]. Computer, 9(12):72–73, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Eckhouse:1976:WRM

[1213] R. H. Eckhouse. Workshop report:

MICRO-9. Computer, 9(12):73–74, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:WOD

[1214] Anonymous. Workshop on Operating and Data Base Management Systems. Computer, 9(12):75, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:OC

[1215] Anonymous. The open channel. Computer, 9(12):76–78, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1976:NP

[1216] D. A. Michalopoulos. New products. Computer, 9(12):79–83, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:NL

[1217] Anonymous. New literature. Computer, 9(12):84, December 1976. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:RIA

[1218] Anonymous. Recent IC announcements. Computer, 9(12):85, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1976:NA

[1219] D. A. Michalopoulos. New applications. Computer, 9(12):86–87, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:JGC

[1220] Anonymous. Join the great computer roundup. Computer, 9(12): 88, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:RA

[1221] Anonymous. Repository [advertisement]. Computer, 9(12):89–95, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:Cb

[1222] Anonymous. Calendar. Computer, 9 (12):102–105, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:CPb

[1223] Anonymous. Call for papers. Computer, 9(12):103–104, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:AID

[1224] Anonymous. Advertisers index — December 1976. Computer, 9(12): 106, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:ICSb

[1225] Anonymous. IEEE Computer Society bylaws. Computer, 9(12):109– 114, December 1976. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:PRD

[1226] Anonymous. Pattern recognition data bases. Computer, 9(12):115, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:BR

[1227] Anonymous. Book reviews. Computer, 9(12):116–118, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1976:POB

[1228] Anonymous. Pick our brains [advertisement]. Computer, 9(12):118, December 1976. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fairman:1977:MRM

[1229] Michael Fairman, A. C. Weaver, and R. W. Catlin. MUMS — a reconfigurable microprocessor architecture.

Computer, 10(1):11–17, January 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wickes:1977:MCD

[1230] W. E. Wickes. Microprocessor chip design with the user in mind. Computer, 10(1):18–22, January 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Johnson:1977:AGC

[1231] Gearold R. Johnson and Robert A. Mueller. Automated generation of cross-system software for microcomputers. Computer, 10(1):23–31, January 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wakerly:1977:MCE

[1232] John F. Wakerly and Edward J. Mc-Cluskey. Microcomputers in the computer engineering curriculum. Computer, 10(1):32–38, January 1977. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Carey:1977:MLU

[1233] Bernard J. Carey. Microprocessor laboratory for a university environment. Computer, 10(1):40–46, January 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rony:1977:TMI

[1234] Peter R. Rony and David G. Larsen. Teaching microcomputer interfacing to non-electrical engineers. Computer, 10 (1):53–57, January 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srini:1977:FDM

[1235] Vason P. Srini. Fault diagnosis of microprocessor systems. Computer, 10 (1):60–65, January 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Krummel:1977:AMD

[1236] Larry Krummel and Gaymond Schultz.

Advances in microcomputer development systems. Computer, 10(2):13– 19, February 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peuto:1977:CIA

[1237] Bernard L. Peuto and Leonard J. Shustek. Current issues in the architecture of microprocessors. Computer, 10(2): 20–25, February 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wakerly:1977:MIA

[1238] John F. Wakerly. Microprocessor input/output architecture. Computer, 10 (2):26–33, February 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Allison:1977:DPM

[1239] Dennis R. Allison. Design philosophy for microcomputer architectures. Computer, 10(2):35–41, February 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morgan:1977:SMA

[1240] D. E. Morgan and D. J. Taylor. Survey of methods of achieving reliable software. Computer, 10(2):44–53, February 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Warren:1977:PHC

[1241] Jim Warren. Personal and hobby computing: an overview. Computer, 10 (3):10–22, March 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kay:1977:PDM

[1242] Alan C. Kay and Adele Goldberg. Personal dynamic media. Computer, 10 (3):31–41, March 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crane:1977:ODE

[1243] Hewitt D. Crane and Robert E. Savoie. On-line data entry system for handprinted characters. Computer, 10(3): 43–50, March 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adkins:1977:CST

[1244] Gerald W. Adkins and Udo W. Pooch. Computer simulation: a tutorial. Computer, 10(4):12–17, April 1977. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jayakumar:1977:SME

[1245] M. S. Jayakumar and Thomas M. Mc-Calla, Jr. Simulation of microprocessor emulation using GASP-PL/I. Computer, 10(4):20–26, April 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Braun:1977:PDE

[1246] Ludwig Braun, Stanley M. Finkelstein, and William B. Blesser. Pulmonary diffusion — a model study. Computer, 10(4):28–36, April 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chattergy:1977:IDV

[1247] Rahul Chattergy and Udo W. Pooch.

Integrated design and verification of simulation programs. Computer, 10 (4):40–45, April 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneidewing:1977:USE

[1248] N. F. Schneidewing. Use of simulation in the evaluation of software. Computer, 10(4):47–53, April 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Russo:1977:ICM

[1249] Paul M. Russo. Interprocessor communication for multi-microcomputer systems. Computer, 10(4):67–76, April 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bulman:1977:SCI

[1250] D. M. Bulman. Stack computers:

an introduction. Computer, 10(5): 18–28, May 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Blake:1977:ESA

[1251] Russell P. Blake. Exploring a stack architecture. Computer, 10(5):30–39, May 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Couch:1977:SSE

[1252] John Couch and Terry Hamm. Semantic structures for efficient code generation on a stack machine. Computer, 10(5):42–48, May 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schoenfeld:1977:MBL

[1253] Robert L. Schoenfeld, William A. Kocsis, Norman Milkman, and Gordon Silverman. Microprocessor in the biological laboratory. Computer, 10 (5):56–67, May 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Poppendieck:1977:MET

[1254] Mary Poppendieck and Edouard J. Desautela. Memory extension techniques for minicomputers. Computer, 10(5):68–75, May 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Su:1977:CHD

[1255] S. Y. H. Su. Computer hardware descriptive languages and their applications: An introduction and prognosis. Computer, 10(6):10–13, June 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lipovski:1977:HDL

[1256] G. Jack Lipovski. Hardware description languages: Voices from the Tower of Babel. Computer, 10(6):14–17, June 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stewart:1977:LCL

[1257] James H. Stewart. LOGAL: A CHDL for logic design and synthesis of computers. Computer, 10(6):18–26, June 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hill:1977:DLB

[1258] Fredrick J. Hill and Ben M. Huey. A design language based approach to test sequence generation. Computer, 10 (6):28–33, June 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ching:1977:ICG

[1259] Shingfat Stephen Ching and James H.

Tracey. An interactive computer graphics language for the design and simulation of digital systems. Computer, 10(6):35–41, June 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jordan:1977:ASH

[1260] Harry Frederick Jordan and Burton J.

Smith. The assignment statement in hardware description languages. Computer, 10(6):43–49, June 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shanker:1977:TCS

[1261] K. S. Shanker. Total computer security problem: an overview. Computer, 10(6):50–62, June 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Diffie:1977:ECN

[1262] Whitfield Diffie and Martin E. Hellman. Exhaustive cryptanalysis of the NBS Data Encryption Standard. Computer, 10(6):74–84, June 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schember:1977:MSS

[1263] Kurt Schember and John R. Rumsey. Minimal storage sorting and searching techniques for RAM applications: a tutorial. Computer, 10(6):92–100, June 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gunn:1977:BRU

[1264] Jeffery F. Gunn, D. J. Lynes, and Robert L. Pritchett. Bipolar 16K ROM utilizing Schottky diode cells. Computer, 10(7):14–17, July 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Panigrahi:1977:IES

[1265] G. Panigrahi. Implications of electronic serial memories. Computer, 10 (7):18–25, July 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Greenwood:1977:RTI

[1266] Walter T. Greenwood, Jr. Reliability testing for industrial use. Computer, 10(7):26–30, July 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Srini:1977:ATR

[1267] Vason P. Srini. API tests for RAM chips. Computer, 10(7):32–35, July 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1977:PTA

[1268] Edward F. Miller, Jr. Program testing: art meets theory. Computer, 10 (7):42–51, July 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Spetz:1977:MN

[1269] William L. Spetz. Microprocessor networks. Computer, 10(7):64–70, July 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Booth:1977:DCD

[1270] John M. Booth and John B. Schroeder.

Design considerations for digital image processing systems. Computer, 10 (8):15–20, August 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gambino:1977:EDI

[1271] Lawrence A. Gambino and Bryce L.

Schrock. Experimental digital interactive facility. Computer, 10(8): 22–28, August 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wells:1977:IIA

[1272] Donald C. Wells. Interactive image analysis for astronomers. Computer, 10 (8):30–34, August 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wilson:1977:MID

[1273] Robert Michael Wilson, Dieter L. Teuber, Douglas T. Thomas, Jimmy Ray Watkins, and Charles M. Cooper. MSFC image data processing system. Computer, 10(8):37–44, August 1977.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Andrews:1977:EDI

[1274] Harry C. Andrews. Educational digital image processing facility. Computer, 10 (8):48–53, August 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rohrbacher:1977:IPS

[1275] Donald Rohrbacher and J. L. Potter. Image processing with the Staran parallel computer. Computer, 10(8): 54–59, August 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Adams:1977:NCD

[1276] John Adams and Bob Wallis. New concepts in display technology. Computer, 10(8):61–68, August 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Caulkins:1977:CFM

[1277] David C. Caulkins. Critique of the F8 microprocessor. Computer, 10(8): 83–86, August 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maples:1977:RMA

[1278] Michael D. Maples and Eugene R. Fisher. Real-time microcomputer applications using LLL Basic. Computer, 10(9):14–20, September 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Zick:1977:RFA

[1279] Gregory L. Zick and Tzafrir T. Sheffer. Remote failure analysis of micro-based instrumentation. Computer, 10(9):30– 35, September 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Strait:1977:DMC

[1280] Bobby G. Strait, Michael E. Thuot, and J. P. Hong. Distributed microcomputer control system for a high-energy gas-laser facility. Computer, 10(9):36– 43, September 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hargreaves:1977:DPC

[1281] James C. Hargreaves and Richard A.

Krakowski. Distributed process control: a micro-mini marriage. Computer, 10(9):44–56, September 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jarvis:1977:SHE

[1282] R. A. Jarvis. Simple hand/eye experimental set-up for computer vision research. Computer, 10(9):65– 69, September 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stanchi:1977:OC

[1283] Luciano Stanchi. Open channel. Computer, 10(9):71–72, September 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fuller:1977:MEA

[1284] Samuel H. Fuller and William E. Burr.

Measurement and evaluation of alternative computer architectures. Computer, 10(10):24–35, October 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barbacci:1977:ECT

[1285] Mario R. Barbacci and Daniel P. Siewiorek. Evaluation of the CFA test programs via formal computer descriptions. Computer, 10(10):36– 43, October 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pollmann:1977:CDS

[1286] Robert E. Pollmann. Control Data 480 series microprogrammable computer family. Computer, 10(10):45– 53, October 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walsh:1977:FMC

[1287] Larry Walsh and Stan Tenold. Family of MIL-SPEC computers. Computer, 10(10):54–63, October 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1977:CSA

[1288] Ware Myers. Computer-aided seismic analysis and discrimination. Computer, 10(10):64–71, October 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Greene:1977:RCS

[1289] William Greene and Udo W. Pooch. A review of classification schemes for computer communication networks. Computer, 10(11):12–21, November 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Falk:1977:ADN

[1290] Gilbert Falk and John M. McQuillan. Alternative for data network architectures. Computer, 10(11):22– 29, November 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mamrak:1977:SMC

[1291] Sandra A. Mamrak and Patricia A.

DeRuyter. Statistical methods for comparing computer services. Computer, 10(11):32–39, November 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morgan:1977:SMI

[1292] D. E. Morgan, D. J. Taylor, and G. Custeau. Survey of methods for improving computer network reliability and availability. Computer, 10(11): 42–50, November 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yajima:1977:LOL

[1293] Shuzo Yajima, Yahiko Kambayashi, Susumu Yoshida, and Kazuo Iwama. Labolink: an optically linked laboratory computer network. Computer, 10 (11):52–59, November 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jones:1977:DMC

[1294] S. K. Jones and R. K. Cavin, III. Dynamic modeling and control of digital communication networks. Computer, 10(11):60–64, November 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rosen:1977:USP

[1295] Charles A. Rosen and David Nitzan.

Use of sensors in programmable automation. Computer, 10(12):12–23, December 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nevins:1977:RAA

[1296] James L. Nevins and Daniel E. Whitney. Research on advanced assembly automation. Computer, 10(12): 24–38, December 1977. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Woo:1977:PSM

[1297] Tony C. Woo. Progress in shape modeling. Computer, 10(12):40–46, December 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Voelcker:1977:GMM

[1298] Herbert B. Voelcker and Aristides A. G. Requicha. Geometric modeling of mechanical parts and processes. Computer, 10(12):48–57, December 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Balde:1977:NCC

[1299] John W. Balde and Daniel I. Amey. New chip carrier package concepts. Computer, 10(12):58–68, December 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mulder:1977:CSE

[1300] Michael C. Mulder. Computer science and engineering education. Computer, 10(12):69–80, 89–133, December 1977. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Enslow:1978:WLD

[1301] Philip H. Enslow, Jr. What is a “distributed” data processing system? Computer, 11(1):13–21, January 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Jensen:1978:HED

[1302] E. Douglas Jensen. Honeywell experimental distributed processor — an overview. Computer, 11(1):28–38, January 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peebles:1978:SAD

[1303] Richard Peebles and Eric Manning.

System architecture for distributed data management. Computer, 11(1): 40–47, January 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Forsdick:1978:OSC

[1304] Harry C. Forsdick, Richard E. Schantz, and Robert H. Thomas. Operating systems for computer networks. Computer, 11(1):48–57, January 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1978:ICG

[1305] Ware Myers. Interactive computer graphics: poised for takeoff? Computer, 11(1):60–74, January 1978. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gosling:1978:ERT

[1306] J. B. Gosling, S. H. Lavington, and J. M. Kelley. Extending registertransfer technology to teach computer architecture. Computer, 11(1):76– 80, January 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1978:NSE

[1307] Ware Myers. Need for software engineering. Computer, 11(2):12–26, February 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Maryanski:1978:SDD

[1308] Fred J. Maryanski. Survey of developments in distributed data base management systems. Computer, 11(2): 28–38, February 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Boyd:1978:IMS

[1309] Donald L. Boyd. Implementing mass storage facilities in operating systems. Computer, 11(2):40–45, February 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Scheuermann:1978:DED

[1310] Peter Scheuermann. On the design and evaluation of data bases. Computer, 11 (2):46–55, February 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crow:1978:SCG

[1311] Franklin C. Crow. Shaded computer graphics in the entertainment industry. Computer, 11(3):11–22, March 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fisher:1978:DCP

[1312] David A. Fisher. DoD’s common programming language effort. Computer, 11(3):24–33, March 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bass:1978:PFS

[1313] Charlie Bass. PLZ: a family of system programming languages for microprocessors. Computer, 11(3):34–39, March 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Brown:1978:MPP

[1314] William L. Brown. Modular programming in PL/M. Computer, 11(3): 40–46, March 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1978:PT

[1315] Edward F. Miller, Jr. Program testing. Computer, 11(4):10–12, April 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fairley:1978:TSA

[1316] Richard E. Fairley. Tutorial: static analysis and dynamic testing of computer software. Computer, 11(4): 14–23, April 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Huang:1978:PIS

[1317] J. C. Huang. Program instrumentation and software testing. Computer, 11(4):25–32, April 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

DeMillo:1978:HTD

[1318] Richard A. DeMillo, Richard J. Lipton, and Frederick G. Sayward. Hints on test data selection: help for the practicing programmer. Computer, 11 (4):34–41, April 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Panzl:1978:AST

[1319] David J. Panzl. Automatic software test drivers. Computer, 11(4): 44–50, April 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Darringer:1978:ASE

[1320] John A. Darringer and James C. King. Applications of symbolic execution to program testing. Computer, 11(4): 51–60, April 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bloch:1978:CPE

[1321] Erich Bloch and Dominick Galage. Component progress: its effect on high-speed computer architecture and machine organization. Computer, 11 (4):64–76, April 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davidson:1978:OFE

[1322] Scott Davidson and Bruce D. Shriver. Overview of firmware engineering. Computer, 11(5):21–33, May 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stockenberg:1978:VMP

[1323] John Stockenberg and Andries van Dam. Vertical migration for performance enhancement in layered hardware/firmware/software systems. Computer, 11(5):35–50, May 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Barbacci:1978:UEV

[1324] Mario R. Barbacci and Alan Parker.

Using emulation to verify formal architecture descriptions. Computer, 11 (5):51–56, May 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Fiala:1978:MS

[1325] Edward R. Fiala. The MAXC systems. Computer, 11(5):57–67, May 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

URL http://research.microsoft.
com/~lampson/Systems.html#maxc.

Glantz:1978:CCM

[1326] Stanton A. Glantz. Computers in clinical medicine: a critique. Computer, 11(5):68–77, May 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Su:1978:HDL

[1327] S. Y. H. Su. Hardware description language applications: An introduction and prognosis. Computer, 11 (6):10–13, June 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

McKenzie:1978:EPD

[1328] Ken McKenzie. Evolution of peripheral devices in microprocessor systems. Computer, 11(6):12–17, June 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Morse:1978:IME

[1329] Stephen P. Morse, William B. Pohlman, and Bruce W. Ravenel. Intel 8086 microprocessor: a 16-bit evolution of the 8080. Computer, 11 (6):18–27, June 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Garland:1978:PCG

[1330] Harry Garland and Alice Ahlgren. Personal computers go to work. Computer, 11(6):28–32, June 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Magers:1978:MSD

[1331] Celeste S. Magers. Managing software development in microprocessor projects. Computer, 11(6):34–42, June 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rauscher:1978:UAM

[1332] Tomlinson G. Rauscher. Unified approach to microcomputer software development. Computer, 11(6):44–54, June 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Alexandridis:1978:SMB

[1333] Nikitas A. Alexandridis. Special monograph: bit-sliced microprocessor architecture. Computer, 11(6):56–80, June 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lehmann:1978:NSB

[1334] John R. Lehmann. NSF support of basic research in computer science. Computer, 11(7):10–13, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siewiorek:1978:ULM

[1335] Daniel P. Siewiorek, Donald E.

Thomas, and Don L. Scharfetter. Use of LSI modules in computer structures: trends and limitations. Computer, 11 (7):16–25, July 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kartashev:1978:DAP

[1336] Steven I. Kartashev and Svetlana P.

Kartashev. Dynamic architectures: problems and solutions. Computer, 11 (7):26–40, July 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Borgerson:1978:MIO

[1337] Barry R. Borgerson, Garold S. Tajden, and Merlin L. Hanson. Mainframe implementation with off-the-shelf LSI modules. Computer, 11(7):42–48, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rennels:1978:RMC

[1338] David A. Rennels. Reconfigurable modular computer networks for spacecraft on-board processing. Computer, 11(7):49–59, July 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Reifer:1978:GST

[1339] Donald J. Reifer and Stephen Trattner. A glossary of software tools and techniques. Computer, 11(7):52–60, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Friedman:1978:ELT

[1340] Arthur D. Friedman and Luca Simoncini. Effect of LSI technology on the theory of modular computer design. Computer, 11(7):60–67, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Chu:1978:LMD

[1341] Yaohan Chu. LSI modular directexecution computer organization. Computer, 11(7):69–76, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Denning:1978:OSP

[1342] Peter J. Denning. Operating systems principles for data flow networks. Computer, 11(7):86–96, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1978:CDP

[1343] Harold S. Stone and Shahid H. Bokhari. Control of distributed processes. Computer, 11(7):97–106, July 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Powers:1978:MAB

[1344] V. Michael Powers and Jose H. Hernandez. Microprogram assemblers for bit-slice microprocessors. Computer, 11 (7):108–120, July 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Herlihy:1978:CHS

[1345] David Herlihy. Computation in history: styles and methods. Computer, 11(8):8–17, August 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lansdown:1978:CC

[1346] John Lansdown. Computer in choreography. Computer, 11(8):19–30, August 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ross:1978:CSL

[1347] Donald Ross, Jr. Computer-aided study of literary language. Computer, 11(8):32–39, August 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Truax:1978:CMC

[1348] Barry Truax. Computer music composition: the Polyphonic POD System. Computer, 11(8):40–50, August 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lipovski:1978:DDC

[1349] G. Jack Lipovski and Keith L. Doty.

Developments and directions in computer architecture. Computer, 11(8): 54–67, August 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1978:SFS

[1350] Ware Myers. Special feature: studying the oceans from space. Computer, 11 (8):68–75, August 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Caswell:1978:BT

[1351] Hollis L. Caswell, George A. Brown, Daniel H. Carter, Edward B. Eichelberger, William C. Holton, S. Krishnaswamy, William Lattin, Paul Losleben, Richard J. Petschauer, Herbert Stopper, and Robert M. Sullivan. Basic technology. Computer, 11(9): 10–19, September 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Apfelbaum:1978:CSO

[1352] Henry Apfelbaum, Richard P. Case, J. Egil Juliussen, Bruce Shriver, Harold S. Stone, and James E. Thornton. Computer system organization: problems of the 1980’s. Computer, 11 (9):20–28, September 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wasserman:1978:SET

[1353] Anthony I. Wasserman, Susan L. Gerhart, Edward F. Miller, Jr., L. A. Belady, William Waite, and William A. Wulf. Software engineering: the turning point. Computer, 11(9):30– 41, September 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Yao:1978:DS

[1354] S. Bing Yao, Philip A. Bernstein, Nathan Goodman, Stewart A. Schuster, David W. Shipman, and Diane C. P. Smith. Data-base systems. Computer, 11(9):46–60, September 1978.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pooch:1978:CSC

[1355] Udo W. Pooch, Richard H. Austing, Rahul Chattergy, and Michael C. Mulder. Computer science and computer engineering education in the 80’s. Computer, 11(9):69–83, September 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Isaacson:1978:PC

[1356] Portia Isaacson, Robert C. Gammill, Richard S. Heiser, Adam Osborne, Larry Tesler, and Jim C. Warren, Jr. Personal computing. Computer, 11(9): 86–97, September 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schnable:1978:RCI

[1357] George L. Schnable, Larry J. Gallace, and Henry L. Pujol. Reliability of CMOS integrated circuits. Computer, 11(10):6–17, October 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Waser:1978:HMM

[1358] Shlomo Waser. High-speed monolithic multipliers for real-time digital signal processing. Computer, 11(10): 19–29, October 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Nutt:1978:CSS

[1359] Gary J. Nutt. Case study of simulation as a computer system design tool. Computer, 11(10):31–36, October 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Trivedi:1978:AMC

[1360] Kishor Trivedi. Analytic modeling of computer systems. Computer, 11(10): 38–56, October 1978. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ohlson:1978:SDC

[1361] Mark Ohlson. System design considerations for graphic input devices. Computer, 11(11):9–18, November 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Preiss:1978:SCD

[1362] Richard B. Preiss. Storage CRT display terminals: Evolution and trends. Computer, 11(11):20–26, November 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Lucido:1978:ODB

[1363] Anthony P. Lucido. An overview of directed beam graphics display hardware. Computer, 11(11):29–36, November 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1978:SPP

[1364] Alan Jay Smith. Sequential program prefetching in memory hierarchies. Computer, 11(12):7–21, December 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1978:SAS

[1365] Ware Myers. Statistical approach to scheduling software development. Computer, 11(12):23–35, December 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Shemer:1978:AEO

[1366] Jack E. Shemer and J. Richard Keddy.

Architecture of an experimental office system: the Soft Display Word Processor. Computer, 11(12):39–48, December 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rochester:1978:CK

[1367] Nathaniel Rochester, Frank C. Bequaert, and Elmer M. Sharp. Chord keyboard. Computer, 11(12):57–63, December 1978. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bersoff:1979:SCM

[1368] Edward H. Bersoff, Vilas D. Henderson, and Stan G. Siegel. Software configuration management: a tutorial. Computer, 12(1):6–13, January 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bhandarkar:1979:CDM

[1369] D. P. Bhandarkar, J. B. Barton, and A. F. Tasch, Jr. Charge-coupled device memories: a perspective. Computer, 12 (1):16–23, January 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Casasent:1979:COC

[1370] David Casasent. Coherent optical computing. Computer, 12(1):27–28, January 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Treleaven:1979:EPC

[1371] Philip C. Treleaven. Exploiting program concurrency in computing systems. Computer, 12(1):42–49, January 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ramsey:1979:MES

[1372] R. L. Ramsey, R. L. Aylor, and R. D. Williams. Microcomputer-aided eating for the severely handicapped. Computer, 12(1):54–61, January 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Peuto:1979:ANM

[1373] Bernard L. Peuto. Architecture of a new microprocessor. Computer, 12(2): 10–21, February 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Wakerly:1979:IMM

[1374] John F. Wakerly. Intel MCS-48 microcomputer family: a critique. Computer, 12(2):22–31, February 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Landau:1979:SDT

[1375] Jack V. Landau. State description techniques applied to industrial machine control. Computer, 12(2):32– 40, February 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stritter:1979:MAC

[1376] Edward Stritter and Tom Gunter. Microprocessor architecture for a changing world: the Motorola 68000. Computer, 12(2):43–52, February 1979.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Cohen:1979:MML

[1377] Harvey A. Cohen and Rhys S. Francis. Macro-assemblers and macrobased languages in microprocessor software development. Computer, 12(2): 53–64, February 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bentley:1979:IAD

[1378] Jon Louis Bentley. Introduction to algorithm design. Computer, 12(2): 66–78, February 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Belady:1979:ES

[1379] Laszlo A. Belady. Evolved software for the 80’s. Computer, 12(2):79– 82, February 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hsiao:1979:DMC

[1380] D. K. Hsiao. Database machines are coming, database machines are coming! Computer, 12(3):7–??, March 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). Reprinted in A. R. Hurson, L. L. Miller, and S. H. Pakzad, Parallel Architectures for Database Systems, IEEE Computer Society Press, Washington, DC, 1989.

Su:1979:CDC

[1381] Stanley Y. W. Su. Cellular-logic devices: Concepts and applications. Computer, 12(3):11–25, March 1979.

CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Smith:1979:RDB

[1382] Diane C. P. Smith and John Miles Smith. Relational data base machines. Computer, 12(3):28–37, March 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hollaar:1979:TRC

[1383] Lee A. Hollaar. Text retrieval computers. Computer, 12(3):40–50, March 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Berra:1979:RAA

[1384] P. Bruce Berra and Ellen Oliver. Role of associative array processors in data base machine architecture. Computer, 12(3):53–61, March 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kerr:1979:DBM

[1385] Douglas S. Kerr. Data base machines with large content-addressable blocks and structural information processors. Computer, 12(3):64–79, March 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Burr:1979:BSM

[1386] William E. Burr, Dennis R. Allison, Alice C. Parker, Charles B. Silio, Jr., and A. Brinton Cooper. Bus system for the Military Computer Family. Computer, 12(4):11–22, April 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Dietz:1979:AEM

[1387] William Dietz and Leland Szewerenko.

Architectural efficiency measures: an overview of three studies. Computer, 12(4):26–33, April 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Stone:1979:LCA

[1388] Harold S. Stone. Life-cycle cost analysis of instruction-set architecture standardization for military computer systems. Computer, 12(4):35–39, 41–47, April 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Schneider:1979:PO

[1389] G. Michael Schneider. Pascal: an overview. Computer, 12(4):61–65, April 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Ravenel:1979:TPS

[1390] Bruce W. Ravenel. Toward a Pascal standard. Computer, 12(4):68–82, April 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1979:BT

[1391] Ware Myers. Barriers to technology.

Computer, 12(4):84–91, April 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mathur:1979:TIH

[1392] Francis P. Mathur. Technology’s impact on hardware reliability. Computer, 12(4):91–93, April 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Atkins:1979:FSC

[1393] D. E. Atkins. Fourth Symposium on Computer Arithmetic: crunching with quality and LSI. Computer, 12 (4):94–97, April 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Slana:1979:CE

[1394] Matthew F. Slana. Computer elements for the 80’s. Computer, 12 (4):98–102, April 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hoagland:1979:STC

[1395] Albert S. Hoagland. Storage technology: Capability and limitations. Computer, 12(5):12–18, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Welch:1979:AMH

[1396] Terry A. Welch. Analysis of memory hierarchies for sequential data access. Computer, 12(5):19–26, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Champine:1979:CTD

[1397] George A. Champine. Current trends in data base systems. Computer, 12 (5):27–41, May 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Astrahan:1979:SRR

[1398] M. M. Astrahan, M. W. Blasgen, D. D. Chamberlin, J. N. Gray, W. F. King, B. G. Lindsay, R. A. Lorie, J. W. Mehl, T. G. Price, G. R. Putzolu, M. Schkolnick, P. P. Selinger, D. R. Slutz, H. R. Strong, and P. Tiberio. System R: a relational data base management system. Computer, 12(5):42–48, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Hansen:1979:KAC

[1399] Per Brinch Hansen. Keynote address on concurrent programming. Computer, 12(5):50–56, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mori:1979:MJE

[1400] Ryoichi Mori, Yoshikuni Okada, Hiroaki Tajima, and Morihiko Tajima. Microprocessors in Japan — status in 1978. Computer, 12(5):58–63, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Mori:1979:MAJ

[1401] Ryoichi Mori, Yuzo Kita, Iwao Morishita, Kiyoto Ohkawa, Seishi Nishikawa, Akio Tojo, Akio Kobubu, Yoshikuni Okada, and Shunichi Uchida. Microcomputer applications in Japan. Computer, 12(5):64–74, May 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Joel:1979:CSU

[1402] Amos E. Joel, Jr. Circuit switching: unique architecture and applications. Computer, 12(6):10–22, June 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Kasson:1979:RCB

[1403] James M. Kasson. Rolm computerized branch exchange: an advanced digital PBX. Computer, 12(6):24–31, June 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Masson:1979:SCS

[1404] Gerald M. Masson, George C. Gingher, and Shinji Nakamura. Sampler of circuit switching networks. Computer, 12 (6):32–48, June 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Siegel:1979:INS

[1405] Howard Jay Siegel. Interconnection networks for SIMD machines. Computer, 12(6):57–65, June 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

El-Ayat:1979:III

[1406] Khaled A. El-Ayat. Intel 8089: an integrated I/O processor. Computer, 12 (6):67–78, June 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1979:ICGa

[1407] Ware Myers. Interactive computer graphics: flying high — 1. Computer, 12(7):8–12, 14, 16, 17, July 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Paul:1979:RMA

[1408] Richard Paul. Robots, models, and automation. Computer, 12(7):19–27, July 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Elmquist:1979:SSS

[1409] Kells A. Elmquist, Howard Fulner, David B. Gustavson, and George Morrow. Standard specification for S100 bus interface devices. Computer, 12 (7):28–52, July 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Frank:1979:PPC

[1410] Robert J. Frank. Patent protection and the computing arts. Computer, 12(7):66–73, July 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Davis:1979:DII

[1411] Ruth M. Davis. DoD initiative in integrated circuits. Computer, 12 (7):74–79, July 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:TCa

[1412] Anonymous. Table of contents. Computer, 12(8):1–2, August 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:SMa

[1413] Anonymous. Special messages. Computer, 12(8):3–4, August 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:ACa

[1414] Anonymous. About the cover. Computer, 12(8):5, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Miller:1979:SQA

[1415] E. Miller. Software quality assurance. Computer, 12(8):7–9, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Geiger:1979:PTT

[1416] W. Geiger, L. Gmeiner, H. Trauboth, and U. Voges. Program testing techniques for nuclear reactor protection systems. Computer, 12(8):10–18, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sorkowitz:1979:CTP

[1417] A. R. Sorkowitz. Certification testing: a procedure to improve the quality of software testing. Computer, 12 (8):20–24, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:SDCa

[1418] Anonymous. Sixth Data Communications Symposium 1979. Computer, 12(8):25, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Gannon:1979:EDU

[1419] C. Gannon. Error detection using path testing and static analysis. Computer, 12(8):26–31, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Holthouse:1979:EAT

[1420] M. A. Holthouse and M. J. Hatch. Experience with automated testing analysis. Computer, 12(8):33–36, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bowen:1979:SSP

[1421] J. B. Bowen. A survey of standards and proposed metrics for software quality testing. Computer, 12 (8):37–42, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Buckley:1979:SSQ

[1422] F. Buckley. A standard for software quality assurance plans. Computer, 12 (8):43–50, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:AIC

[1423] Anonymous. Announcing … the IEEE Computer Society’s Tutorial Week79. Computer, 12(8):49, August 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1979:ICGb

[1424] W. Myers. Interactive computer graphics: Flying high. Part II. Computer, 12 (8):52–67, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Tjaden:1979:SFS

[1425] G. Tjaden and M. Cohn. Special feature some considerations in the design of mainframe processors with microprocessor technology. Computer, 12 (8):68–74, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:ICD

[1426] Anonymous. The 1st International Conference on Distributed Computing Systems. Computer, 12(8):75–78, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Myers:1979:NWS

[1427] W. Myers. NCC ’79 wrap-up: The social implications of computers. Computer, 12(8):79–86, August 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:PCA

[1428] Anonymous. Public concerned about privacy. Computer, 12(8):82–83, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:OCa

[1429] Anonymous. The open channel. Computer, 12(8):87–88, August 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1979:NPa

[1430] D. Michalopoulos. New products. Computer, 12(8):91–95, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:NLa

[1431] Anonymous. New literature. Computer, 12(8):96, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:RIAa

[1432] Anonymous. Recent IC announcements. Computer, 12(8):97, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:RMAa

[1433] Anonymous. Recent microsystems announcements. Computer, 12(8):98, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1979:NARa

[1434] D. Michalopoulos. New applications & recent research. Computer, 12 (8):99, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:Ua

[1435] Anonymous. Update. Computer, 12(8):

100–103, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:Ca

[1436] Anonymous. Calendar. Computer, 12 (8):106–107, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:FIC

[1437] Anonymous. Fifth International Conference on Very Large Data Bases. Computer, 12(8):108–109, August 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:CPa

[1438] Anonymous. Call for papers. Computer, 12(8):110, August 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:BRa

[1439] Anonymous. Book reviews. Computer, 12(8):111, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:Cb

[1440] Anonymous. Computer. Computer, 12(8):c1, August 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Sunshine:1979:FTP

[1441] C. A. Sunshine. Formal techniques for protocol specification and verification. Computer, 12(9):20–27, September 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Walden:1979:EHP

[1442] D. C. Walden and A. A. McKenzie. The evolution of host-to-host protocol techniques. Computer, 12(9):29– 37, September 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:TCb

[1443] Anonymous. Table of contents. Computer, 12(10):1–2, October 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:SMb

[1444] Anonymous. Special message. Computer, 12(10):3, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:ACb

[1445] Anonymous. About the cover. Computer, 12(10):4, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anderson:1979:HTT

[1446] K. R. Anderson and H. A. Perkins.

Hardware test technology. Computer, 12(10):7–8, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Williams:1979:TLN

[1447] T. W. Williams and K. P. Parker.

Testing logic networks and designing for testability. Computer, 12(10): 9–21, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:CPb

[1448] Anonymous. Call for papers. Computer, 12(10):22, October 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Crafts:1979:TMT

[1449] J. M. Crafts. Techniques for memory testing. Computer, 12(10):23– 31, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Foley:1979:EMR

[1450] E. Foley. The effects of the microelectronics revolution on systems and board test. Computer, 12(10): 32–38, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Plice:1979:AAT

[1451] W. A. Plice. Automatic analog testing — 1979 style. Computer, 12(10): 40–47, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:CPc

[1452] Anonymous. Call for papers. Computer, 12(10):48, October 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Clary:1979:STC

[1453] J. B. Clary and R. A. Sacane. Selftesting computers. Computer, 12(10): 49–59, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:TW

[1454] Anonymous. Tutorial Week79. Computer, 12(10):56, October 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:SDCb

[1455] Anonymous. Sixth Data Communications Symposium 1979. Computer, 12 (10):60, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Korn:1979:MSF

[1456] G. A. Korn. Microdare special feature: Microdare: a fast, directexecuting high-level language system for small computers. Computer, 12(10): 61–71, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:CPd

[1457] Anonymous. Call for papers. Computer, 12(10):72–73, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:IMM

[1458] Anonymous. 1979 International Micro & Mini Computer Conference. Computer, 12(10):74, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Bork:1979:SFL

[1459] A. Bork. Special feature learning with computer simulations. Computer, 12 (10):75–84, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:ICS

[1460] Anonymous. 1980–1981 IEEE Computer Society scholarships for members of student branch chapters of the Computer Society. Computer, 12(10): 85–86, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:OCb

[1461] Anonymous. The open channel. Computer, 12(10):87–88, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1979:NPb

[1462] D. Michalopoulos. New products.

Computer, 12(10):89–93, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:NLb

[1463] Anonymous. New literature. Computer, 12(10):94, October 1979. CO-DEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:RIAb

[1464] Anonymous. Recent IC announcements. Computer, 12(10):95, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:RMAb

[1465] Anonymous. Recent microsystems announcements. Computer, 12(10): 96, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Michalopoulos:1979:NARb

[1466] D. Michalopoulos. New applications & recent research. Computer, 12 (10):97, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:MPP

[1467] Anonymous. Micro-12 preliminary program. Computer, 12(10):98–99, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:Ub

[1468] Anonymous. Update. Computer, 12 (10):100–103, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:Cc

[1469] Anonymous. Calendar. Computer, 12 (10):107–110, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:CPe

[1470] Anonymous. Call for papers. Computer, 12(10):108–109, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:BRb

[1471] Anonymous. Book reviews. Computer, 12(10):111, October 1979. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Anonymous:1979:Cd

[1472] Anonymous. Computer. Computer, 12 (10):c1, October 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Pferd:1979:IGT

[1473] W. Pferd, L. A. Peralta, and F. X.

Prendergast. Interactive graphics teleconferencing. Computer, 12(11):62– 67, November 1979. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).